A Complete Bibliography of Publications in Marine Policy (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 October 2023 Version 1.06

Title word cross-reference

$108.50 [Gos93b]. $250.00 [Gra90]. $40 [Tan90]. $49.50 [Gla91b]. $65.00 [Goo97]. $92.50 [Gla91a].

0 [Goo97, Low94]. 0-7748-0605-2 [Goo97]. 0-7923-2281-9 [Low94].

1 [Gau94, Pic94a]. 1-85333-7881 [Gau94]. 103-180 [Ano90l]. 105 [Kwi93a]. 111-196 [Ano97i]. 117-190 [Ano99d]. 117.00 [Lan90a]. 12 [Bar90b]. 136pp [Pre91]. 147-228 [Ano91j]. 147-239 [Ano92h]. 155-232 [Ano93p]. 17 [Nau94]. 17.55 [Fel90b]. 171-248 [Ano95d]. 175-280 [Ano98g]. 183-281 [Ano96k]. 183-296 [Ano90m]. 19.95 [Til91]. 191-270 [Ano99e]. 1960s [Sch98b]. 197-288 [Ano97j]. 1970s [Sch98b]. 1987 [Wat90c, Wat90d]. 1988 [Dub91b, Gla91a]. 1989 [Ano90g, Dub91a, Fre90a, Gul90, Kim90, Law90b, Phi90, Red90e, Wec90]. 1990 [Bar90b, KW91, Kwi90a, Pon91, Red91c, Sto91a]. 1991 [Bir92, BC92, Haa92, McE91b, Nei91]. 1992 [Ano92g, Bro93c, EHP93, Law93b, McE93a, Nei92]. 1993 [Ano93j]. 1994 [Jis95, Pic94a]. 1995 [Gil96].

2 [Goo97, Gul90, Ijl90]. 200-mile [Sco90]. 21 [Ano94h, Lau93, Nol93, OZ94, Won21]. 219-288 [Ano94n]. 21st [Kwi93b, Ste93a]. 231-300 [Ano91l]. 235-326 [Ano93q]. 243-335 [Ano92i]. 257pp [Pre90]. 26 [Jis95, McE93a]. 27 [EHP93]. 271-536 [Ano99f]. 272pp [Gla92]. 281-436 [Ano21b]. 29 [Nei91]. 291-359 [Ano94o]. 299-376 [Ano90o]. 29th [Ano91r, Bar90b]. 2nd [Kwi90a].

3 [App94b]. 30.00 [Law90a]. 303-380 [Ano91m]. 321pp [Kim93a]. 324 [Nau94]. 332-488 [Ano93s]. 338-408 [Ano92j]. 347pp [Law93c]. 35.00 [Pre90, Wat90a]. 357-438 [Ano96l]. 362pp [Tre90]. 363-439 [Ano94q]. 371-448 [Ano95f]. 378-463 [Ano90p]. 383-467 [Ano91n]. 39 [Bla91].

4 [Ano90g]. 409-491 [Ano97k]. 411-496 [Ano92k]. 41st [Phi90]. 437-544 [Ano98h]. 439-524 [Ano96m]. 44/225 [Ano91u]. 467-544 [Ano90q]. 484pp [Kwi93a]. 493-549 [Ano97l]. 497pp [Low90]. 4th [Nei92].

537-600 [Ano99g]. 572pp [Van92].

6 [Wat90c]. 6.50 [Gra91].

7 [Wat90d]. 73 [AO93]. 75-144 [Ano91o, Ano92l]. 75-152 [Ano93t]. 7881 [Gau94].

8 [Kwi90a]. 83-166 [Ano95g]. 83-174 [Ano98i].

9 [Low94]. 92 [App94b]. 92-64-13958-3 [App94b]. ’95 [McC99]. 99 [Gos93a]. 99-181 [Ano96n].

aan [Wec90]. abandonment [SBH93, Sid97]. Aberdeen [Ano93a]. Access [HRS97, Lod92, Sha97]. accountability [Mac93a]. achievements [Hey92]. achilles’ [Day95b]. ACOPS [Ano90g]. Act [AO91, Gar93, Ket95]. Action [CVM95, Col90, SE97a, Ros90, Wil96]. activities [Bud99, GY21, Wil96]. acts [Day95b]. Adalberto [Pre91]. adaptive [Fre90a]. address [oCKG90]. adjustment [GW91, Nie92]. administration [Kes97b, Sta94, Sun94]. advanced [HKP91]. Advancement [Lon91]. advantage [CM97]. Advisory [GSG93]. Aegean [Kar90, Van96]. Affairs [HP98, Gla91a, Ijl90, Kwi90b, Kwi90c, Kwi93b]. affecting [MA91]. Africa [Dio98, Cis92, CP99, Dev93, Keb90, Kwi93b, Nau95, SA98]. African [Dev94, Ekp98]. After [CK93b, Bis93, Pic94a, Rim97]. against [Gil91]. age [MT96]. Agency [Law90b, Hey92]. Agenda [Ano94h, Bár94a, BM96, Kim93b, Tol90, Lau93, Nol93, OZ94, Won21].

Agreement [Ano96a, CVM95, Bro95a, Gil96]. agreements [Coo94, Lan90a, WC99]. AGREH [Dio98]. aircraft [Roa96]. al [Gos93a]. Alaskan [Law93a]. Alastair [Gla92]. albacore [Ano91s, FS98]. Albatross [Ber97, BE99]. Alexander [Lon91]. alliance [dV90b]. allocation [Arm98, FGS95, Mor95]. allocations [Cad96]. Allowable [Kar96]. Almquist [Tre90]. along [Bav96]. Alpha [Pic94a]. alternative [dVdLE97]. Alternatives [Pic94b]. America [Bro92, Gab90, Tan90]. American [Lon91, Bis95]. among [NR99]. Amsterdam [Bla93]. anachronism [Gau95]. analysis [Arm98, DV90a, FGS95, Hil97, Kes97a, MU92, SN96, SN97, Smi95b, SP97b, SE97b, SM96, Thé97, Pre91]. Andresen [BC92]. Angeles [Smi94]. Angeles/Oxford [Smi94]. Anglo [But90b]. Anglo-Soviet [But90b]. Anniversary [Law90b]. Announcement [Ano94a]. Annual [Nei92, Phi90, Wec90, WLD98, Gla91a, Ijl90]. anomalies [Lan90b]. Antarctic [Sut90, Hof93, Sto91a]. apathy [Won21]. appendices [Bir94b]. application [BB96, Bud99, FGS95]. applied [Arm98]. Apportunity [Won21]. appraisal [Fog95]. approach [AM98, Cad96, CVM95, Hey91, Kes97b, Le 92, Mit21, Noe93, Nor94, Pet99, Pom91, Pot99]. approaches [ZC90]. April [Nei92]. aquaculture [BW94, Sor92]. aquatic [Nau94]. arbitrated [Her95]. arbitration [Gol90, Mar93]. Archaeological [Hay96, Hut96]. archipelagic [KA91, Rot90]. archipelago [Val94]. archipelagos [Cre99]. Arctic [Gla90a, JD92, But90a, Hue95, KV90, NSS96, RL92]. area [CK93a, Ano91r]. areas [BB96, Bro93c, Far96, Kba94, Tow94, dVF94]. arising [Lei93]. armaments [Bar90a]. arms [Bat21]. arrangement [AB97, Lod98]. arrangements [QJZ97]. Arroyo [Bro93a]. article [Bro96b, Ver97]. Artificial [Pic96]. Artisanal [Haa92]. ASEAN [Jus98, Tan98]. Asia [Bat21, Ken92, McE92, Sos94]. Asian [Smi94, LV95, McE93a]. aspects [Bir94b, Dav90c, Don93, Eik93, Gwi93, KAP96, Lau93, Lon90, Pla90, Sor92]. assembly [Ano91u]. Assessing [Lei93]. Assessment [Gos93a, Bud99, CK93b, JD92, McE91b, Pay94b, Sau98]. Assistance [Töp90]. Association [Lon91, Nei92, O’K96, Gra91, Noe93]. Atlantic [Dev96, Fel90a, FS98, HPB92, O’R95, Sch95, Wet94]. atlas [Bla93, Gla92, Smi94]. Atmospheric [Ald92]. attitudes [Ber90b]. attitudinal [YSM96]. August [Kim90, KW91]. Australia [BE99, BM96, Jon97, RH96, SP96]. Australian [CM97, Eva96]. authorities [Ano94h, BW94, Kba94]. authority [Noe93, Ram90]. authors [Ano90j, Ano90k, Ano91h, Ano91i, Ano93k, Ano93l, Ano93m, Ano93n, Ano93o, Ano94i, Ano94j, Ano94k, Ano94l, Ano94m]. Autonomy [Høn98]. available [Hey92, SP96]. avenues [Van96]. average [WLD98]. award [Mar93]. awareness [Bin91].

B [Bro93b, Lon91, Low90, Sco90]. Back [Boi94]. background [Cor92]. balance [Böh92, Cou99]. Ball [Bal99a]. Baltic [SE97a, Töp92]. Baltica’ [Ber90a]. ban [Bai97]. Bandung [Kwi90a]. Bank [App94a, Law93b, Nau94, McE93b]. Bankable [Tow92]. Banks [Day95b, Mit97]. Barents [Bro93c, Eik93]. Bargaining [Arm98]. Barrier [Kri91]. Barriers [Fel90a, Nie92]. based [Boy92a, CS97, Cre99, Dah93, Dou93, GP99a, GP99b, Küt94, NV97, Nol93, Pom91, PC97, oUotCEP92, SHG96, Wet94, Wil96, Won21, WT94]. Basic [Boc91]. basics [Boi94]. basin [Mar91]. basis [GY21]. Basle [Cot90]. Bay [PP95, Rim97]. be [Dev93, Law93a, NSS96, Won96]. beaches [GRBC98]. bed [Nol91, Tre90, Kwi93a]. Bedford [Kap99]. before [Rim97]. Belgium [Sme90]. Belhaven [Gos93a]. Benefit [Gos93a, FGS95, SM96]. benefit-cost [SM96]. benefits [HRS97, WNJ99, tCCT91]. Benin [Haa92]. Berkeley [Smi94, Sut90]. Berkeley/Los [Smi94]. best [SP96]. between [Hil97]. Beyond [Kin94]. biodiversity [Sch98a]. bluefin [BH94]. Board [Ano90h, Ano91e, Ano92f, Ano93i, Ano94f, Ano95b, Ano96g, Ano97f, Ano98e, Rei21]. Bombay [Mut94]. Book [AO93, Ano93a, App94b, App94a, Bir92, Bir94b, Bla91, Bla93, BC92, Bro93b, Bro93a, Bro93c, Clé96, Dub91a, Dub91b, Fel90b, Gau94, Gla90a, Gla90b, Gla91a, Gla91c, Gla91b, Gla92, Gol91, Goo97, Gos93b, Gos93a, Gra91, Gra90, Hey92, Ijl90, JD92, Kim93a, Kwi93a, Lan90a, Law90a, Law93c, Law93b, Lon91, Low90, Low94, McE91c, Pre91, Pre90, Sco90, Smi94, Sut90, Tan90, Til91, Tre90, Van92, Wat90a, Wat90c, Wat90d]. border [vdM94]. Borgese [Wat90c, Wat90d]. Boston [Bro93b, Bro93a, Kwi93a, Low90]. Boundaries [BGW91, Sch96, Bla91]. boundary [Chu93, Her95, McE92]. Box [Hey92]. Braer [Ano93a, Lar93, Lei93, Rit93]. Brassey [Til91]. breakthrough [Bro95a]. Brent [Sid97]. Brest [MP97]. Britain [Bud99, Jon99]. British [Fel90b, Gol91, GPT96, Rei21, Smi99b]. Brown [Kwi93a]. Brunswick [CM91]. Budislav [Gla90b]. Building [LV95, DN98, Kul98, Oke98, Per98, SA98, SS98]. bulk [MT96, Pic96]. Bureau [Sco92]. burning [Sun96]. Butterworth [Ano93b, Ano94b]. Butterworth-Heinemann [Ano93b, Ano94b]. Butterworths [AO93]. buyback [HGG99]. by-products [Pic96]. bycatch [Ber97].

C [Gos93a, Low94, Mai98, Bro97a]. C$22.95 [Gol91]. C$39.95 [Gol91]. C-SPODP [Mai98]. CA [Phi90, Sut90]. cables [Wag95]. Calcutta [Mut94]. Calendar [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano91a, Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94c, Ano94d, Ano94e, Ano95a, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano21a]. California [Smi94, Sut90]. call [SL99]. Cambridge [Dub91b, Gla90a, Gla91c]. Canada [JD92, Jus98, Tan98, Cri95, Day95a, Day95b, HPB92, Mai98, Mar93, Nor94, O’R95, Sch95, Gla90a]. Canada-France [Mar93]. Canadian [Bar92, tCCT91, Fel90a, Hue95, Mit21, Mur94, Sch98b, Gol91, AO93]. Canary [LRA96]. canned [Pay94a]. Capacity [Kul98, DN98, Oke98, Per98, SA98, SS98]. capitalism [Jak99]. cargo [MT96]. cargoes [Gol90]. Caribbean [Bla98, BP99, CMM98, Gab90, SHG96]. CARICOM [BP99, CMM98, Sau98]. Cartagena [oUotCEP92]. cartel [Cam96]. case [AH97, Bav96, Bro93c, Bro21, Cre99, DMV99, GGIMSRVL96, Gla90b, HPB92, Kap21, Kar90, Ken97, Mac93a, Mai98, Mar92, PP95, Sch96, SP96, Sue94, Wet94, WNJ99, WBN91, WC99]. casebound [Gla91a]. Catch [Kar96, MU92, WLD98]. catchability [WLD98]. cause [Gar90]. causes [Kin95]. cautious [CP99]. CEC [Law93b, Nau94]. celebrate [Law90b]. cement [Pic96]. Central [McE91b]. Centralised [Cre99]. Centre [Bir94b]. century [Kwi93b, Ste93a]. Certificates [Gar93]. CFRAMP [Sau98]. Chairman [Dav90b]. challenge [JD92, SP97c, Tol90, Won96]. challenges [GPT96, Hey92]. change [Fre90a, Kul94, Pre91, Rit93, Tab94, Tho94, Whi90]. changes [Law90a, Pic94a, Pre91]. Changing [Cou99, Sto91b, Rid97]. Chaos [Fog95, FHG97, HG96, Kes95, WAMK94, WAK96, PM96]. Chapter [Lau93, Ano94h]. chart [KPG95]. charterers [Hol93]. Chicago [Wat90c, Wat90d]. Chief [Bir92, Low90]. China [HS96, Sto94, Sun95, Ano90f, Gao91, Key99, hBS90, hS99, Sun95, WyZ92, Yu91, ZC90, ZC91, Zhi91, Kim93a]. choices [BJ90]. Churchill [Bro93c]. CIDA [BKLM98, HB98, HBMV98]. CIMDS [Ano90f]. circumstances [Pre90]. Citizen [GSG93]. citizens [Sko94]. city [Kba94, Sue94]. civil [Mak92]. claims [GY21, Nad92, Red90e, RH96]. Clarendon [Kim93a, Pre90]. Classification [Boi94]. clean [Bar92]. climate [Fre90a]. climatic [Tab94]. Co [BP99, DvV96, JMW21, Kap99, Kap21, Kes97a, KA94, NV97, PP95, PB97, SN96, SN97, dVdLE97, Fel90b]. Co-management [BP99, DvV96, JMW21, Kap99, Kap21, Kes97a, KA94, NV97, PP95, PB97, SN96, SN97, dVdLE97]. Co-operative [Fel90b]. coal [Pic96]. coal-fired [Pic96]. coast [Bav96, CM97, Stu94]. Coastal [Ano96f, Bal99b, Edw94, Gib93, MA91, Rot92, Str94, ZC90, BW94, Dah93, GRBC98, Gos93a, Kba94, KA94, KAP96, Nor94, Pet99, Pic94b, PC97, Pot94, Pot99, SE97b, Tab94, Tow94, WNJ99, WT94, EHP93, Pot99]. Coasts [Gra91]. cod [NSS96]. cold [VH97]. collaboration [App94a]. collection [Lan90a]. collective [Bav96, Hat97]. College [Bir94b]. Columbia [Fel90b, Gol91]. combat [Sta93]. commentaries [Lan90a]. Commentary [SN97, Bir92, Low90]. Comments [PM96]. commercial [BE99, Bir94b, CM91, Kap99, Red91c]. Commission [Ano91r, Kim90, KW91, Phi90, Oke98, dM92]. Committee [Bar90b]. committees [SP97c]. Common [Til91, Bav96, Eyt96, Gra91, Gwi93, Mat92, McE91a]. Commonwealth [Bla98]. Community [BP99, CMM98, Dou93, PC97, CK93b, Cre99, Fel90b, Fog95, Kes95, PM96, Pom91, WAMK94]. Community-based [Dou93, PC97, Cre99, Pom91]. Comparative [CM97, Gla90b, Kes97a, MU92, SN96, SN97]. compared [DvV96]. Compensation [Red92, RG90, Wal93]. competition [Dav90a, Han95, Stu90]. competitiveness [Tha21]. compiled [Ano93a, Gla91a]. complexity [Fog95, HH98, Kes95, PM96, WAMK94]. compliance [Kap21]. components [Bis95]. Comprehensive [SE97a, JB90]. concept [KA91, Wat90b]. concern [Gar90]. concerning [Ano91r, Clé96]. concerns [Red91c]. Conch [Kap21]. Conclusions [Ano90g]. conditions [Lod92]. Conference [Jis95, Nei92, Nei91, Now93, Smi93, Ano91r, Bar95, Law90b]. conferences [Clé96]. confidence [FS98]. conflict [Bis93, Rid97]. conflicting [PW92]. Conflicts [Hil97, Cha92b, Sto94, HS96]. Congressional [Ano91f]. Consequences [GP99a]. conservation [Ano91s, Cam94, CK93a, Hof93, Sah95, SP96]. conservation-related [SP96]. considerations [Now92]. containerized [Mar91]. Contaminated [vdM94]. Continental [Kwi93a, Gra90]. continuing [Rid97, Sut93a]. contractual [Mak92]. contribution [Dev96, Hun99]. Contributions [Gla90b, Joy96]. Control [Dun93, McI93, dV90b, Bav96, Bel93, Küt94, Low94, Mac93a, Pay94b, Sme90, Sut93b]. Controlling [Mut94, Bar90a]. controls [SBH93]. controversy [Sut93a, WD91]. convenience [Mar91, Egi90]. Convention [Ano91d, Bir92, Hof93, Kim90, KW91, Low90, Ago91, And94, Ano91p, Bir94a, Bro93e, Bro97b, Dav90c, Gao91, Gil91, Kol93, Nad92, oUotCEP92, Won96, Bro95a, Bro96b, Cot90, Hay96, KK91, O’K96, Sha92, Ver97]. conventions [Pay94b, Hey91, Bro93a]. conversion [Bar90a]. cool [Sun96]. cooperation [Bla98, Han95, Küt94, Kwi90b, Kwi90c, Sor92, Sta93, vE94, BC92]. Cooperative [Jus98, Coo94, QJZ97, Tow95a, Val90]. Copenhagen [Jis95]. Cornish [FS98]. Coromandel [Bav96]. corporate [Tow95a]. corresponding [Won96]. Cost [Pot94, FGS95, SM96]. cost-benefit [FGS95]. Cotonou [Haa92]. Council [Gil96, Bár94a, GSG93]. countries [DMV99, Mar97, Töp90, App94b, Law93c]. Couper [Gla92, Law90a]. CPMS [Jus98]. CPMS-II [Jus98]. creation [Sut90]. Crisis [Gar93, O’R95, Dic92, FS98, Sch95, Gra91]. critical [dVdLE97]. Cross [vdM94]. Cross-border [vdM94]. cruise [Mar92]. cultural [BB96, Bri96, Bro96b, Clé96, Cou96, Hut96, O’K96]. Current [CMM98, Red90a, Red90b, Red90c, Red91a, Red91b, SBH93, Gar90, Jon99, Sch95, Boy92a, YR94, Clé96, Gos93b]. Cushing [Dub91b]. cycle [Boy90, Jus98]. Cynthia [JD92].

D [Ano93a, Dub91b, Kwi93a, Law90a, Lon91, Pre90]. Dakar [Kba94]. Dallmeyer [Bla91]. damage [Kor92, Red92]. Dangerous [Gol90, Gab90]. Danish [Nie92, Nie94, NV97]. data [Lau93]. database [HP98]. David [JD92, AO93]. Davies [Ano93a]. day [McE91a]. DC [App94a, Lon91]. deadly [Bat21]. death [Sue94]. decade [Bis93]. December [Fre90a, Haa92]. Decentralization [dVdLE97]. Deciding [Bla91]. decision [Hil97, Sid97, PW92]. declaration [Ano91t, Ano94h]. decline [MG97]. decommissioned [DV90a]. Deep [Nol91, Tre90, Zhi91]. defence [Bri94, Pet99, Pot99]. Degli [Pre91]. delicate [Böh92]. delimitation [Chu93, Gra90, Mar93, Pre90]. Dell’Instituto [Pre91]. demurrage [Lax92]. Department [Sun94]. departments [Mar90]. dependent [Cre99]. design [Wet94]. detained [Bro21]. developed [Dur94]. developing [Kwi90b, Kwi90c, Mar97, Töp90]. Development [Ano90f, HPB92, Jak99, Kit90, Law90b, Mai98, Mak92, Sah95, Ano96f, BJ90, Bla98, CS97, Dur94, Hin92, Hun99, Joh92, Ken97, LS99, Lod98, MP97, Mit21, Nij90, Pom91, Rit93, Tri90, Whi90, ZC91, App94a, Fel90b, Law90a]. Developments [Blo94, Ago91, Arn93, BH94, Boy92b, Bud99, Clé96, Fra90a, GW91, Jon97, Lax92, Red90a, Red90b, Red90c, Red91a, Red91b, Sun95, Thé97, Won96, YCS95]. Deventer [Bro93a]. devolved [Hat97]. Dfl [Bro93a]. DFL170 [Gos93b]. Dfl395 [Bro93b]. Diego [Phi90]. dimension [Cox93]. Direct [FGS95]. directions [Fel90b, Kin99]. discards [CS94]. Discharge [Mor99]. disciplinary [Le 92]. discrimination [MT96]. Discussion [App94a]. Dispersion [RB99]. disposal [Gab90, Keb90, Lon90, SBH93, VH97]. Dispute [Bro97b, Val94, Van96]. disputed [Bro93c]. disputes [McE92]. distant [CM97]. distress [Dev96]. distribution [Mat92]. Distributional [Eik93]. Do [HGG99]. Dobbs [Gra90]. documentation [Boc91]. Documents [Ijl90, Gla91a, Lan90a]. doing [Hem98]. Dolphin [BE99]. Don [Gra91]. Donald [Gol91]. Donaldson [Cor95, Pla95, Smi95b]. Donat [Gla90a]. Dordrecht [Bir92, Bla91, Bro93b, Gos93b, Kwi93a, Lan90a, Low90, Sco90, Van92]. Dordrecht/Boston/London [Low90]. Dorinda [Bla91]. Dr [Hem98]. Draft [O’K96, Bro96b]. drawing [Bla91]. drawn [JM95]. dredged [vdM94]. Drift [WD91, Ano91r]. Drift-net [WD91, Ano91r]. driftnets [Ano91d, Ano91p]. Drug [Gil91]. drugs [Gil91]. dry [MT96]. dry-bulk [MT96]. Dublin [Bir94b]. Dumping [DV90a, Fre90b, CB91, Skj92]. dunes [GRBC98]. during [Sch98b]. dynamic [Tow95b].

EAFE [Nei92]. Earth [Bár94a, CS96]. Earthscan [Gra91]. East [FS98, Smi94, Bat21, CM97, MU92, Wet94, LV95, Sos94]. EC [Don93, Dur94, Mac93b, Nie94, Red90a, Red90b, Red90c, Red91a, dM90]. Ecological [Kor92]. Economic [Bat21, Hug92, Lon90, MS93, CM91, Gla91c, HRS97, Kar90, Kor90, MP97, Nad92, Nie92, Now92, Sco90, Gwi93]. economics [Cam94, Far96, Gos93a, Gos93b]. Economists [Nei92]. ecosystem [Bis95, Eik93]. ecosystems [Ale93, Lon91]. Edited [Bro93b, AO93, Bla91, Bro93a, Dub91a, Fel90b, Gau94, Gla90b, Gla91a, Gla92, Gol91, Gos93b, Hey92, JD92, Lan90a, Law90a, Lon91, McE91c, Smi94, Wat90c, Wat90d]. edition [Tan90]. Editor [Bir92, Low90, Kim94, Nic90, Seb90, Seb92]. Editor-in-Chief [Bir92, Low90]. Editorial [Ano98c, Ano98d, Ano99a, Bro93d, Bro94, Bro95b, Bro96a, Ano90h, Ano91e, Ano92f, Ano93i, Ano94f, Ano95b, Ano96g, Ano97f, Ano98e]. editors [Low90]. Eduardo [App94a]. Edward [Dub91a]. EEC [Bir94b]. EEZs [Ist93]. effective [SHG96, Skj92]. effectiveness [Pot94, BC92]. effects [CK93b]. efficiency [Fra90b]. effort [WLD98]. EFJ [Gar90]. eighth [KW91]. electronic [Ano93b, Ano94b, Edw94, HKP91, KPG95]. elements [Thé97]. Elisabeth [Wat90c, Wat90d]. elusive [Lév93]. Emerging [RL92, Boy90, Far96]. emphasis [HP98]. employment [MG97]. enclosed [Gla90b]. encyclopaedia [Gla92]. end [Skj92]. energy [Kwi93a]. enforcement [Mur94]. England [Kap21, Bal99b, Gib93, Pet99, PPJ95, Pot99, WC99]. entails [Sco94]. enterprise [Cha92a]. Environment [Bar90b, Gra91, Per98, Ano90g, Boy92b, Don93, Ekp98, Hug92, Kor92, Mah97, MA91, Noe93, Nol91, Smi93, Smi95b, Van90, Wil96]. Environmental [BW94, Dur94, McI95, Oki93, Red91c, Rit93, Sco92, SE97a, WMH99, Bud99, CVM95, Fen94, Gau94, GSG93, God94, Ist93, Joh92, JD92, Joy96, Ken97, Key99, Law93a, Lei93, MH92, Owe99, Pla96, Red92, RL92, Smi95a, Sue94, SP97c, dV90b]. environments [Kul94, Tab94, Gla91b]. era [MP97, Pic99]. Eric [Til91]. escalating [HH98]. establishment [Ist93]. estate [Kin94]. Estonia [Ano96a]. estuary [God94, Now93]. EU [Blo94, DMV99, Gez99]. Europe [Gil96, CD94, SM96]. European [Bir94b, Jis95, Nei92, Til91, dM92, Ber90b, Col90, CS94, HKP91, Kar96, Noe93, Pla95, Sco92, Sta93, Ste93b, dV90b]. Evaluating [WLD98]. Evaluation [KPG95, PC97, Sau98]. Evans [Pre90]. Evelyn [Fel90b]. events [Ano96i]. evidence [HGG99]. Evolution [Cot90, Low94, QJZ97, Sut90, Bis95]. evolving [Bal99b]. example [Skj92]. exclusion [Van91]. exclusive [Bar92, Gla91c, Kar90, Kor90, Nad92, Sco90]. exhibition [Nei91]. Expansion [Wat90a]. experience [AM98, Cla93, Law93a, Pot94, Ram90, SS98]. experiences [JM95]. exploitation [MH92, Zhi91]. explored [Ber90b]. export [WBN91]. exports [dV90b]. exposures [Lei93]. Expresso [Rot92]. Extended [Sch95, Dub91a, QJZ97]. extent [Lan90b]. externalisation [Gez99]. extinction [Sen96].

face [PW92]. facilities [Bal99a]. facility [Joh92, Per98]. factors [Nie92]. Faculty [Gla90b]. Failure [McE92, Sch98b]. fairness [HH98]. Falklands [Bis93]. Falklands/Malvinas [Bis93]. FAO [Law93b, Nau94]. farming [Jak99, Nei91]. Fatalities [NR99]. February [Clé96, Kim90]. federal [Mur94, RH96]. fees [HRS97]. Ferry [Gra90]. Fiji [Hun99]. financial [Kap99, Gar93]. Financing [Joh92]. finfish [WBN91]. Finnmark [AH97]. fired [Pic96]. first [BS99, Bro21, Clé96, Jis95]. fish [Arm98, Bar95, Cou99, Jak99, NSS96, Nei91]. Fisher [Bav96]. Fisheries [Bis93, Chu93, Her95, HP98, Kes97a, Mar97, McE91b, MH92, Mit97, Miz91, Nau94, Nei92, Sau98, SN96, SN97, Sen96, Tan98, Tow95a, AJ91, App94b, App94a, Arn93, BJ90, BW94, BH92, BP99, CMM98, CP99, Coo94, Cre99, Dio98, DvV96, Dub91a, Fel90b, Fog95, FGS95, GW91, Goo93, Gwi93, Hat97, Hey92, HG96, Hil97, Hin92, Høn98, Hun99, JM95, JMW21, Kes97b, KG21, KA94, Law90b, Law93c, Law93b, Lod92, Mac93b, Mar90, McE93a, McE93b, Moh91, Mor95, NV97, O’R95, ONAK92, PM96, PW92, Pom91, PP95, PB97, Pon91, Rob93, RB99, Sah95, Sch95, SCC^{+}98, SE97b, SP97c, Thé97, WLD98, Whi90, WAMK94, WC99, dVdLE97, lW90, oA93]. Fishermen [Haa92, Bin91, CS97, Nie94, NV97, Sch98b, YSM96]. fishermen-based [CS97]. fishers [SP96]. Fishery [Cha92b, FGA93, Aka93, AB97, BH94, CM97, CM91, Dou93, Fel90a, FS98, HPB92, HH98, HRS97, Hun97, McE91a, McE91b, Nie94, Now92, PPJ95, WyZ92, Yu91, Kap21, Lod98, Rei21]. fishes [Fel90a]. Fishing [Han95, Ano91d, Ano91p, Ano91r, AH97, BE99, Bir94b, CM97, DMV99, DT90, Eyt96, GGIMSRVL96, HGG99, Kap99, LRA96, Nie92, QJZ97, Squ94, Sto91b, Sum99, WD91]. FL [Fre90a, Gla91a, Ijl90]. flag [Bel93, Tha21, Egi90, Mar91]. fleet [HPB92, MG97]. fleets [CM97, Mur94]. Florida [Gos93a]. focus [Ale93]. Fonseca [Sco94]. food [Oki93]. force [MA91]. forecast [Cri95]. foreign [DT90, Mar91]. foreword [Seb90, Seb92]. Form [Sha92]. formulating [Sko94]. formulation [Mai98]. Forum [Law90b, Dev94, Hey92, McE93a]. forward [CK93b, Kim93b]. found [Hut96]. Foundation [Ber90a]. framework [Bal99b, Cha92b]. Frameworks [SP97b]. France [App94b, Mar93]. Francis [Gla91b, Tan90]. Francisco [Gla91c]. Freeze [Sun96]. Freight [Haw91, Böh92]. Fridjof [BC92]. Fridtjof [Sto91a]. frozen [Sut90]. Fundy [Coo94, Now92]. future [Aka95, AM98, AJ91, Boi94, Hun99, Jon99, Ket95, Law93a, Pay94a, Ros90, Sen96, Sid97, YCS95, BH92].

G [Ano93a, Bla91]. game [Sum99]. game-theoretic [Sum99]. Gap [Sut93a]. Garden [Jon97]. gas [Hug94, Pic99]. gear [Nau95]. Geir [Bro93c]. general [Ano91u, dC90, Kwi90b]. Genova [Pre91]. gentleman [WC99]. Geografiche [Pre91]. Geographical [Pre91]. geography [Smi99b]. geopolitics [SS97]. George [Low94]. Geostrategy [Bro97a, Jon97, MP97, Rid97, SS97, SP97b, Rim97]. geotactics [Jon97]. Gerard [Tan90]. gets [Nau95]. Ginsburg [Wat90c, Wat90d]. glance [Pay94a]. Glasgow [Nei91]. Global [Dav90c, Tab94, CD94, Fre90a, Pre91, SM96, Tol90, Wil96, Woo94, Joh92, Per98]. globalization [AH97]. Gold [Lon91]. good [KT93]. goods [Bar92]. Gordon [Gol91]. governance [AB97, CMM98, Tow95a]. Government [Byw95, Ste93a, Ano91s, Bel93, Lee90, Mur94, PB97, Sta94, WT94, Pla95]. governments [GN90]. Graduate [BKLM98]. Graham [Bro93b, Gau94]. Grand [Day95b, Mit97]. Great [Bud99, Kri91]. Greek [Kar90]. Greenfield [Kim93a]. greening [Red90d]. Greenpeace [Cur92]. Greenwich [Clé96]. groundfish [HPB92]. grounds [ONAK92]. Groups [Pot99]. Grove [Til91]. growth [Bat21, Kap99, Squ94]. guard [Stu94]. Guest [Seb90, Seb92]. guide [Ano96i]. guidelines [Ano93b, Ano94b, BB96, Bro96b]. Gulf [CVM95, Dic92, Her95, Mah97, Sco94]. Gwilliam [Gos93b].

H [Bro93b, Dub91b, Low90]. Hague [Gul90]. harbour [Noe93]. harbours [WMH99]. Hard [BJ90]. hardback [Gol91]. harm [Red92]. Harper [Gla92]. Harvest [Tan98]. Harvesting [PW92, Sha97]. harvests [KT93]. Hawkins [Ano93a]. Hazardous [Bar92, dM90, Ano90g, Gol90, Lon90, Tol90, Van90]. hazards [Sos94]. Headquarters [Kim90]. health [Cox93, Fen94, Sos94]. hedging [Haw91]. heel [Day95b]. Heinemann [Ano93b, Ano94b]. heritage [BB96, Bri96, Bro96b, Clé96, Cou96, Hut96, O’K96]. Herr [Hey92]. Herring [Rei21, Coo94, Now92, Rei21]. high [Han95, Pla96, Van91]. highly [Bar95]. Hinrichsen [Gra91]. Hiran [Van92]. historic [Hut96]. historical [Hay96, Smi99b]. history [Law93a]. Hole [Pon91]. Honiara [Hey92, Law90b]. Hormuz [Mah91]. House [Ano91f]. Hull [Now93]. human [JD92, Lei93]. Humber [Now93]. hunt [BC98]. hydrocarbon [YCS95].

IAEA [CB91]. IBC [Red90e, Red91c]. ice [Sun96]. ice-cool [Sun96]. Iceland. [Eyt96]. ICES [Gul90]. ICJ [Jan92, Sco94]. ICONA [Bla93]. If [Fel90a]. Ignacio [Bro93a]. II [Jus98, Tan98, Ano91p]. III [Gla91a]. illicit [Gil91]. imbalance [GP99a]. IMO [Clé96, Bar90b, BB96, Dij93, Won96]. impact [Bro93e, CK93b, HH98, SP96, Bir94a, Tab94]. impacts [Fre90a, GSG93, MS93, Oki93, Rit93]. Implementation [Pay94b, Ano94h, Bro96b, CS96, McC99, OZ94, Töp90, WMH99, Bro95a, Tan98]. Implementing [Key99, And90, BM96, SE97a, Lau93]. Implication [Dub91a]. Implications [Her95, Fre90a, GP99a, Kin95, MG97, Hey91, SS97, SE97b]. implies [Sco94]. Import [lW90]. improved [Fel90b]. Improving [God94]. inadequate [WLD98]. incentives [PW92, SHG96]. incident [Lar93, Lei93, Rot90]. including [Red91c]. Index [Ano90i, Ano91g, Ano92g, Ano93j, Ano94g, Ano95c, Ano96h, Ano97g, Ano99b, Ijl90]. India [YR94]. Indian [Bro97a, Kwi90c]. indicator [WLD98]. indicators [KG21]. indigenous [Sch96]. Individual [Cla93, SCC^{+}98, FGA93, Hat97, Tow92, YSM96]. Indonesia [McE91b, Ago91, Bai97, KAP96, Kwi90a, MU92]. Indonesian [KA91, Rot90, SS98]. industries [Squ94]. Industry [Rei21, AH97, Gar93, Hol93, Kap99, Ket95, Mor99, Nie92, Pic99, SE97a, Sid97, SL99, Ste93a]. informatics [Sch98a]. Information [Hey92, Edw94, LS99, Lau93, PW92, SP96]. initiative [HB98]. Initiatives [PP95, BKLM98, HBMV98, vdM94]. Innocent [Fra90a, HM99, Rot92]. innovation [Rei21]. inquiry [Kin95]. Inshore [SP97c, HPB92]. insights [Sha97]. instability [Gar90]. installations [SBH93]. Institute [Ano90f, BC92, Gla90b, Hey92, McE91c, McE91b, Sto91a, Wec90]. Institution [Pon91]. Institutional [KT93, Kwi90b, Kwi90c, SE97b, Don93, KAP96, Mar97, Wet94, Nie92]. institutionalize [Ken92]. institutions [ZC90]. instrument [Stu94]. instrumentation [HKP91]. Insurance [RG90, Sch98b, Gau94]. Integrated [GRBC98, Tow94, Ano96f, Dah93, Wat90b, Bal99b]. intellectual [GY21]. interdiction [Gil96]. interests [AO91]. interface [Cox93]. Intergovernmental [Wal93, Kri91, Oke98]. Interim [Ano93a]. International [tCCT91, Coo94, Gla90b, Nau94, O’K96, Pla90, Pre90, Sch96, Sor92, Tre90, Ano96i, Arn93, BH94, BC92, Boy90, Cam96, Dav90c, Dic92, Gla90a, Gla91c, Goo97, JM95, KT93, Kwi93a, Law93b, Pay94b, Per98, Pol90, RH96, Sco95, Skj92, hS99, Sue94, Töp90, Van92, And90, BGW91, Bro21, Küt94, Phi90, Sri90, Bro93b, Bro93a]. Introduction [Bar94b, Seb94]. Introductions [oCKG90]. Introductory [Pre91]. Investing [Cam94]. Involvement [Kba94]. IOMAC [Kwi90c]. Irish [Bir94b]. ISBN [App94b, Gau94, Goo97, Low94]. Island [Jon97, Ken97]. Islands [Hey92, Van92, LRA96, Law90a]. Isles [Law90b, Hal93, Smi99b]. issue [Mit97]. issues [Bir94a, CMM98, Chu93, Cou96, Dev93, Dev94, Gos93b, Jon99, KV90, McC99, Sun96, Val90]. Italy [Nei92, BGS92, GN90]. itineraries [Mar92]. ITQs [Eyt96]. IV [Bir92, Gla91b]. Iwan [Bal99a]. IWC [KT93]. ix [Gla91c].

J [Ano93a, Goo97, Smi94, Sut90, Van92]. Jakarta [McE91b]. Jamaica [Kim90, KW91]. January [Ano90n, Ano91k, Ano93r, Ano94p, Ano95e, Ano96j, Ano97h, Ano98f, Ano99c, Clé96, Kwi90a, Nei91, Ano92g, Ano93j]. Japan [Aka95, Ano91s, LV95, Sto91b, Val90]. Japanese [Aka93, BH92]. Japanese-Russian [Aka93]. Java [McE91a, McE91b]. Jayewardene [Van92]. Jeanette [Kim93a]. jobs [GP99a, GP99b]. Johnstone [Ano93a]. Joint [But90a, SE97a, Aka93]. Jorgen [BC92]. Joseph [Smi94, Wat90d]. judgment [Bro21]. judicial [Dev96]. July [Ano90o, Ano91l, Ano92g, Ano92i, Ano93j, Ano93q, Ano94o, Ano99f, Ano21b]. June [Phi90, Wec90]. jurisdiction [Low94, QJZ97, Sch95, Dub91a].

Kasoulides [Low94]. Kenchington [Gla91b]. Kingdom [Hat97, MG97, Smi99a]. Kingston [JD92, Kim90, KW91]. Kitakyushu [Sue94]. Kluwer [Gos93b, Bro93a]. knowledge [SL99]. Korea [Hon95]. Korean [Lee90]. kowska [Bro93b]. Kroner [Gau94]. Kwiat [Bro93b]. Kwiat-kowska [Bro93b]. Kwiatkowska [Sco90].

L [JD92, Lon91]. Laboratory [Ano93a]. Lamson [JD92]. Land [Boy92a, Dah93, oUotCEP92, SHG96, WT94, GP99a, GP99b, Küt94, Oki93, Wet94, Wil96]. Land-based [Boy92a, Dah93, oUotCEP92, SHG96, WT94, GP99a, Küt94, Wet94, Wil96]. lanes [Rot90]. Large [Ale93, Bis95, Eik93, Lon91]. Latin [Gab90]. Latvian [Ber90a]. Law [Bir92, Bir94b, Bro21, Gla90b, Gla91a, Ijl90, Kim90, KW91, O’K96, Pre90, RK94, AO93, Ago91, Bir94a, Boc91, Boy92b, Bro93b, Bro93e, Bro96b, Bro97b, But90b, But92, Dev93, Dev94, Gau95, Gib93, Gol91, Jan92, Joy96, Key99, Kim93a, Kit90, Kol93, Kwi90a, Kwi93b, Lan90a, Low90, Red90d, Sco90, hS99, Tre90, Ver97, Wag95, Won96, Boy90, Bro95a, Gla90b, Hay96, Wec90, Gla90a, Gla91c, Van92]. Lawson [Nau94]. lawyer [Pla95]. Laytime [Lax92]. lead [LV95]. leadership [Sue94]. Leanza [Gra90]. learned [HB98, NSS96]. learning [Hem98]. legacy [VH97]. Legal [KAP96, Mah97, Red91c, Bir94a, Boy92a, But90b, DV90a, Gla90b, Gla91c, GY21, Gra90, KV90, Kwi93a, Pic96, Pla90, Red90e, Don93, Sor92, Bir94b].

Legislation [Sta94, Git92, KK91, Mak92]. Legislative [Pic94a]. less [Dur94]. less-developed [Dur94]. lesson [NSS96]. Lessons [Bai97, JM95, Jon99, HB98]. Letter [Kim94]. level [Fen94, Fre90a]. liability [Gau94, Gau95, Red90e, RG90, Wal93]. light [Sid97]. Limitation [Gau95]. Limited [Red91c]. Limiting [Bar90a]. Limits [Gez99]. line [MU92, Mar92]. liner [Dav90a, Fra93, Stu90]. Lisbon [Ano94h, Ano94h, God94]. litter [Hal93]. liv [Kwi93a]. living [Cad96, Gul90, Hof93]. Lloyd [Sha92]. Loayza [App94a]. lobster [tCCT91, CP99]. Local [OZ94, Ram90, Ano94h, BW94, Fel90b, Fen94, GN90, Jak99, Kba94, Mac93a, MP97, WT94]. London [Bro93b, Bro93c, Clé96, Gos93a, Gra91, Kwi93a, Law90a, Low90, Red90e, Red91c, Tan90, Til91, Wat90a, Bar90b, Pic94a]. long [Ano91d, Ano91p]. longline [CM97]. longlining [Ber97]. LOSC [Eva96]. Louis [Bla91, Low90]. Ltd [Bir94b]. Lucian [App94a]. Luigi [Gra90]. Lusitania [Rot92]. Lysaker [BC92].

M [Ano93a, App94a, Bir92, Lon91, Sut90, Bro21]. M/V [Bro21]. Macmillan [Wat90a]. made [Clé96]. Madrid [Joy96]. Mahmoudi [Tre90]. Main [Ijl90, Gla91a]. Maine [CVM95, Ken97, Her95]. major [Ano96i, KV90]. making [Hue95, Nie94, PW92]. Malaysian [Moh91, ONAK92, Sah95]. Malcolm [Pre90]. Malvinas [Bis93]. mammals [Bir93]. Management [CM91, Hun97, Law90b, Pot99, Sch98a, Sco95, Sme90, Sut93b, Ago91, Ale93, Arn93, BW94, BS99, BH94, Bis93, Bla98, Bro93c, BP99, CP99, Cor95, Cre99, Dah93, Day95a, Dij93, Dio98, Dou93, DvV96, EHP93, Eik93, Ekp98, Fel90b, Fel90a, Fog95, FHG97, FGA93, Fra93, Gib93, Goo93, GRBC98, Gos93a, Gul90, HPB92, Hat97, HH98, HG96, Hol93, Høn98, Ist93, JM95, JMW21, JB90, Kap99, Kap21, Kar96, Kes95, Kes97a, Kes97b, KG21, Kit90, KA94, KT93, KA91, KAP96, Law93a, Lod92, Lod98, Mac93b, Mit97, Moh91, Mor95, NSS96, Nie94, NV97, Nij90, Owe99, Par93, PM96, PW92, Pic94b, Pol90, Pom91, PP95, PC97, PB97, Pot94, Rob93, Sah95, SN96, SN97, Smi95a, SP96, Smi99a, Ste93a, SE97b, SP97c, Thé97, Tow94]. management [Tri90, Val95, WyZ92, WAMK94, WAK96, WT94, WC99, WMH99, WD91, Yu91, dVdLE97, oA93, Bal99b, Hil97, Hin92, Sau98, ZC90, Bla93, Dub91a]. Managing [HS96, Rei21, Smi93, Sto94, Ber97, Gla91b, Sut90]. Mangone [Tan90]. Mann [Wat90c, Wat90d]. Manual [Gos93a]. manuscripts [Ano93b, Ano94b]. mapping [Ver97]. maps [Bla93]. March [Ano90l, Ano91o, Ano92g, Ano92l, Ano93j, Ano93t, Ano95g, Ano96n, Ano97i, Ano98i, Ano99d, Bar90b, Kim90, KW91, McE91b, Pic94a]. Marine [Ano90f, Bár94a, CK93a, Cox93, Far96, GY21, HP98, Hon95, Jon99, Jus98, Kri91, Nau94, SA98, Sos94, Sun94, WyZ92, YR94, Yu91, Ale93, Ano93a, Ano94h, Bat21, Bir93, Bis95, Boy92a, Boy92b, Bud99, Byw95, CVM95, Dav90c, Dev94, DN98, Don93, Eik93, Eva96, Gla91b, Hin92, HKP91, Hof93, Joy96, Ket95, Key99, Kin94, KG21, KK91, Kor92, Kul98, Kwi90b, Kwi90c, LS99, Lon91, Mah97, McE91c, McE92, Mur94, Nol91, Nol93, Per98, RL92, Sah95, Sch98a, SHG96, Smi93, Smi94, SL99, SS98, Squ94, Sta93, Ste93b, SP97c, Thé97, Val90, Wat90b, Whi90, Wil96, WT94, ZC91, dVF94, Bar90b, Dev93, dVF94, Bro93c, Tan90]. Maritime [And90, Bir94a, Mur94, Sri90, Bla91, Blo94, Böh92, Bri94, Bro93a, But92, Chu93, Clé96, CD94, Cri95, GPT96, Gau95, Gol90, Gos93b, Her95, Ket95, LV95, Mak92, Mar93, Pre90, RH96, Sch96, Til91]. Mark [Smi94, McC99]. Market [DvV96, Dav90a, GP99a, GP99b]. marketing [McE91c]. Markets [BGS92, Böh92]. Martinus [Bir92, Bla91, Bro93b, Kwi93a, Low90, Sco90, Van92, Lan90a]. Masters [Noe93]. material [GN90, vdM94]. materials [Pic96]. May [Ano90m, Ano91j, Ano92g, Ano92h, Ano93j, Ano93p, Ano94n, Ano95d, Ano96k, Ano97j, Ano98g, Ano99e, Sto91a]. McGill [JD92]. McGill-Queen [JD92]. McLay [Ano93a]. McRae [Gol91]. McVicar [Ano93a]. meaningful [Wat90b]. means [GSG93]. measures [HPB92, Hug92, Lod92, Sah95]. Mediterranean [Gra90, Ant92, Küt94, Rid97, Sco95, Val95, Gla90b]. Medunarodno [Gla90b]. meeting [Ano90g, Phi90, Wec90]. member [App94b, Law93c]. membership [YSM96]. Merchant [AO91, Ber90b, Git92, MG97, NR99, Pla96]. Message [dC90]. Miami [Fre90a]. Michael [Wat90a]. Migrations [Haa92]. migratory [Bar95]. Miguel [PP95]. mile [Sco90]. Miles [Dub91a]. Military [Van91, Bar90a, Roa96]. minerals [Kwi93a]. Minimum [Lod92]. mining [Bro93e, Nol91, Tre90]. Minke [BC98]. Miquelon [Mar93]. missile [hS99]. mission [Mur94]. mitigating [GSG93]. mix [BE99]. mixed [KT93, RB99]. MLA [Ano93a]. models [Gul90, Sum99, dVdLE97]. modern [McE91a]. modern-day [McE91a]. module [HP98]. monitoring [Ano93a, McI95, Rit93, Sau98, WMH99]. Monographs [Pre90]. Montreal [JD92]. Morgan [Smi94, Wat90d]. Morris [Wat90a]. most [Hug94]. movements [Sme90]. MS [Ano96a]. MS/Estonia [Ano96a]. MSY [Kes97b]. Muddling [Aka95]. multi [Dev94, Gul90, Le 92]. Multi-disciplinary [Le 92]. multi-party [Dev94]. multi-species [Gul90]. multispecies [Eik93, SCC^{+}98]. Munro [Gol91]. museum [Clé96]. mutually [Bar92]. Myron [Low90].

N [Hem98]. NAFO [Day95b, Gez99]. Nairobi [Gra91]. Nansen [BC92, Sto91a]. narcotic [Gil91]. Narcotics [Gil96]. nascent [CVM95]. Natalino [Lan90a]. National [Gol91, Moh91, Clé96, Lév93, Stu94, Nie94]. Nations [Bar95, Bir92, Bro96b, Gla91a, Gra91, Hay96, Low90, Ver97, Gil91, Won96]. natural [KA91, PW92, Sch98a]. nature [Gla91c, Jon99, QJZ97]. naval [Bar90a, Bat21, Jon97, Lan90a, MP97, Rid97, SS97, SP97b, vE94]. navies [Wat90a]. Navigation [Cor95, Pla96]. Navy [Mur94, SS97]. navyports [Bro97a]. necessary [Bro93e]. need [Dav90c, Ken92, Pol90, Tho94, Law93a]. needed [Nau94]. negotiation [Cad96]. negotiations [Dev94]. Neither [Bro93e]. net [Ano91r, WD91]. Netherlands [Bir92, Bla93, Gos93b, Wec90]. network [HP98]. Nijhoff [Bir92, Bla91, Bro93b, Kwi93a, Lan90a, Low90, Low94, Sco90, Van92]. No [App94a, Nau94, Til91]. nodules [Zhi91]. non [Pot99]. non-statutory [Pot99]. Noordwijk [Wec90]. nor [Bro93e]. Nordquist [Bir92, Low90]. norms [RK94]. North [Sme92, AJ91, Bro92, FS98, Ist93, LV95, MH92, Sid97, Skj92, SA98, Str94, Wet94, Bla93]. North-East [Wet94, LV95]. Northeast [NSS96]. northwestern [Høn98]. Norton [Wat90c, Wat90d]. Norway [BC92, AJ91]. Norwegian [BC98, Jak99]. nose [Day95b]. note [Gar90, Nic90]. Notes [Ano90j, Ano90k, Ano91h, Ano91i, Ano93k, Ano93l, Ano93m, Ano93n, Ano93o, Ano94i, Ano94j, Ano94k, Ano94l, Ano94m]. notions [Jan92]. Nova [AJ91, Bin91, HPB92]. novel [CVM95]. November [Ano90q, Ano91n, Ano92g, Ano92k, Ano93j, Ano96m, Ano97l, Ano98h, Ano99g, EHP93, Fre90a, Red90e, Red91c, Now93]. noxious [Gol90]. NSW [SP96]. Nuclear [VH97, DV90a, Mor99]. number [Gla90b]. NY [Gra90].

objective [Cad96]. objects [Hay96, Hut96]. observations [Kul98]. observing [SM96, Woo94]. Ocean [AB97, Arn93, Böh92, Bro93c, Kwi93b, Mai98, AO93, Ago91, Aka95, CB91, Cri95, Day95a, Ekp98, Gla91a, HB98, Ijl90, JB90, Kit90, Kul98, Lau94, Lév93, PW92, SM96, VH97, Woo94, Bro97a, Kwi90c, MH92, Pre91, Wat90c, Wat90d]. Oceana [Gra90]. oceanic [Bla91]. Oceanographic [Oke98, Pon91]. oceans [BKLM98, BM96, Bla98, Bur93, Gol91, HBMV98, Kim93b, McI95, Mit21, HP98]. October [Ano90g, Gul90, Jis95, McE93a, Pon91]. OECD [App94b, Law93c, Pot94, App94b]. Office [Gla91a, Ste93b]. Officer [Hey92]. Offshore [Hug94, Bin91, Eva96, Pic99, SBH93]. Oil [Ket95, Gar93, Hol93, Hug94, Owe99, Pic99, Red92, Rit93, Sid97, Sta93, Wal93, Gar93, Red90e, Red91c, Ano93a]. oil-spill [Ano93a]. OMAF [HP98]. one [Dic92, Sco94, YSM96]. Open [Sha92, Egi90]. operations [Aka93, GSG93, MU92]. operative [Fel90b]. Opportunities [Aka93]. Optimal [Mor95, KT93, NSS96]. options [BW94, Fre90a, Hil97, Van96]. Organisation [Sri90]. organisational [Bal99b]. organisations [Bro93b, NV97]. Organization [And90]. organizational [Tho94]. organizations [Hat97]. organized [Red90e, Red91c]. origins [Sch95]. Orrego [Gla91c]. Oslo [Sto91a, Hey91]. other [Fre90a]. Our [Gra91]. Overcapitalization [GW91]. overlapping [SE97b]. overview [Per98]. owners’ [Cam96]. Oxford [Kim93a, Pre90, Smi94].

P [Gau94, McE91c]. Pacific [Ano91r, HRS97, Hey92, Law90a, Law90b, Lod98, Mai98, McE91c, MH92, Ano91d, Ano91p, Ano91s, AB97, BM96, DT90, Dou93, FGS95, Haw91, Hun97, Lod92, Mar91, Miz91, Rim97, Tab94, WD91, McE91c]. Packaging [WBN91]. Pages [Ano90n, Ano90o, Ano90l, Ano90m, Ano90q, Ano90p, Ano91k, Ano91l, Ano91o, Ano91j, Ano91n, Ano91m, Ano92i, Ano92l, Ano92h, Ano92k, Ano92j, Ano93r, Ano93q, Ano93t, Ano93p, Ano93s, Ano94p, Ano94o, Ano94n, Ano94q, Ano95e, Ano95g, Ano95d, Ano95f, Ano96j, Ano96n, Ano96k, Ano96m, Ano96l, Ano97h, Ano97i, Ano97j, Ano97l, Ano97k, Ano98f, Ano98i, Ano98g, Ano98h, Ano99c, Ano99f, Ano99d, Ano99e, Ano99g, Ano21b]. Palau [AB97, Lod98]. Pan [Ekp98]. Pan-African [Ekp98]. Panikkar [Hem98]. Papers [App94a]. paradigms [HG96]. paradox [HH98]. parametric [FHG97, WAK96]. Paris [Law93c, Hey91]. Park [Kri91]. Parliament [Col90]. Part [Bro95a, Kwi90b, Kwi90c]. Participation [Nie94, Don93, Fel90a, JM95, Mar91]. particular [BB96, Gla90b]. parties [Won96]. party [Dev94]. pass [McC99]. Passage [Mah91, Fra90a, HM99, Lan90b, Rot90, Rot92]. Passenger [Mar92]. past [HM99, Jon99]. Patagonian [Bis93]. path [Law93a]. Patterns [Lon91]. peacetime [Pla96]. pelagic [McE91b, Ano91r]. Penataran [Kwi90a]. penetration [lW90]. Penning [Gos93a]. Penning-Rowsell [Gos93a]. people [Sch96]. Perceptions [PPJ95, Mar92]. perestroika [Git92, But92]. Performance [KG21, Dav90a, MU92]. Persian [Mah97]. perspective [Bar92, CM91, Hal93, HKP91, Kul98, Pol90, Red91c, YSM96]. perspectives [DvV96, RH96, NV97]. Peter [Goo97]. Peterson [Sut90]. Pharand [Gla90a]. phase [Jus98, Tan98]. phases [Jus98]. Philippines [PP95, PC97]. Philipson [McE91c]. Pierre [Mar93]. Pierre-Miquelon [Mar93]. Pinkerton [Fel90b]. Piper [Pic94a]. Plan [CVM95]. planning [EHP93, Gib93, Tow94]. plans [AM98, Pot99]. points [NSS96]. Polar [Hue95, YCS95, RK94]. pole [MU92]. pole-and-line [MU92]. policies [CS97, Sko94, hBS90]. Policy [Gar90, Law93b, Nau94, Sun95, Ago91, Aka95, Bir94a, Bla93, Blo94, Bro93c, CP99, CK93a, Dav90c, Day95a, Dur94, Eva96, Fre90a, GPT96, GP99a, Gez99, Gla91a, Gol91, Gwi93, Hon95, Hue95, Ijl90, Kin99, Kit90, LRA96, Lee90, Lév93, Nie94, ONAK92, Smi94, Tan90, Tri92, VH97, Wat90b, WMH99, dVF94, lW90, Sha97, Tol92, AO93]. politics [Goo97, Pla96, Sto91a, Wet94]. Pollution [Gar93, Ket95, Ald92, Boy92a, Byw95, Dah93, Dav90c, Dun93, Hal93, Ist93, Joy96, Kin95, KK91, Kor92, Küt94, Lei93, Mac93a, McI93, Mut94, Nol93, oUotCEP92, Red90e, Red91c, Red92, SHG96, Sta93, Wal93, Wet94, Won21, WT94, ZC91]. polymetallic [Zhi91]. pool [Bav96]. Population [ZC91]. Poredbeno [Gla90b]. Port [Bal99a, Bel93, Dij93, Hol93, Par93, Jon97, Ken92, Low94, Pay94b, SS97, SP97b, vdM94, Low94]. ports [Bal99a, MP97, WMH99]. position [Bel93, Goo93, Mac93b, Rob93]. possible [BB96, Val90]. Post [Tan98, Aka95, Cor95, GPT96, Smi95b]. post-Donaldson [Cor95, Smi95b]. Post-Harvest [Tan98]. post-war [Aka95, GPT96]. Postharvest [Nau95]. potent [MA91]. potential [Sto94, HS96]. power [Pic96, Rid97]. pp [AO93, App94b, Bir94b, Bla91, Fel90b, Gau94, Gla90a, Gla91a, Gla91b, Gol91, Goo97, Gra91, Hey92, Ijl90, Lan90a, Law90a, Low94, Nau94, Sco90, Tan90, Til91, Wat90a]. pp. [App94a, Gla90b, Gla91c, JD92, Lon91, McE91c, Smi94, Sut90]. PR [hBS90, WyZ92, Yu91, ZC90]. Practical [Pic94b, Lan90b]. practice [Eyt96, FGA93, Hol93, Kim93a, Red91b, Wal93]. practices [Dou93]. Pragmatic [Stu90]. Prava [Gla90b]. Pravo [Gla90b]. precautionary [Hey91]. predicting [Kul94]. Preface [SP97a]. Preparatory [Kim90, KW91]. prepare [Boi94]. prescription [Sun96]. present [GP99b, HM99]. Presentation [Clé96, Dun93, Lar93, Par93, oA93]. preservation [Ken97]. President [oCKG90]. Presidents [Ano90g]. Press [Dub91a, Dub91b, Fel90b, Gla90a, Gla91c, Gol91, Goo97, Gos93a, Kim93a, Pre90, Smi94, Sut90, Wat90c, Wat90d, JD92]. prevention [Kin95, Nol93, KK91]. price [Gar90, Tha21]. principal [Cou96]. Principles [Mat92, Bro96b]. Prinosi [Gla90b]. private [DMV99, SE97a]. Privatization [Eyt96]. problem [CS94, Lau94, Skj92]. Problems [Hal93, Aka93, Boy92b, Fen94, Hin92, Hut96, Key99, Kwi90b, Miz91, Nie92, Dic92]. process [Kim93b]. processes [Lon91]. processors [AJ91]. Producers [Hat97]. production [Bar90a, Cou99]. products [McE91c, Pic96]. Professor [Lan90a]. Programme [Gra91, Jus98, Mai98, Sau98, Ano93a, Dio98, Ekp98, Wil96, AM98, Nau95, SE97a]. programs [HGG99, PC97]. progress [CS96, Key99]. prohibition [Ano91d, Ano91p]. Project [Tan98, Jus98, McE91b, Sau98, Tan98]. projects [PC97]. Promoting [Dav90a, Fra90b]. prompt [Bro21]. Property [DMV99, GY21]. Prospects [Cam96, Kol93, Aka95, And94, JB90, Key99, Kwi90b, SHG96, Hey92]. protect [Day95b, Hug92]. Protected [dVF94, Far96, Nor94]. Protecting [Boy92b, O’K96, Van90, Wil96, Ano90g]. Protection [Bar90b, Bro96b, Ket95, BB96, CVM95, Clé96, Don93, Dur94, GRBC98, Joy96, Kba94, Key99, Mah97, Nol91, Pla96, RL92, WNJ99, AO91]. protections [Wag95]. protocol [Joy96, oUotCEP92, SHG96]. Protocols [Ano91p]. Proucavanje [Gla90b]. provided [Wag95]. provident [Dub91b]. prudent [Bro93e]. PSSAs [BB96]. psychotropic [Gil91]. Pubblicazione [Pre91]. Public [Day95a, Cox93, Ste93a]. Publications [Ano90r, Ano90s, Ano91q, Ano92m, Gra91, Gra90]. Publishers [Gla91a, Ijl90]. Purse [Lod98, AB97, HRS97, McE91a, Now92]. pursuit [Kor90].

quality [Fra93, God94, Sko94, Str94]. Quantifying [Red92]. Queen [JD92]. quest [Lév93]. question [Val94]. questions [hS99]. quota [BS99, Hat97, Kar96, Mor95, PW92]. quotas [Cla93, FGA93, SCC^{+}98, Tow92].

R [Ano93a, Gau94, Wat90d]. Radioactive [CB91]. raises [McE92]. rate [Haw91]. rational [Ano91s, KA91]. realistic [Kes97b]. reappraisal [MP97]. reception [Bal99a]. Reconsidering [RK94]. recovery [Rit93]. Recreational [WNJ99, KG21, SP96]. Reduction [Lal90]. Reef [Kri91]. reefs [Pic96]. Reference [NSS96, Per98, Sco95]. reform [Mar97, Sch98b]. regarding [Ano96a, GGIMSRVL96, GRBC98]. Regime [Gla91c, Boy92a, Bro93e, Bro97b, Gla90b, KV90, Kwi93a, Pla96, Van92, Wet94, Gra90, Low94]. regimes [LV95, Sco95]. region [Lod92, Miz91]. Regional [AM98, Nau95, BKLM98, DT90, Dur94, GN90, HB98, HBMV98, Oke98, SHG96, Smi99a, SM96]. regionalisation [Høn98]. regionalization [dVdLE97]. regions [Dur94, Kwi90b, Kwi90c, RK94]. register [DT90]. registration [Git92, Egi90]. regulated [Sha97]. regulating [DvV96]. Regulation [Kap21, MT96, DvV96, GGIMSRVL96, Mor99, Nie94, PW92, Sta94, dM90]. regulations [Bav96, PPJ95, SE97b]. regulatory [Won96]. related [SP96]. Relations [Gla90b, But90b, Fra90a, Her95, Kri91, Mak92]. relationship [Nol93]. release [Bro21]. relevant [Gul90, Val94, hS99, Pre90]. rent [Mat92]. report [Ano93a, BC98, McE93b, BKLM98, Pla95]. Representing [YSM96]. Republic [Hon95, Sun95]. requirements [Won96]. Research [Ano93a, BGW91, Law93b, McE91b, GY21, Hem98, Kul98, Nau94, Law93b]. reserves [Jon99]. Resolution [Ano91r, Ano91u, Ano91f]. resolve [McE92]. resolved [Dev93]. Resource [Sau98, Arm98, BC92, Cam96, Hug94, KA91, Nau94, PC97, Val90, WyZ92]. resources [Ale93, Ano91s, Ano96f, Bat21, Bav96, Bla91, Cad96, Eva96, FGS95, Gul90, Hof93, Ket95, KA91, Sch98a, SS97]. response [BE99, Byw95, Nau94, Pla95]. responses [SP96, Val90]. responsibility [Sco95, Gar93, Kor92]. Responsible [Lod92]. restoration [Dou93]. revenue [RB99]. Review [AO93, Ano93a, App94b, App94a, Bir92, Bir94b, Bla91, Bla93, BC92, Bro93b, Bro93a, Bro93c, Clé96, Dub91a, Dub91b, Fel90b, Gau94, Gla90a, Gla90b, Gla91a, Gla91c, Gla91b, Gla92, Gol91, Goo97, Gos93b, Gos93a, Gra91, Gra90, Hey92, Ijl90, JD92, Kim93a, Kwi93a, Lan90a, Law90a, Law93c, Law93b, Lon91, Low90, Low94, McE91c, Pre91, Pre90, Sco90, Smi94, Sut90, Tan90, Til91, Tre90, Van92, Wat90a, Wat90c, Wat90d, Dav90b, GPT96, Nau94, Pic96, PC97, Sum99, Law93c]. revision [Hey91]. revisited [Kes97b]. Richard [Gla91b, Hey92]. right [Kor90]. rights [DMV99, Eyt96, KA91, Sha97, Tow95b, Bla91]. Rim [Haw91]. Rio [CS96]. rise [Fre90a]. Risk [Jis95, Nij90]. risks [PW92]. Robin [Bro93c]. rock [CP99]. Role [Sko94, And90, CB91, Cha92a, Dij93, GN90, Gol90, Hol93, Lee90, Noe93, Per98, PB97, Rei21, SE97a, Sri90, Bár94a, BW94, Dur94, Gib93, WT94, dVF94]. Ronzitti [Lan90a]. Rosenne [Low90]. Rotterdam [vdM94]. routes [Owe99]. Routledge [Bro93c, Law90a]. Row [Gla92]. Rowena [Nau94]. Rowsell [Gos93a]. Royal [Mar92, Pic94a]. rule [Arm98]. rules [Töp90]. Russia [Høn98]. Russian [Aka93, VH97].

S. [Dev94]. safe [Bar92]. safety [Bin91, Boi94, PPJ95, Smi95b, Pla96]. Said [Tre90]. ‘Saiga’ [Bro21]. Salerno [Nei92]. salmon [BGS92, CM91, Fel90a]. Salvage [Bri96, Red90d, Sha92]. San [Phi90, PP95]. Sarasota [Gla91a, Ijl90]. scale [AJ91, KA94, Pom91]. scenarios [Rid97]. Science [JD92, Jus98, Lon91, Wet94, Ste93b, YR94, Tri92]. Scientific [Cor92, GY21, Pic96, WMH99]. Scienze [Pre91]. Scotia [AJ91, Coo94, HPB92, Now92]. Scotia-Fundy [Coo94]. Scotian [Bin91]. Scotland [Cou99]. Scottish [Hal93, Nei91]. SE [McE92]. Sea [Bir92, Bla93, Gla91a, Goo93, Kim90, KW91, Mac93b, Rob93, SE97a, SP97c, Wec90, oA93, Bar92, Bir94b, BB96, Boc91, Boi94, Boy92b, Bro93e, Bro96b, Bro97b, But90b, DV90a, Dun93, Fre90a, Fre90b, Gil91, Gil96, Hut96, Jan92, Kol93, Kwi90a, Kwi93b, Low90, McI93, Nol91, Nol93, O’R95, Rot90, RK94, SBH93, SL99, Sue94, Tre90, Ver97, Wag95, Won96, Won21, Boy90, Bro95a, Bro21, Eik93, Hay96, HS96, Ist93, Kar90, McE91a, McE91b, Sid97, Skj92, Sme92, Sto94, Str94, Sun95, Val90, Val95, Van96, Gla92, Bro93c, Gol91, Bro93b, Dub91b, Kim93a, Sco90, Kwi93a]. sea-based [Nol93, Won21]. sea-bed [Nol91, Tre90, Kwi93a]. sea-fishing [Bir94b]. seabed [Bro93e, Zhi91]. seabird [Ber97]. seabirds [Hil97]. Seafarers [GP99b, GP99a, NR99]. seafloor [Ver97]. Seaport [Ken97]. seaports [Tho94]. Search [ONAK92]. Sears [Ken97]. Seas [AM98, Jis95, Bar92, Gla90b, Gra91, Hal93, Han95, Smi99b, Smi99a, Ste93a, Sun96, Van91, Smi94]. seasonal [Sch98b]. seasonality [tCCT91]. Seattle [Dub91a]. Seawatch [SM96]. second [Clé96, Hug94, Tan90]. secretary [dC90]. secretary-general [dC90]. sector [Ken92]. security [Cri95, MH92, Stu94, Til91, Til91]. Seine [Lod98, AB97, HRS97, McE91a, Now92]. self [Tow95a]. self-governance [Tow95a]. semi [Gla90b]. semi-enclosed [Gla90b]. Seminar [Haa92, McE91b]. sensitive [BB96]. September [Ano90p, Ano91m, Ano92g, Ano92j, Ano93j, Ano93s, Ano94q, Ano95f, Ano96l, Ano97k, Ano21b, Kim90, Law90b]. Series [App94a, Law93b, Bro93c]. Services [Red91c, Pla90, Red90e]. Session [Kim90, KW91, Bar90b]. settlement [Bro97b]. settlers [Kap99]. Seventh [Kim90]. Shabtai [Low90]. shared [Cad96]. Sharing [Arm98]. shelf [Bis93, Bis95, Gra90, Kwi93a]. Sherman [Lon91]. Shetland [Goo93, Hal93, Smi93, Sut93b]. Shifting [GPT96]. shipbuilding [Kin99, hBS90]. shipments [dM90]. Shipping [AO91, Smi95b, hBS90, Ber90b, Böh92, Dav90a, Dij93, Fra90b, Fra93, Hol93, JD92, Lee90, Par93, Smi95a, Stu90, Sut93b, Tha21]. Ships [Dev96, Bar92, Byw95, Egi90, Git92, KK91, Pla96]. shipwrecks [Kin95]. shoreline [Kin94, Pot99]. Sico [Gra90]. since [Bro97a, CS96]. Singapore [McE93a]. sites [Hut96]. situation [GP99b]. Skipjack [MU92, Pol90]. Slipi [McE91b]. Small [KA94, Pom91, AJ91]. Small-scale [KA94, Pom91, AJ91]. Social [JMW21, Law90a, MS93]. socially [KT93]. societies [Boi94]. society [MA91, Pic94a]. socio [Nie92]. socio-economic [Nie92]. softback [Gol91]. Sohn [Low90]. soil [Oki93]. Solomon [Hey92, Law90b]. solution [CS94]. Solving [Fen94, Skj92]. Some [Won96, Ano96i, SP96, Ber90b, Boy92b, Jan92, Lan90b, Now92]. Soons [Bro93b]. source [Ist93, Lal90]. Sources [Squ94, Boc91, Boy92a, Küt94]. South [Bis95, Hey92, Law90b, Mai98, McE91c, Sut90, Ano91d, Ano91p, Ano91r, Ano91s, BM96, CP99, Dev93, Dev96, DT90, Dou93, HS96, HP98, Hun97, Lod92, Miz91, Sos94, Sto94, Sun95, Tab94, WD91]. South-east [Sos94]. Southeast [Bis95]. Southern [BH94, PPJ95]. Southwest [WC99]. Sovereign [KA91]. sovereignty [Rot92, Val94]. Soviet [Ber90b, But90a, But90b, But92, Fra90a, Git92, KV90, KK91, Mak92]. spaces [Nor94]. Spain [BGS92, dVdLE97]. Spanish [GGIMSRVL96, dVF94]. Spar [Sid97]. special [Sco95]. species [Gul90, WBN91]. spill [Rit93, Ano93a]. SPODP [Mai98]. Sprague [App94a]. Spratly [Sun96, Val94]. spurious [Nol93]. St [Mar93]. stabilised [Pic96]. Stadsvitgeverij [Bla93]. stage [Bro93e]. standard [KPG95]. standards [Tri90]. state [Bel93, DMV99, Low94, Pay94b, WT94, Kwi90b, Kwi90c, Nad92, Rot92]. Statement [Ano91s, Che94]. states [Won96, HRS97]. stations [Pic96]. statistics [Pon91]. status [Jon99, YR94, And94]. statutory [Pet99, Pot99]. STCW [McC99]. Steinar [BC92]. Stewardship [Nor94]. still [Val94]. stock [Mit97, NSS96, Tow95b, WLD98]. Stockholm [Tre90]. stocks [Bar95]. Stoett [Goo97]. straddling [Bar95, Mit97]. Strait [Mah91, Sch96, hS99]. straits [Pla96, Rot90, Sco95, Sco95]. strategic [Cri95, MP97]. Strategies [Law90b, Cre99, Gol91, Pic94b, RL92]. Strategy [Ano90f, App94a, Bla98, Rit93, Sto91b, Til91]. Structural [Nie92, GW91]. structure [Sun94]. structures [Tow95a]. Studi [Pre91]. Studies [Hey92, McE91c, Red91c, Cre99, Red90e, SP96, Til91]. Study [Gla90b, Nau94, DMV99, HPB92, Law93b, Mai98, Mar92, WNJ99, WBN91, WC99]. Subic [Rim97]. Submarine [Wag95]. submarines [DV90a]. submission [Ano93b, Ano94b]. substances [Ano90g, Gil91, Van90]. successfully [Aka95]. suitability [Pic96]. Summary [Ros90]. summit [CS96]. Sunken [Roa96]. supply [tCCT91]. Support [Dio98]. survey [HGG99]. Suspicion [Kap99]. Sustainability [Sen96]. Sustainable [Mit21, Nij90, Bla98, KA91, Val95]. Symposium [Gul90]. system [BS99, CD94, God94, Kar96, Mor95, SS97, SM96, Woo94, Sto91a, Sut90]. systems [Edw94, Hat97, KPG95, LS99, Nau94, PW92, SP97b].

tables [Law93c]. TAC [Kar96]. Tagus [God94]. tail [Day95b]. Taiwan [hS99]. take [LV95]. tango [PB97]. tanker [GSG93, Owe99]. Tarawa [Ano91t]. Taylor [Gla91b, Tan90]. technical [Cor92, DV90a, Sah95]. Techniques [Gos93a]. Technological [Whi90, KPG95]. technologies [YCS95]. Technology [Tan98, Gab92, Ste93b, Ver97, YR94]. ten [BS99]. Tending [Day95b]. tensions [McE92]. Tenth [Law90b]. terminal [GSG93]. terms [Lod92, Tha21]. terrestrial [Cox93]. tests [hS99]. Thailand [Kit90]. their [Hun99, BB96]. theoretic [Sum99]. theoretical [Arm98]. Theory [Eyt96, FGA93, GRBC98, JMW21, Wal93]. there [Won96]. Third [Mar90, McE93a, Wat90a, dV90b]. threat [MH92]. Threats [Hut96]. three [DMV99, Sco94]. times [Gla92]. Timor [Sut93a]. Topping [Ano93a]. Toronto [AO93]. Torres [Sch96]. Torry [Ano93a]. Total [Fra93, Kar96]. tourism [MA91]. toxic [Gab90]. TQ [Mor95]. trade [Mar91, Pay94a]. traders [Dic92]. trades [MT96]. trading [AO91, CD94]. traditional [Dou93, RK94]. traffic [Gil91, Pla90]. trafficking [Gil91]. tragedy [McE91a]. Trainee [BKLM98]. training [Ken92]. Transboundary [Thé97, CVM95]. Transfer [Gab92]. Transferable [Tow95b, Cla93, Mor95, Sha97, SCC^{+}98, Tow92]. transfrontier [Sme90]. transit [Lan90b, Rot90]. Transition [Dev93]. Transnational [Sto91b, Val90, Gau94]. transport [Blo94, Lon90, Mak92, Sun94]. trawler [Bai97]. treatment [God94]. Treaty [Sto91a, Sut90, Sut93a]. trends [BH92, WLD98]. triangle [Bat21]. Tribunal [Bro21, Jan92]. Trinity [Bir94b]. Tropical [Kul94, Sun96]. Trotman [Bro93b, Gau94]. Trotman/Martinus [Bro93b]. tuna [BH92, BH94, Cam94, Cam96, CM97, HRS97, Hun97, Pay94a, Pol90]. tunnel [Hue95]. Turkish [Pla96]. Two [PB97, Cre99, Sco94, DvV96].

UBC [Goo97]. UK [Dub91b, Gau94, Gla91c, Pre90, AO91, Bal99a, Bel93, Bir94a, Bri94, CD94, EHP93, GP99b, GW91, Hug94, LS99, Mor99, Now93, Owe99, Pic94a, Ram90, Red91b, Rob93, SL99, SE97b, Sun94, SP97c, lW90]. UKCS [Pic99]. Ulfstein [Bro93c]. Umberto [Gra90]. UNCED [Ano94h, Bár94a, BM96, Bir93, Bir94a, Bur93, Kim93b]. UNCLOS [And90]. Understanding [Kul94]. underutilized [WBN91]. underwater [BB96, Bri96, Bro96b, Clé96, Cou96, Hut96, O’K96]. UNDP [Law93b, Nau94, Che94]. unemployment [Sch98b]. UNEP [AM98, Tol92]. UNESCO [Clé96, Oke98]. unified [Cha92b]. UNIFO [Gla91a, Ijl90]. Union [Kar96]. unique [Stu94]. unit [WLD98, BGW91]. United [Gla91a, Gra91, HRS97, Bar95, Bro96b, Gil91, Hat97, Hay96, MG97, Smi99a, Ver97, Won96, Bir92, Low90]. universality [Bro93e, Bro95a, Kol93]. universita [Pre91]. University [Dub91a, Dub91b, Gla90a, Gla90b, Gla91c, Gol91, Hey92, McE91c, Smi94, Sut90, Wat90c, Wat90d, JD92, Fel90b]. unpalatable [BE99]. unprotected [Hut96]. Update [Kas90]. US$159 [Van92]. US$186 [Kwi93a]. US$237 [Bro93a]. US$24 [Hey92]. US$49.50 [Gla90a]. US$65.00 [Gla92]. US$67.50/£ [Bla91]. US-Soviet [Fra90a]. US/Pacific [Mar91]. USA [Gla90a, Kim90, KW91, Red91c, Ken97, WT94]. User [JM95, Fel90a]. Using [Lau94, Hil97, Pic96]. utilisation [Ano91s]. utility [Jan92].

v [Bel93, Bro21]. Valencia [Smi94]. Vallega [Pre91]. valuable [Hug94]. values [JD92]. Vancouver [AO93, Fel90b, Gol91]. VanderZwaag [AO93, JD92]. variability [PW92]. ventures [Aka93, But90a]. versus [DvV96, Ken97]. vessel [HGG99, Pla90, PPJ95, Ist93]. Vessel-source [Ist93]. vessels [Bro21, DT90]. vi [Smi94]. viability [dVdLE97]. Vice [Ano90g]. Vicuña [Gla91c]. view [Ano91s, Pla95, dVdLE97, dV90b]. Views [Cur92, Sco92]. Viking [Mar92]. vision [Hue95]. Vol [Bir92, Kwi93a, Nau94, Bro93b, Bro93b]. Vols [Ijl90]. Volume [Gla90b, Gla91a, Low90, Ano90i, Ano91g, Ano92g, Ano94g, Ano95c, Ano96h, Ano97g, Ano93j]. volumes [Gra90]. Vorsey [Bla91]. voyage [Rot92]. Vukas [Gla90b].

W [Gos93b, McE91c]. WA [Dub91a]. Wales [Bal99b, Gib93, Pet99, Pot99]. war [Aka95, GPT96, VH97]. warfare [Lan90a]. warning [Van91]. warships [Fra90a, Mah91, Roa96]. Washington [App94a, Dub91a, Law93b, Lon91]. Waste [GN90, Keb90, Tri90, Bal99a, CB91, Gab90, Lal90, Lau94, Pic96, Sme90, dM90, dV90b]. wastes [Lon90, Tol90]. wastewater [God94]. Water [Mac93a, CM97, Sko94, Str94]. watercourses [Boy90]. waters [Gla90a, Hue95, Nor94, Owe99, Per98]. way [CK93b]. weak [PW92]. were [Fel90a]. West [Dio98, Nau95, SA98]. Western [Jon97, Rim97, AB97, HRS97, Lod98]. Wettestad [BC92]. WEU [vE94]. whale [BC98, KT93]. Whaling [Phi90, Goo97]. which [SP96, Won96]. whiting [FGS95]. Whittle [Ano93a]. Wicksell [Tre90]. wildlife [Oki93]. Will [LV95]. Winchester [EHP93]. wisely [Lau94]. wishes [Fel90a]. withdrawal [Rim97]. within [dM90]. WMO [DN98]. Woods [Pon91]. work [HGG99]. working [Sha92]. Workshop [BKLM98, Fre90a, Pon91, Dun93, HB98, HBMV98, Lar93, Par93, oA93]. World [App94a, Hin92, Law93b, Dub91a, NR99, Pay94a, Sto91a, Wat90a, Mar90, McE93b, Nau94, dV90b]. wrecks [Hut96].

xi [Gla91a, Bro95a]. xiii [Wat90a]. xv [Kim93a]. xvi [Hey92]. xviii [Sco90]. xxxiv [AO93].

year [Dic92]. yearbook [Wat90c, Wat90d]. years [BS99]. yellowfin [Cam94, Pol90]. yields [Lon91]. York [Bro93c, Gla91b, Gla92, Kim90, KW91, Tan90].

za [Gla90b]. Zagreb [Gla90b]. Zealand [BS99, Cla93, Sha97]. Zee [Wec90]. zone [Gib93, Gla91c, Kar90, Kor90, KAP96, MA91, Nad92, Pic94b, Pot94, Sco90, WT94, ZC90]. zones [QJZ97, Van91].

References

Aqorau:1997:OGW

[AB97] Transform Aqorau and Anthony Bergin. Ocean governance in the Western Pacific purse seine fishery — The Palau arrangement. Marine Policy, 21(2):173–186, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0308597X96000541.

Agoes:1991:ICR

[Ago91] Etty R. Agoes. Indonesia and the LOS convention: Recent developments in ocean law, policy and management. Marine Policy, 15(2):122–131, March 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190011Y.

Arbo:1997:GFI

[AH97] P. Arbo and B. Hersoug. The globalization of the fishing industry and the case of Finnmark. Marine Policy, 21(2):121– 142, March 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97832241.

Apostle:1991:NSN

[AJ91] Richard Apostle and Svein Jentoft. Nova Scotia and North Norway fisheries: the future of small-scale processors. Marine Policy, 15(2):100–110, March 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190009Z.

Akaha:1993:JRF

[Aka93] Tsuneo Akaha. Japanese-Russian fishery joint ventures and operations: Opportunities and problems. Marine Policy, 17(3): 199–212, May 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390077G.

Akaha:1995:MTS

[Aka95] Tsuneo Akaha. Muddling through successfully: Japan’s postwar ocean policy and future prospects. Marine Policy, 19(3): 171–183, May 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9400012H.

Alders:1992:AP

[Ald92] Hans Alders. (2) Atmospheric pollution. Marine Policy, 16 (1):59, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290061S.

Alexander:1993:LME

[Ale93] Lewis M. Alexander. Large marine ecosystems: a new focus for marine resources management. Marine Policy, 17(3): 186–198, May 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390076F.

Akiwumi:1998:URS

[AM98] P. Akiwumi and T. Melvasalo. UNEP’s Regional Seas Programme: approach, experience and future plans. Marine Policy, 22(3):229–234, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X98000098.

Andrianov:1990:RIM

[And90] V. I. Andrianov. The role of the International Maritime Organization in implementing the 1982 UNCLOS. Marine Policy, 14(2):120–124, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090097B.

Anderson:1994:CSP

[And94] D. H. Anderson. LOS convention: Status and prospects. Marine Policy, 18(6):494–497, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9490071X.

Anonymous:1990:Ca

[Ano90a] Anonymous. Calendar. Marine Policy, 14(1):ibc, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X90900502.

Anonymous:1990:Cb

[Ano90b] Anonymous. Calendar. Marine Policy, 14(2):180, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90901062.

Anonymous:1990:Cc

[Ano90c] Anonymous. Calendar. Marine Policy, 14(4):376, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900672.

Anonymous:1990:Cd

[Ano90d] Anonymous. Calendar. Marine Policy, 14(5):460, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090014I.

Anonymous:1990:Ce

[Ano90e] Anonymous. Calendar. Marine Policy, 14(6):542, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090029Q.

Anonymous:1990:CIM

[Ano90f] Anonymous. China Institute for Marine Development Strategy (CIMDS). Marine Policy, 14(5):455, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090009G.

Anonymous:1990:CMV

[Ano90g] Anonymous. Conclusions of a meeting of the Vice Presidents of ACOPS on 4 October 1989: Protecting the environment from hazardous substances. Marine Policy, 14(3): 295–296, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900937.

Anonymous:1990:EB

[Ano90h] Anonymous. Editorial Board. Marine Policy, 14(1):ifc, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090031L.

Anonymous:1990:IV

[Ano90i] Anonymous. Index to volume 14. Marine Policy, 14(6):539– 541, November 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090027O.

Anonymous:1990:NAa

[Ano90j] Anonymous. Notes for authors. Marine Policy, 14(5):461– 462, September 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090015J.

Anonymous:1990:NAb

[Ano90k] Anonymous. Notes for authors. Marine Policy, 14(6):543– 544, November 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090030U.

Anonymous:1990:PMa

[Ano90l] Anonymous. Pages 103-180 (March 1990). Marine Policy, 14 (2):??, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:PMb

[Ano90m] Anonymous. Pages 183-296 (May 1990). Marine Policy, 14 (3):??, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:PJa

[Ano90n] Anonymous. Pages 2–100 (January 1990). Marine Policy, 14 (1):??, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:PJb

[Ano90o] Anonymous. Pages 299-376 (July 1990). Marine Policy, 14 (4):??, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:PS

[Ano90p] Anonymous. Pages 378-463 (September 1990). Marine Policy, 14(5):??, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:PN

[Ano90q] Anonymous. Pages 467-544 (November 1990). Marine Policy, 14(6):??, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1990:Pa

[Ano90r] Anonymous. Publications. Marine Policy, 14(4):375, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090066Z.

Anonymous:1990:Pb

[Ano90s] Anonymous. Publications. Marine Policy, 14(6):542, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090028P.

Anonymous:1991:Ca

[Ano91a] Anonymous. Calendar. Marine Policy, 15(1):72, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190049H.

Anonymous:1991:Cb

[Ano91b] Anonymous. Calendar. Marine Policy, 15(2):141, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X91900154.

Anonymous:1991:Cc

[Ano91c] Anonymous. Calendar. Marine Policy, 15(3):228, IBC, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9190068M.

Anonymous:1991:CPF

[Ano91d] Anonymous. Convention for the prohibition of fishing with long driftnets in the South Pacific. Marine Policy, 15(5):332– 334, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190086Q.

Anonymous:1991:EB

[Ano91e] Anonymous. Editorial Board. Marine Policy, 15(1):ifc, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9190037C.

Anonymous:1991:HCR

[Ano91f] Anonymous. House Congressional resolution 214. Marine Policy, 15(5):330–331, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190084O.

Anonymous:1991:IV

[Ano91g] Anonymous. Index to volume 15. Marine Policy, 15(6):466– 467, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190058J.

Anonymous:1991:NAa

[Ano91h] Anonymous. Notes for authors. Marine Policy, 15(2):143– 144, March 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X91900176.

Anonymous:1991:NAb

[Ano91i] Anonymous. Notes for authors. Marine Policy, 15(5):380, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190099W.

Anonymous:1991:PMb

[Ano91j] Anonymous. Pages 147-228 (May 1991). Marine Policy, 15 (3):??, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PJa

[Ano91k] Anonymous. Pages 2–72 (January 1991). Marine Policy, 15 (1):??, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PJb

[Ano91l] Anonymous. Pages 231-300 (July 1991). Marine Policy, 15 (4):??, July 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PS

[Ano91m] Anonymous. Pages 303-380 (September 1991). Marine Policy, 15(5):??, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PN

[Ano91n] Anonymous. Pages 383-467 (November 1991). Marine Policy, 15(6):??, November 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PMa

[Ano91o] Anonymous. Pages 75-144 (March 1991). Marine Policy, 15 (2):??, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1991:PIC

[Ano91p] Anonymous. Protocols I and II of the convention for the prohibition of fishing with long driftnets in the South Pacific. Marine Policy, 15(5):335–336, September 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190087R.

Anonymous:1991:P

[Ano91q] Anonymous. Publications. Marine Policy, 15(2):142, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X91900165.

Anonymous:1991:RSP

[Ano91r] Anonymous. Resolution of the 29th South Pacific Conference concerning Pelagic Drift-net fishing in the South Pacific Commission Area. Marine Policy, 15(5):329–330, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9190083N.

Anonymous:1991:SVG

[Ano91s] Anonymous. Statement: the view of the government of Japan on the conservation and rational utilisation of albacore resources in the South Pacific. Marine Policy, 15(5):336–337, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190088S.

Anonymous:1991:TD

[Ano91t] Anonymous. Tarawa declaration. Marine Policy, 15(5):329, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190082M.

Anonymous:1991:GAR

[Ano91u] Anonymous. UN general assembly resolution 44/225. Marine Policy, 15(5):331–332, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190085P.

Anonymous:1992:Ca

[Ano92a] Anonymous. Calendar. Marine Policy, 16(2):144, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290042N.

Anonymous:1992:Cb

[Ano92b] Anonymous. Calendar. Marine Policy, 16(3):238–239, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X92900876.

Anonymous:1992:Cc

[Ano92c] Anonymous. Calendar. Marine Policy, 16(4):334, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290048T.

Anonymous:1992:Cd

[Ano92d] Anonymous. Calendar. Marine Policy, 16(5):408, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290010M.

Anonymous:1992:Ce

[Ano92e] Anonymous. Calendar. Marine Policy, 16(6):491, November 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X92900762.

Anonymous:1992:EB

[Ano92f] Anonymous. Editorial Board. Marine Policy, 16(1):ifc, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290049U.

Anonymous:1992:IVJ

[Ano92g] Anonymous. Index to volume 16: January, March, May, July, September and November 1992. Marine Policy, 16(6):492– 496, November 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900773.

Anonymous:1992:PMb

[Ano92h] Anonymous. Pages 147-239 (May 1992). Marine Policy, 16 (3):??, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1992:PJ

[Ano92i] Anonymous. Pages 243-335 (July 1992). Marine Policy, 16 (4):??, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1992:PS

[Ano92j] Anonymous. Pages 338-408 (September 1992). Marine Policy, 16(5):??, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1992:PN

[Ano92k] Anonymous. Pages 411-496 (November 1992). Marine Policy, 16(6):??, November 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1992:PMa

[Ano92l] Anonymous. Pages 75-144 (March 1992). Marine Policy, 16 (2):??, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1992:P

[Ano92m] Anonymous. Publications. Marine Policy, 16(2):143, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290041M.

Anonymous:1993:BRI

[Ano93a] Anonymous. Book review: Interim report of the marine monitoring programme on the Braer oil-spill: Marine Laboratory Aberdeen, from a report compiled by J. M. Davies, I. M. Davies, A. D. Hawkins, R. Johnstone, A. McVicar, A. McLay, G. Topping (all MLA) and K. Whittle (Torry Research Laboratory). Marine Policy, 17(5):441–448, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900639.

Anonymous:1993:BHG

[Ano93b] Anonymous. Butterworth-heinemann guidelines for submission of electronic manuscripts. Marine Policy, 17(5):488, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390072B.

Anonymous:1993:Ca

[Ano93c] Anonymous. Calendar. Marine Policy, 17(1):69, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390010Z.

Anonymous:1993:Cb

[Ano93d] Anonymous. Calendar. Marine Policy, 17(2):149–150, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900385.

Anonymous:1993:Cc

[Ano93e] Anonymous. Calendar. Marine Policy, 17(3):230, May 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390079I.

Anonymous:1993:Cd

[Ano93f] Anonymous. Calendar. Marine Policy, 17(4):326, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390090P.

Anonymous:1993:Ce

[Ano93g] Anonymous. Calendar. Marine Policy, 17(5):486, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390071A.

Anonymous:1993:Cf

[Ano93h] Anonymous. Calendar. Marine Policy, 17(6):576, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390021T.

Anonymous:1993:EB

[Ano93i] Anonymous. Editorial Board. Marine Policy, 17(1):ifc, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390001J.

Anonymous:1993:IVJ

[Ano93j] Anonymous. Index to Volume 17 January, March, May, July, September and November 1993. Marine Policy, 17(6):573– 575, November 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X93900204.

Anonymous:1993:NAa

[Ano93k] Anonymous. Notes for authors. Marine Policy, 17(1):ibc, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390011Q.

Anonymous:1993:NAb

[Ano93l] Anonymous. Notes for authors. Marine Policy, 17(2):ibc, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900396.

Anonymous:1993:NAc

[Ano93m] Anonymous. Notes for authors. Marine Policy, 17(4):ibc, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390091G.

Anonymous:1993:NAd

[Ano93n] Anonymous. Notes for authors. Marine Policy, 17(5):ibc, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390073C.

Anonymous:1993:NAe

[Ano93o] Anonymous. Notes for authors. Marine Policy, 17(6):ibc, November 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390022U.

Anonymous:1993:PMb

[Ano93p] Anonymous. Pages 155-232 (May 1993). Marine Policy, 17 (3):??, May 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1993:PJb

[Ano93q] Anonymous. Pages 235-326 (July 1993). Marine Policy, 17 (4):??, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1993:PJa

[Ano93r] Anonymous. Pages 3–72 (January 1993). Marine Policy, 17 (1):??, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1993:PS

[Ano93s] Anonymous. Pages 332-488 (September 1993). Marine Policy, 17(5):??, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1993:PMa

[Ano93t] Anonymous. Pages 75-152 (March 1993). Marine Policy, 17 (2):??, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1994:A

[Ano94a] Anonymous. Announcement. Marine Policy, 18(6):541, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900809.

Anonymous:1994:BHG

[Ano94b] Anonymous. Butterworth-heinemann guidelines for submission of electronic manuscripts. Marine Policy, 18(6):538, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900787.

Anonymous:1994:Ca

[Ano94c] Anonymous. Calendar. Marine Policy, 18(3):288, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900086.

Anonymous:1994:Cb

[Ano94d] Anonymous. Calendar. Marine Policy, 18(4):359, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900507.

Anonymous:1994:Cc

[Ano94e] Anonymous. Calendar. Marine Policy, 18(6):539, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900795.

Anonymous:1994:EB

[Ano94f] Anonymous. Editorial Board. Marine Policy, 18(1):ifc, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900833.

Anonymous:1994:IV

[Ano94g] Anonymous. Index to volume 18. Marine Policy, 18(6):542– 544, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900817.

Anonymous:1994:LDL

[Ano94h] Anonymous. Lisbon declaration: Lisbon declaration on implementation by local authorities of the marine chapter of UNCED’s Agenda 21. Marine Policy, 18(2):94–98, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900116.

Anonymous:1994:NAa

[Ano94i] Anonymous. Notes for authors. Marine Policy, 18(2):ibc, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900329.

Anonymous:1994:NAb

[Ano94j] Anonymous. Notes for authors. Marine Policy, 18(3):ibc, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900094.

Anonymous:1994:NAc

[Ano94k] Anonymous. Notes for authors. Marine Policy, 18(4):ibc, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900515.

Anonymous:1994:NAd

[Ano94l] Anonymous. Notes for authors. Marine Policy, 18(5):ibc, September 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900434.

Anonymous:1994:NAe

[Ano94m] Anonymous. Notes for authors. Marine Policy, 18(6):ibc, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900825.

Anonymous:1994:PM

[Ano94n] Anonymous. Pages 219-288 (May 1994). Marine Policy, 18 (3):??, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1994:PJb

[Ano94o] Anonymous. Pages 291-359 (July 1994). Marine Policy, 18 (4):??, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1994:PJa

[Ano94p] Anonymous. Pages 3–81 (January 1994). Marine Policy, 18 (1):??, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1994:PS

[Ano94q] Anonymous. Pages 363-439 (September 1994). Marine Policy, 18(5):??, September 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1995:C

[Ano95a] Anonymous. Calendar. Marine Policy, 19(4):365, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9590006G.

Anonymous:1995:EB

[Ano95b] Anonymous. Editorial Board. Marine Policy, 19(1):ifc, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9590003D.

Anonymous:1995:IV

[Ano95c] Anonymous. Index to volume 19. Marine Policy, 19(6):509– 510, November 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9590007H.

Anonymous:1995:PMb

[Ano95d] Anonymous. Pages 171-248 (May 1995). Marine Policy, 19 (3):??, May 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1995:PJ

[Ano95e] Anonymous. Pages 3–74 (January 1995). Marine Policy, 19 (1):??, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1995:PS

[Ano95f] Anonymous. Pages 371-448 (September 1995). Marine Policy, 19(5):??, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1995:PMa

[Ano95g] Anonymous. Pages 83-166 (March 1995). Marine Policy, 19 (2):??, March 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1996:ARM

[Ano96a] Anonymous. Agreement regarding MS/Estonia. Marine Policy, 20(4):355–356, July 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X96900111.

Anonymous:1996:Ca

[Ano96b] Anonymous. Calendar. Marine Policy, 20(3):i, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96900172.

Anonymous:1996:Cb

[Ano96c] Anonymous. Calendar. Marine Policy, 20(4):I, July 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96900135.

Anonymous:1996:Cc

[Ano96d] Anonymous. Calendar. Marine Policy, 20(5):I, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X96900196.

Anonymous:1996:Cd

[Ano96e] Anonymous. Calendar. Marine Policy, 20(6):III, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X96900214.

Anonymous:1996:CRI

[Ano96f] Anonymous. Coastal resources and integrated coastal development. Marine Policy, 20(2):171–173, March 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00800019.

Anonymous:1996:EB

[Ano96g] Anonymous. Editorial Board. Marine Policy, 20(1):ifc, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X96900226.

Anonymous:1996:IV

[Ano96h] Anonymous. Index to volume 20. Marine Policy, 20(6):I– II, November 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X96900202.

Anonymous:1996:IGS

[Ano96i] Anonymous. An international guide to some of the major events in 1996. Marine Policy, 20(2):I, March 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96900032.

Anonymous:1996:PJ

[Ano96j] Anonymous. Pages 1–97 (January 1996). Marine Policy, 20 (1):??, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1996:PMb

[Ano96k] Anonymous. Pages 183-281 (May 1996). Marine Policy, 20 (3):??, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1996:PS

[Ano96l] Anonymous. Pages 357-438 (September 1996). Marine Policy, 20(5):??, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1996:PN

[Ano96m] Anonymous. Pages 439-524 (November 1996). Marine Policy, 20(6):??, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1996:PMa

[Ano96n] Anonymous. Pages 99-181 (March 1996). Marine Policy, 20 (2):??, March 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1997:Ca

[Ano97a] Anonymous. Calendar. Marine Policy, 21(1):I, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97900075.

Anonymous:1997:Cb

[Ano97b] Anonymous. Calendar. Marine Policy, 21(2):I, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97900038.

Anonymous:1997:Cc

[Ano97c] Anonymous. Calendar. Marine Policy, 21(3):I, May 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97900051.

Anonymous:1997:Cd

[Ano97d] Anonymous. Calendar. Marine Policy, 21(4):I, July 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97845952.

Anonymous:1997:Ce

[Ano97e] Anonymous. Calendar. Marine Policy, 21(5):I, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97900014.

Anonymous:1997:EB

[Ano97f] Anonymous. Editorial Board. Marine Policy, 21(1):ifc, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97900063.

Anonymous:1997:IV

[Ano97g] Anonymous. Index to volume 21 (1997). Marine Policy, 21(6): I–II, November 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9780036X.

Anonymous:1997:PJ

[Ano97h] Anonymous. Pages 1–109 (January 1997). Marine Policy, 21 (1):??, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1997:PMa

[Ano97i] Anonymous. Pages 111-196 (March 1997). Marine Policy, 21 (2):??, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1997:PMb

[Ano97j] Anonymous. Pages 197-288 (May 1997). Marine Policy, 21 (3):??, May 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1997:PS

[Ano97k] Anonymous. Pages 409-491 (September 1997). Marine Policy, 21(5):??, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1997:PN

[Ano97l] Anonymous. Pages 493-549 (November 1997). Marine Policy, 21(6):??, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1998:Ca

[Ano98a] Anonymous. Calendar. Marine Policy, 22(1):I, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98900070.

Anonymous:1998:Cb

[Ano98b] Anonymous. Calendar. Marine Policy, 22(2):I, March 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98900033.

Anonymous:1998:Ea

[Ano98c] Anonymous. Editorial. Marine Policy, 22(3):175–181, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000049.

Anonymous:1998:Ec

[Ano98d] Anonymous. Editorial. Marine Policy, 22(3):183–184, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000141.

Anonymous:1998:EB

[Ano98e] Anonymous. Editorial Board. Marine Policy, 22(1):ifc, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98900057.

Anonymous:1998:PJ

[Ano98f] Anonymous. Pages 1–81 (January 1998). Marine Policy, 22 (1):??, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1998:PMb

[Ano98g] Anonymous. Pages 175-280 (May 1998). Marine Policy, 22 (3):??, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1998:PN

[Ano98h] Anonymous. Pages 437-544 (November 1998). Marine Policy, 22(6):??, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1998:PMa

[Ano98i] Anonymous. Pages 83-174 (March 1998). Marine Policy, 22 (2):??, March 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1999:E

[Ano99a] Anonymous. Editorial. Marine Policy, 23(4–5):271–273, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X9900010X.

Anonymous:1999:I

[Ano99b] Anonymous. Index. Marine Policy, 23(6):I–II, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000202.

Anonymous:1999:PJa

[Ano99c] Anonymous. Pages 1–116 (January 1999). Marine Policy, 23 (1):??, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1999:PMa

[Ano99d] Anonymous. Pages 117-190 (March 1999). Marine Policy, 23 (2):??, March 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1999:PMb

[Ano99e] Anonymous. Pages 191-270 (May 1999). Marine Policy, 23 (3):??, May 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1999:PJb

[Ano99f] Anonymous. Pages 271-536 (July 1999). Marine Policy, 23 (4–5):??, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1999:PN

[Ano99g] Anonymous. Pages 537-600 (November 1999). Marine Policy, 23(6):??, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:C

[Ano21a] Anonymous. Calendar. Marine Policy, 22(4–5):I, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X9890001X.

Anonymous:2021:PJS

[Ano21b] Anonymous. Pages 281-436 (July–September 1998). Marine Policy, 22(4–5):??, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Antoine:1992:M

[Ant92] M. Serge Antoine. (2) The Mediterranean. Marine Policy, 16(1):51–53, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290056U.

Ademuni-Odeke:1991:UMS

[AO91] Ademuni-Odeke. UK Merchant Shipping Act, 1988: Protection of shipping and trading interests. Marine Policy, 15(4): 272–288, July 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190004U.

Ademuni-Odeke:1993:BRC

[AO93] Ademuni-Odeke. Book review: Canadian ocean law and policy: edited by David VanderZwaag, Butterworths, Toronto and Vancouver, 1992, xxxiv and 546 pp, £73. Marine Policy, 17(2): 147–148, March 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X93900363.

Appleyard:1994:BRS

[App94a] W. Philip Appleyard. Book review: A strategy for fisheries development: by Eduardo A. Loayza in collaboration with Lucian M. Sprague, World Bank Discussion Papers, Fisheries Series, No 135, World Bank, Washington, DC, 1992, 88 pp. Marine Policy, 18(3):287, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900078.

Appleyard:1994:BRO

[App94b] W. Philip Appleyard. Book review: OECD fisheries review review of fisheries in OECD member countries: OECD France 1993 395 pp £67.00 ISBN 92-64-13958-3. Marine Policy, 18(5): 439, September 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900426.

Armstrong:1998:SFR

[Arm98] Claire W. Armstrong. Sharing a fish resource: Bargaining theoretical analysis of an applied allocation rule. Marine Policy, 22(2):119–134, March 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97800012.

Arnason:1993:OFM

[Arn93] Ragnar Arnason. Ocean fisheries management: recent international developments. Marine Policy, 17(5):334–339, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390041Z.

Bailey:1997:LIT

[Bai97] Conner Bailey. Lessons from Indonesia’s 1980 trawler ban. Marine Policy, 21(3):225–235, May 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000031.

Ball:1999:PWR

[Bal99a] Iwan Ball. Port waste reception facilities in UK ports Iwan Ball. Marine Policy, 23(4–5):307–327, July 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000578.

Ballinger:1999:EOF

[Bal99b] Rhoda C. Ballinger. The evolving organisational framework for Integrated Coastal Management in England and Wales. Marine Policy, 23(4–5):501–523, July 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000542.

Barabolia:1990:LNA

[Bar90a] P. D. Barabolia. Limiting naval armaments and controlling conversion from military production. Marine Policy, 14(2):

125–129, March 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090098C.

Barston:1990:SME

[Bar90b] R. P. Barston. 29th Session of Marine Environment Protection Committee of IMO: London, 12–16 March 1990. Marine Policy, 14(6):533–534, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090023K.

Barker:1992:HGS

[Bar92] Simon Barker. Hazardous goods at sea: Are safe ships and clean seas mutually exclusive? A Canadian perspective. Marine Policy, 16(4):306–332, July 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290046R.

Barcena:1994:MAU

[Bár94a] Alicia Bárcena. Marine agenda of UNCED: Role of the Earth Council. Marine Policy, 18(2):99–103, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900124.

Barston:1994:I

[Bar94b] Ronald Barston. Introduction. Marine Policy, 18(6):443–444, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900639.

Barston:1995:UNC

[Bar95] Ronald Barston. United Nations Conference on straddling and highly migratory fish stocks. Marine Policy, 19(2):159– 166, March 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9500001M.

Bateman:2021:EGM

[Bat21] Sam Bateman. Economic growth, marine resources and naval arms in east Asia — a deadly triangle? Marine Policy, 22(4– 5):297–306, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000419.

Bavinck:1996:FRA

[Bav96] Maarten Bavinck. Fisher regulations along the Coromandel coast: a case of collective control of common pool resources. Marine Policy, 20(6):475–482, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000395.

Blanco-Bazan:1996:IGP

[BB96] Agustín Blanco-Bazán. The IMO guidelines on particular sensitive sea areas (PSSAs). Their possible application to the protection of underwater cultural heritage. Marine Policy, 20(4): 343–349, July 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9600019X.

Boehmer-Christiansen:1992:BRE

[BC92] Sonja Boehmer-Christiansen. Book review: Effectiveness of international resource cooperation: Steinar Andresen and Jorgen Wettestad, Fridjof Nansen Institute, Lysaker, Norway, 1991. Marine Policy, 16(3):234–235, May 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X92900843.

Bjorndal:1998:RNM

[BC98] Trond Bjørndal and Jon M. Conrad. A report on the Norwegian Minke whale hunt. Marine Policy, 22(2):161–174, March 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000365.

Bache:1999:DAC

[BE99] Sali Jayne Bache and Nathan Evans. Dolphin, albatross and commercial fishing: Australia’s response to an unpalatable mix. Marine Policy, 23(3):259–270, May 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000384.

Bell:1993:PSC

[Bel93] Douglas Bell. Port state control v. flag state control: UK government position. Marine Policy, 17(5):367–370, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900499.

Bergholtsas:1990:LBF

[Ber90a] J. J. Bergholtsas. The Latvian ‘Baltica’ Foundation. Marine Policy, 14(2):177–179, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090105Z.

Bergstrand:1990:SMS

[Ber90b] Simon Bergstrand. Soviet merchant shipping: Some European attitudes explored. Marine Policy, 14(2):112–119, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090096A.

Bergin:1997:ALM

[Ber97] Anthony Bergin. Albatross and longlining — Managing seabird bycatch. Marine Policy, 21(1):63–72, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000401.

Bjorndal:1992:MSS

[BGS92] Trond Bjørndal, Daniel V. Gordon, and Kjell G. Salvanes. Markets for salmon in Spain and Italy. Marine Policy, 16(5):338–344, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X92900016.

Blake:1991:IBR

[BGW91] G. H. Blake and C. Grundy-Warr. International Boundaries Research Unit. Marine Policy, 15(1):67–68, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190045D.

Bergin:1992:JTF

[BH92] Anthony Bergin and Marcus Howard. Japanese tuna fisheries: Future trends. Marine Policy, 16(3):147–166, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X92900784.

Bergin:1994:SBT

[BH94] Anthony Bergin and Marcus Haward. Southern bluefin tuna fishery: Recent developments in international management. Marine Policy, 18(3):263–273, May 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900043.

Binkley:1991:NSO

[Bin91] Marian Binkley. Nova Scotian offshore fishermen’s awareness of safety. Marine Policy, 15(3):170–182, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190060O.

Birnie:1992:BRU

[Bir92] Patricia Birnie. Book review: United Nations Convention on the Law of the Sea 1982: a Commentary, Vol IV: Editorin-Chief M. Nordquist, Martinus Nijhoff, Dordrecht, The Netherlands, 1991. Marine Policy, 16(3):232–234, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X92900832.

Birnie:1993:UMM

[Bir93] Patricia Birnie. UNCED and marine mammals. Marine Policy, 17(6):501–514, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390013S.

Birnie:1994:MPL

[Bir94a] Patricia Birnie. Maritime policy and legal issues: Impact of the LOS convention and UNCED on UK maritime law and policy. Marine Policy, 18(6):483–493, November 1994.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900701.

Birnie:1994:BRL

[Bir94b] Patricia W. Birnie. Book review: Legal aspects of commercial sea-fishing in the EEC: Irish Centre for European Law Ltd. Trinity College, Dublin, 1992, pp 39 + 41 pp appendices. Marine Policy, 18(1):81, January 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900930.

Bisbal:1993:FMP

[Bis93] Gustavo A. Bisbal. Fisheries management on the Patagonian shelf: a decade after the 1982 Falklands/Malvinas conflict. Marine Policy, 17(3):213–229, May 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390078H.

Bisbal:1995:SSA

[Bis95] Gustavo A. Bisbal. The Southeast South American shelf large marine ecosystem: Evolution and components. Marine Policy, 19(1):21–38, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9592570W.

Bailey:1990:HCF

[BJ90] Conner Bailey and Svein Jentoft. Hard choices in fisheries development. Marine Policy, 14(4):333–344, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090055V.

Bantongbacal:1998:WCR

[BKLM98] Jay Bantongbacal, Lisa Kruger, Stefan Leslie, and Laurence Mazaudier. Workshop on CIDA’s regional oceans initiatives: CIDA Graduate Trainee Report. Marine Policy, 22(6):527– 535, November 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9800027X.

Blake:1991:BRR

[Bla91] Gerald Blake. Book review: Rights to oceanic resources: Deciding and drawing maritime boundaries: edited by Dorinda G. Dallmeyer and Louis de Vorsey, Martinus Nijhoff, Dordrecht, 1989, 207 pp. US$67.50/£ 39. Marine Policy, 15(1): 71–72, January 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190048G.

Blake:1993:BRN

[Bla93] Gerald Blake. Book review: North Sea atlas for Netherlands policy and management: ICONA, Stadsvitgeverij, Amsterdam, 1992, 96 maps. Marine Policy, 17(2):148–149, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900374.

Blake:1998:SCS

[Bla98] Byron Blake. A strategy for cooperation in sustainable oceans management and development, Commonwealth Caribbean. Marine Policy, 22(6):505–513, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000256.

Blonk:1994:DEM

[Blo94] Wim A. G. Blonk. Developments in EU maritime transport policy. Marine Policy, 18(6):476–482, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900698.

Bergin:1996:ASP

[BM96] Anthony Bergin and Frances B. Michaelis. Australia and the South Pacific: implementing the UNCED oceans agenda. Marine Policy, 20(1):47–62, January 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500023Y.

Boczek:1991:BSL

[Boc91] Boleslaw A. Boczek. Basic sources of law of the sea documentation. Marine Policy, 15(6):455–459, November 1991.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190053E.

Bohme:1992:OSD

[Böh92] Hans Böhme. Ocean shipping: delicate balance in maritime freight markets. Marine Policy, 16(4):333, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290047S.

Boisson:1994:CSS

[Boi94] Philippe Boisson. Classification societies and safety at sea: Back to basics to prepare for the future. Marine Policy, 18(5):363–377, September 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900337.

Boyle:1990:LSI

[Boy90] Alan E. Boyle. The Law of the Sea and international watercourses: an emerging cycle. Marine Policy, 14(2):151– 157, March 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090102W.

Boyle:1992:LBS

[Boy92a] Alan E. Boyle. Land-based sources of marine pollution: Current legal regime. Marine Policy, 16(1):20–35, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290052Q.

Boyle:1992:PME

[Boy92b] Alan E. Boyle. Protecting the marine environment: Some problems and developments in the law of the sea. Marine Policy, 16(2):79–85, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290028N.

Brown:1999:CMC

[BP99] David N. Brown and Robert S. Pomeroy. Co-management of Caribbean Community (CARICOM) fisheries. Marine Policy, 23(6):549–570, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000402.

Brigstocke:1994:UMD

[Bri94] J. R. Brigstocke. UK maritime defence. Marine Policy, 18 (6):530–533, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X94900760.

Brice:1996:SUC

[Bri96] Q. C. Geoffrey Brice. Salvage and the underwater cultural heritage. Marine Policy, 20(4):337–342, July 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9600022X.

Browes:1992:NA

[Bro92] Pauline Browes. (4) North America. Marine Policy, 16(1):55– 56, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290058W.

Brown:1993:BRIb

[Bro93a] E. D. Brown. Book review: International maritime conventions: edited by Ignacio Arroyo, Kluwer, Deventer and Boston, 1991, Dfl 455, US$237. Marine Policy, 17(2):146, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390034Z.

Brown:1993:BRIa

[Bro93b] E. D. Brown. Book review: International organisations and the law of the sea, vol 5 1989 and vol 6 1990: Edited by A. H. A Soons and B. Kwiat-kowska Graham and Trotman/Martinus Nijhoff, London, Dordrecht, Boston, 1991 and 1992, Dfl395 per vol. Marine Policy, 17(1):67, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390008Q.

Brown:1993:BRM

[Bro93c] E. D. Brown. Book review: Marine management in disputed areas the case of the Barents Sea: Robin Churchill and Geir Ulfstein, Ocean management and policy series, Routledge, London and New York, 1992. Marine Policy, 17(1): 67, January 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390009R.

Brown:1993:E

[Bro93d] E. D. Brown. Editorial. Marine Policy, 17(1):3, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390002K.

Brown:1993:NNN

[Bro93e] E. D. Brown. ‘Neither necessary nor prudent at this stage’: the regime of seabed mining and its impact on the universality of the UN convention on the law of the sea. Marine Policy, 17(2): 81–107, March 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390024W.

Brown:1994:E

[Bro94] E. D. Brown. Editorial. Marine Policy, 18(1):3, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900841.

Brown:1995:AIP

[Bro95a] E. D. Brown. The 1994 Agreement on the Implementation of Part XI of the UN Convention on the Law of the Sea: breakthrough to universality? Marine Policy, 19(1):5–20, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9592569S.

Brown:1995:E

[Bro95b] E. D. Brown. Editorial. Marine Policy, 19(1):3, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9590005F.

Brown:1996:E

[Bro96a] E. D. Brown. Editorial. Marine Policy, 20(1):1–2, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X96900238.

Brown:1996:PUC

[Bro96b] E. D. Brown. Protection of the underwater cultural heritage. Draft principles and guidelines for implementation of article 303 of the United Nations Convention on the law of the sea, 1982. Marine Policy, 20(4):325–336, July 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000140.

Broeze:1997:GNI

[Bro97a] Frank Broeze. Geostrategy and navyports in the Indian Ocean since c. 1970. Marine Policy, 21(4):345–362, July 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000134.

Brown:1997:DSL

[Bro97b] E. D. Brown. Dispute settlement and the law of the sea: the UN convention regime. Marine Policy, 21(1):17–43, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X96000437.

Brown:2021:MVS

[Bro21] E. D. Brown. The M/V ‘Saiga’ case on prompt release of detained vessels: the first judgment of the International Tribunal for the Law of the Sea. Marine Policy, 22(4–5):307– 326, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000189.

Batstone:1999:NZQ

[BS99] C. J. Batstone and B. M. H. Sharp. New Zealand’s quota management system: the first ten years. Marine Policy, 23(2): 177–190, March 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000360.

Budd:1999:AEA

[Bud99] M. Budd. The application of environmental assessment to marine developments and activities in Great Britain. Marine Policy, 23(4–5):439–451, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000505.

Burke:1993:UO

[Bur93] William T. Burke. UNCED and the oceans. Marine Policy, 17(6):519–533, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390015U.

Butler:1990:JVS

[But90a] W. E. Butler. Joint ventures and the Soviet Arctic. Marine Policy, 14(2):169–176, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090104Y.

Butler:1990:LSA

[But90b] W. E. Butler. The law of the sea and Anglo-Soviet legal relations. Marine Policy, 14(2):103–105, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900948.

Butler:1992:PSM

[But92] W. E. Butler. Perestroika and Soviet maritime law. Marine Policy, 16(2):75–78, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290027M.

Barg:1994:EMO

[BW94] U. Barg and U. N. Wijkström. Environmental management options for coastal fisheries and aquaculture: Role of local authorities. Marine Policy, 18(2):127–137, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900183.

Bywater:1995:GRM

[Byw95] John Bywater. Government response to marine pollution from ships. Marine Policy, 19(6):487–496, November 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9500031Z.

Caddy:1996:OAN

[Cad96] J. F. Caddy. An objective approach to the negotiation of allocations from shared living resources. Marine Policy, 20(2): 145–155, March 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X95000414.

Campbell:1994:IYT

[Cam94] Harry F. Campbell. Investing in yellowfin tuna: the economics of conservation. Marine Policy, 18(1):19–28, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900868.

Campbell:1996:PIT

[Cam96] H. F. Campbell. Prospects for an international tuna resource owners’ cartel. Marine Policy, 20(5):419–427, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X96000279.

Calmet:1991:RWO

[CB91] D. P. Calmet and J. M. Bewers. Radioactive waste and ocean dumping: the role of the IAEA. Marine Policy, 15(6):413– 430, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190051C.

Clinton-Davis:1994:UEG

[CD94] Lord Clinton-Davis. The UK, Europe and the global maritime trading system. Marine Policy, 18(6):472–475, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9490068X.

Chambolle:1992:RE

[Cha92a] Thierry Chambolle. (6) The role of enterprise. Marine Policy, 16(1):63–64, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290065W.

Charles:1992:FCU

[Cha92b] Anthony T. Charles. Fishery conflicts: a unified framework. Marine Policy, 16(5):379–393, September 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290006B.

Cheema:1994:US

[Che94] G. Shabbir Cheema. UNDP statement. Marine Policy, 18(2): 104–106, March 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900132.

Churchill:1993:FIM

[Chu93] R. R. Churchill. Fisheries issues in maritime boundary delimitation. Marine Policy, 17(1):44–57, January 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390005N.

Cissokho:1992:A

[Cis92] Cheikh Cissokho. (5) Africa. Marine Policy, 16(1):56–57, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290059X.

Cole-King:1993:MCN

[CK93a] Adam Cole-King. Marine conservation: a new policy area. Marine Policy, 17(3):171–185, May 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390075E.

Cooper:1993:AEW

[CK93b] Alastair Cooper and Ian Kinniburgh. After effects and the way forward: community impact assessment. Marine Policy, 17(5):479–482, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390069F.

Clark:1993:ITQ

[Cla93] Ian Clark. Individual transferable quotas: the New Zealand experience. Marine Policy, 17(5):340–342, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900422.

Clement:1996:BRC

[Clé96] Etienne Clément. Book review: Current developments at UN-ESCO concerning the protection of the underwater cultural heritage. Presentation made at the first and the second national maritime museum conferences on the protection of underwater cultural heritage (Greenwich, 3 and 4 February 1995) (London, IMO, 25 and 26 January 1996). Marine Policy, 20(4): 309–323, July 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000152.

Cook:1991:MNB

[CM91] B. A. Cook and Richard McGaw. Management of the New Brunswick commercial salmon fishery: an economic perspective. Marine Policy, 15(1):33–38, January 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X91900419.

Campbell:1997:CAD

[CM97] H. F. Campbell and A. McIlgorm. Comparative advantage and distant water fishing fleets: the Australian east coast tuna longline fishery. Marine Policy, 21(6):493–499, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000146.

Chakalall:1998:CIF

[CMM98] Bisessar Chakalall, Robin Mahon, and Patrick McConney. Current issues in fisheries governance in the Caribbean Community (CARICOM). Marine Policy, 22(1):29–44, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97800024.

Collins:1990:AEP

[Col90] Ken Collins. Action by the European Parliament. Marine Policy, 14(3):275–277, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X90900889.

Cook:1994:ICA

[Coo94] B. A. Cook. International cooperative agreements: Scotia-Fundy herring fisheries. Marine Policy, 18(3):275–283, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900051.

Cormack:1992:STB

[Cor92] Douglas Cormack. Scientific and technical background. Marine Policy, 16(1):5–19, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290051P.

Corbet:1995:NMP

[Cor95] A. G. Corbet. Navigation management: post-Donaldson. Marine Policy, 19(6):477–486, November 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500030A.

Cotti:1990:EBC

[Cot90] Flavio Cotti. Evolution of the Basle Convention. Marine Policy, 14(3):210–213, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X90900753.

Couper:1996:PIU

[Cou96] Alastair Couper. The principal issues in underwater cultural heritage. Marine Policy, 20(4):283–285, July 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000231.

Coull:1999:CBF

[Cou99] J. R. Coull. Changing balance of fish production in Scotland. Marine Policy, 23(4–5):347–358, July 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98900483.

Cox:1993:MTI

[Cox93] Derek Cox. Marine — terrestrial interface: the public health dimension. Marine Policy, 17(5):453–456, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390065B.

Cockcroft:1999:CFM

[CP99] Andrew C. Cockcroft and Andrew I. L. Payne. A cautious fisheries management policy in South Africa: the fisheries for rock lobster. Marine Policy, 23(6):587–600, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000451.

Crean:1999:CCB

[Cre99] Kevin Crean. Centralised and community-based fisheries management strategies: case studies from two fisheries dependent archipelagos. Marine Policy, 23(3):243–257, May 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000372.

Crickard:1995:COM

[Cri95] Fred Crickard. Canada’s ocean and maritime security a strategic forecast: a strategic forecast. Marine Policy, 19(4):335– 342, July 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9500012U.

Crean:1994:DPT

[CS94] Kevin Crean and David Symes. The discards problem: towards a European solution. Marine Policy, 18(5):422–434, September 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9490037X.

Cicin-Sain:1996:ESI

[CS96] Biliana Cicin-Sain. Earth summit implementation: progress since Rio. Marine Policy, 20(2):123–143, March 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000024.

Couper:1997:DFB

[CS97] A. D. Couper and H. D. Smith. The development of fishermenbased policies. Marine Policy, 21(2):111–119, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000498.

Curtis:1992:VG

[Cur92] Clifton Curtis. (8) Views of Greenpeace. Marine Policy, 16(1): 67–68, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290067Y.

Chircop:1995:GMA

[CVM95] Aldo Chircop, David VanderZwwag, and Peter Mushkat. The Gulf of Maine Agreement and Action Plan: a novel but nascent approach to transboundary marine environmental protection. Marine Policy, 19(4):317–333, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9500013V.

Dahl:1993:LBP

[Dah93] Arthur Lyon Dahl. Land-based pollution and integrated coastal management. Marine Policy, 17(6):561–572, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390019Y.

Davies:1990:PCM

[Dav90a] J. E. Davies. Promoting competition and market performance in liner shipping. Marine Policy, 14(6):477–483, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090017L.

Davis:1990:CR

[Dav90b] Stanley Clinton Davis. Chairman’s review. Marine Policy, 14 (3):188–190, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090071X.

Davis:1990:GAM

[Dav90c] W. Jackson Davis. Global aspects of marine pollution policy: the need for a new international convention. Marine Policy, 14 (3):191–197, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090072Y.

Day:1995:PPO

[Day95a] Douglas Day. Public policy and ocean management in Canada. Marine Policy, 19(4):251–256, July 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X95000193.

Day:1995:TAH

[Day95b] Douglas Day. Tending the achilles’ heel of NAFO: Canada acts to protect the nose and tail of the Grand Banks. Marine Policy, 19(4):257–270, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X95000207.

deCuellar:1990:MSG

[dC90] Javier Perez de Cuellar. Message from UN secretarygeneral. Marine Policy, 14(3):187–188, May 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900708.

Devine:1993:TNS

[Dev93] D. J. Devine. Transition to a new South Africa: Marine law issues to be resolved? Marine Policy, 17(4):303–308, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390085H.

Devine:1994:AMP

[Dev94] D. J. Devine. S. African multi-party negotiations forum: marine law issues. Marine Policy, 18(1):79, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900914.

Devine:1996:SDJ

[Dev96] D. J. Devine. Ships in distress — a judicial contribution from the South Atlantic. Marine Policy, 20(3):229–234, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9400000I.

Dickie:1992:GCO

[Dic92] Paul Dickie. The Gulf crisis one year on: Problems for international traders. Marine Policy, 16(2):107–110, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290033L.

Dijxhoorn:1993:PSM

[Dij93] O. H. J. Dijxhoorn. Port and shipping management: the role of IMO. Marine Policy, 17(5):363–366, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900488.

Dioh:1998:SPF

[Dio98] Bernard Codou Dioh. Support programme for fisheries management in West Africa (AGREH). Marine Policy, 22(6):441– 453, November 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000207.

diMeana:1990:HWS

[dM90] Carlo Ripa di Meana. Hazardous waste shipments: Regulation within the EC. Marine Policy, 14(3):271–274, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900878.

diMeana:1992:EC

[dM92] Carlo Ripa di Meana. (3) The European Commission. Marine Policy, 16(1):60, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290062T.

Davidse:1999:PRF

[DMV99] W. P. Davidse, L. V. McEwan, and N. Vestergaard. Property rights in fishing: from state property towards private property?: a case study of three EU countries. Marine Policy, 23(6):537–547, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000396.

Dexter:1998:WMC

[DN98] Peter E. Dexter and Gus V. Necco. WMO and marine capacity building. Marine Policy, 22(3):217–227, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000086.

Donoghue:1993:EPP

[Don93] Joan E. Donoghue. EC participation in protection of the marine environment: Legal and institutional aspects. Marine Policy, 17(6):515–518, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390014T.

Doulman:1993:CBF

[Dou93] David J. Doulman. Community-based fishery management: Towards the restoration of traditional practices in the South Pacific. Marine Policy, 17(2):108–117, March 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390025X.

Doulman:1990:SPR

[DT90] David J. Doulman and Peter Terawasi. The South Pacific regional register of foreign fishing vessels. Marine Policy, 14(4): 324–332, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090054U.

Dubinksy:1991:BRM

[Dub91a] William Dubinksy. Book review: Management of world fisheries: Implication of extended fisheries jurisdiction: edited by Edward Miles, University of Washington Press, Seattle, WA, 1989. Marine Policy, 15(4):297–300, July 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190007X.

Dubinksy:1991:BRP

[Dub91b] William Dubinksy. Book review: The provident sea: D. H. Cushing, Cambridge University Press, Cambridge, UK, 1988. Marine Policy, 15(4):297, July 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190006W.

Dunnet:1993:CPS

[Dun93] George M. Dunnet. Control of pollution of the sea: workshop presentation. Marine Policy, 17(5):426–429, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390060G.

Durand:1994:EPE

[Dur94] Guy Durand. Environmental protection in EC less-developed regions: Role of regional development policy. Marine Policy, 18(2):148–152, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X94900205.

Davis:1990:DDN

[DV90a] W. Jackson Davis and Jon M. Van Dyke. Dumping of decommissioned nuclear submarines at sea: a technical and legal analysis. Marine Policy, 14(6):467–476, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090016K.

duVivier:1990:CWE

[dV90b] Francois Roelants du Vivier. Control of waste exports to the Third World: the view of the European environmental alliance. Marine Policy, 14(3):264–270, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900867.

deVivero:1997:DRC

[dVdLE97] J. L. Suárez de Vivero, M. Frieyro de Lara, and J. Jurado Estévez. Decentralization, regionalization and co-management. A critical view on the viability of the alternative management models for fisheries in Spain. Marine Policy, 21(3): 197–206, May 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000018.

deVivero:1994:SMP

[dVF94] J. L. Suárez de Vivero and M. C. Frieyro. Spanish marine policy: Role of Marine Protected Areas. Marine Policy, 18(4): 345–352, July 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900477.

Dubbink:1996:MRV

[DvV96] Wim Dubbink and Martijn van Vliet. Market regulation versus co-management?: Two perspectives on regulating fisheries compared. Marine Policy, 20(6):499–516, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000358.

Edwards:1994:CEI

[Edw94] Simon Edwards. Coastal electronic information systems. Marine Policy, 18(4):357–358, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X94900493.

Egiyan:1990:FCO

[Egi90] G. S. Egiyan. ‘Flag of convenience’ or ‘open registration’ of ships. Marine Policy, 14(2):106–111, March 1990.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900959.

Edwards:1993:CPM

[EHP93] Simon Edwards, John Hobson, and Helen Pickering. ‘Coastal planning and management’: Winchester, UK, 27 November 1992. Marine Policy, 17(2):139–140, March 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900307.

Eikeland:1993:DAM

[Eik93] Per Ove Eikeland. Distributional aspects of multispecies management: the Barents Sea large marine ecosystem. Marine Policy, 17(4):256–271, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9390082E.

Ekpere:1998:PAO

[Ekp98] Johnson Ekpere. Pan-African ocean environment management programme. Marine Policy, 22(6):515–525, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000268.

Evans:1996:LOR

[Eva96] Nathan Evans. LOSC, offshore resources and Australian marine policy. Marine Policy, 20(3):223–227, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000085.

Eythorsson:1996:TPI

[Eyt96] Einar Eythórsson. Theory and practice of ITQs in Iceland. Privatization of common fishing rights. Marine Policy, 20(3): 269–281, May 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000097.

Farrow:1996:MPA

[Far96] Scott Farrow. Marine protected areas: emerging economics. Marine Policy, 20(6):439–446, November 1996.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000346.

Felt:1990:BUP

[Fel90a] Lawrence Felt. Barriers to user participation in the management of the Canadian Atlantic salmon fishery: If wishes were fishes. Marine Policy, 14(4):345–360, July 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090056W.

Felt:1990:BRC

[Fel90b] Lawrence F. Felt. Book review: Co-operative management of local fisheries: New directions for improved management and community development: edited by Evelyn Pinkerton, University of British Columbia Press, Vancouver, 1989, 229 pp, £17.55. Marine Policy, 14(6):536–537, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090025M.

Fenger:1994:SEH

[Fen94] Bent H. Fenger. Solving environmental health problems at local level. Marine Policy, 18(2):116–120, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900167.

Francis:1993:FMI

[FGA93] R. I. C. C. Francis, D. J. Gilbert, and J. H. Annala. Fishery management by individual quotas: theory and practice. Marine Policy, 17(1):64–65, January 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390007P.

Freese:1995:DAR

[FGS95] Stephen Freese, James Glock, and Dale Squires. Direct allocation of resources and cost-benefit analysis in fisheries: an application to Pacific whiting. Marine Policy, 19(3):199–211, May 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9400005D.

Fogarty:1997:CPM

[FHG97] Michael J. Fogarty, Ray Hilborn, and Donald Gunderson. Chaos and parametric management. Marine Policy, 21(2):187– 194, March 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X96000504.

Fogarty:1995:CCC

[Fog95] Michael J. Fogarty. Chaos, complexity and community management of fisheries: an appraisal. Marine Policy, 19(5):437– 444, September 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9598048K.

Franckx:1990:IPW

[Fra90a] Erik Franckx. Innocent passage of warships: Recent developments in US-Soviet relations. Marine Policy, 14(6):484–490, November 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090018M.

Frankel:1990:PES

[Fra90b] Ernst G. Frankel. Promoting efficiency in shipping. Marine Policy, 14(5):438–448, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090005C.

Frankel:1993:TQM

[Fra93] Ernst G. Frankel. Total quality management in liner shipping. Marine Policy, 17(1):58–63, January 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390006O.

Freestone:1990:WAO

[Fre90a] David Freestone. Workshop on adaptive options and policy implications of sea level rise and other impacts of global climate change: Miami, FL, 27 November–1 December 1989. Marine Policy, 14(5):456–457, September 1990.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090011F.

Fretheim:1990:DS

[Fre90b] Atle Fretheim. Dumping at sea. Marine Policy, 14(3):247– 250, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900823.

Findlay:1998:NEA

[FS98] M. Findlay and A. E. Searle. The North East Atlantic albacore fishery: a Cornish crisis of confidence. Marine Policy, 22(2): 95–108, March 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000341.

Gabaldon:1990:DTD

[Gab90] Arnoldo José Gabaldón. The disposal of toxic or dangerous waste in Latin America and the Caribbean. Marine Policy, 14 (3):254–258, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900845.

Gabaldon:1992:TT

[Gab92] Arnoldo José Gabaldón. (2) Transfer of technology. Marine Policy, 16(1):71–72, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X92900692.

Gao:1991:CC

[Gao91] Zhiguo Gao. China and the LOS convention. Marine Policy, 15 (3):199–209, May 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190063H.

Garrod:1990:PNE

[Gar90] Brian Garrod. Policy note: EFJ and price instability: a current cause for concern? Marine Policy, 14(1):86–88, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090041O.

Garick:1993:COI

[Gar93] Jason A. Garick. Crisis in the oil industry: Certificates of Financial Responsibility and the Oil Pollution Act of 1990. Marine Policy, 17(4):272–293, July 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390083F.

Gauci:1994:BRT

[Gau94] Gotthard Gauci. Book review: Transnational environmental liability and insurance: edited by R. P. Kroner, Graham and Trotman UK 1993 300 pp £78.00 ISBN 1-85333-7881. Marine Policy, 18(5):438, September 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9490040X.

Gauci:1995:LLM

[Gau95] Gotthard Gauci. Limitation of liability in maritime law: an anachronism? Marine Policy, 19(1):65–74, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9592573P.

Gezelius:1999:LEE

[Gez99] Stig S. Gezelius. Limits to externalisation: the EU NAFO policy 1979–1997. Marine Policy, 23(2):147–159, March 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000347.

Garza-Gil:1996:SCR

[GGIMSRVL96] Dolores Garza-Gil, Carlos Iglesias-Malvido, Juan C. Suris-Regueiro, and Manuel M. Varela-Lafuente. The Spanish case regarding fishing regulation. Marine Policy, 20(3):249–259, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X96000073.

Gibson:1993:CZP

[Gib93] John Gibson. Coastal zone planning law: Role of law in management of the coastal zone in England and Wales.

Marine Policy, 17(2):118–129, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390026Y.

Gilmore:1991:DTS

[Gil91] William C. Gilmore. Drug trafficking by sea: the 1988 United Nations convention against illicit traffic in narcotic drugs and psychotropic substances. Marine Policy, 15(3):183– 192, May 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190061F.

Gilmore:1996:NIS

[Gil96] William C. Gilmore. Narcotics interdiction at sea: the 1995 Council of Europe Agreement. Marine Policy, 20(1):3–14, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9500022X.

Gitsu:1992:RMS

[Git92] M. A. Gitsu. The registration of merchant ships under, perestroika: Soviet legislation. Marine Policy, 16(2):111–114, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290034M.

Glassner:1990:BRC

[Gla90a] Martin Ira Glassner. Book review: Canada’s Arctic waters in international law: by Donat Pharand, Cambridge University Press, Cambridge, USA, 1988, 288 pp. US$49.50. Marine Policy, 14(1):98–99, January 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090048V.

Glassner:1990:BRL

[Gla90b] Martin Ira Glassner. Book review: The legal regime of enclosed or semi-enclosed seas: the particular case of the Mediterranean: Volume 19, number 22 of Prinosi za Poredbeno Proučavanje Prava i Medunarodno Pravo (Contributions to the Study of Comparative and International Law) edited by Budislav Vukas, The Institute of International Law and International Relations, Faculty of Law, University of Zagreb, 1988, 515 pp. Marine Policy, 14(1):98–99, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090047U.

Glassner:1991:BRA

[Gla91a] Martin Ira Glassner. Book review: Annual review of ocean affairs: Law and policy, main documents 1988: Volume III compiled and edited by the United Nations Office for Ocean affairs and the Law of the Sea, UNIFO Publishers, Sarasota, FL 1990, xi + 473 pp, $92.50 casebound. Marine Policy, 15(6):464– 465, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190056H.

Glassner:1991:BRM

[Gla91b] Martin Ira Glassner. Book review: Managing marine environments: Richard A. Kenchington, Taylor and Francis, New York, 1990, 248 + iv pp, $49.50. Marine Policy, 15(5):374– 375, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190095S.

Glassner:1991:BRE

[Gla91c] Martin Ira Glassner. Book review: The exclusive economic zone: Regime and legal nature under international law, by Francisco Orrego Vicuña, Cambridge University Press, Cambridge, UK, 1989, 309 + ix pp. Marine Policy, 15(5):374– 375, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190094R.

Glassner:1992:BRT

[Gla92] Martin Ira Glassner. Book review: The times atlas and encyclopaedia of the sea: edited by Alastair Couper, Harper and Row, New York, 1990, 272pp, US$65.00. Marine Policy, 16 (2):141, March 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290039R.

Gavioli:1990:WMR

[GN90] Giuseppe Gavioli and Giovanni Nespoli. Waste material and the role of regional and local governments in Italy. Marine Policy, 14(3):283–285, May 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090090E.

Godinho:1994:IEQ

[God94] Rui Godinho. Improving the environmental quality of the Tagus estuary: Lisbon’s wastewater treatment system. Marine Policy, 18(2):186–188, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900264.

Gold:1990:DHN

[Gol90] Edgar Gold. Dangerous, hazardous and noxious cargoes: a new role for maritime arbitration? Marine Policy, 14(5):378– 384, September 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900018.

Gold:1991:BRC

[Gol91] Edgar Gold. Book review: Canadian oceans policy National strategies and the new law of the Sea: edited by Donald McRae and Gordon Munro, University of British Columbia Press, Vancouver, 1989, 268 pp, hardback C$39.95, softback C$22.95. Marine Policy, 15(1):69–71, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190047F.

Goodlad:1993:SFM

[Goo93] John Goodlad. Sea fisheries management: the Shetland position. Marine Policy, 17(5):350–351, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900455.

Goodman:1997:BRI

[Goo97] Dan Goodman. Book review: The international politics of whaling: Peter J. Stoett UBC Press, 1997, 228 pp. $65.00.

ISBN 0-7748-0605-2. Marine Policy, 21(6):547–549, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97800358.

Goss:1993:BRE

[Gos93a] P. O. Goss. Book review: The economics of coastal management: a Manual of Benefit Assessment Techniques, by E. C. Penning-Rowsell, et al., Belhaven Press, London and Florida, 1992, £99. Marine Policy, 17(2):145–146, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390033Y.

Goss:1993:BRC

[Gos93b] Richard Goss. Book review: Current issues in maritime economics: edited by K. W. Gwilliam, Kluwer, Dordrecht, the Netherlands, 1993, $108.50, £66, DFL170. Marine Policy, 17 (4):325, July 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390089L.

Gardner:1999:LBJ

[GP99a] B. M. Gardner and S. J. Pettit. The land-based jobs market for seafarers: Consequences of market imbalance and policy implications. Marine Policy, 23(2):161–175, March 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000359.

Gardner:1999:SLB

[GP99b] B. M. Gardner and S. J. Pettit. Seafarers and the land based jobs market: the present UK situation. Marine Policy, 23(1): 103–115, January 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000335.

Gardner:1996:SCB

[GPT96] B. M. Gardner, S. J. Pettit, and H. A. Thanopoulou. Shifting challenges for British maritime policy: a post-war review. Marine Policy, 20(6):517–524, November 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000383.

Grant:1990:BRM

[Gra90] John P. Grant. Book review: The Mediterranean continental shelf delimitation and legal regime: by Umberto Leanza and Luigi Sico, Oceana Publications, Dobbs Ferry NY, 1988, 4 volumes, $250.00. Marine Policy, 14(4):374–375, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090065Y.

Graham:1991:BRO

[Gra91] Gerald Graham. Book review: Our common seas: Coasts in crisis, by Don Hinrichsen, Earthscan Publications, London, in association with United Nations Environment Programme, Nairobi, 1990, 184 pp, £6.50. Marine Policy, 15(5):378–379, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190098V.

Gordon:1998:ITC

[GRBC98] Melissa Lee Gordon, Margaret A. Reams, and Tina Bernd-Cohen. Integrated theory of coastal management regarding protection of beaches and dunes. Marine Policy, 22(2):83– 93, March 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9700033X.

Ginsburg:1993:CAC

[GSG93] Patty Ginsburg, Scott Sterling, and Sheila Gotteherer. The Citizens’ Advisory Council as a means of mitigating environmental impacts of terminal and tanker operations. Marine Policy, 17(5):404–411, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X93900569.

Gulland:1990:ISM

[Gul90] John Gulland. ICES Symposium on multi-species models relevant to management of living resources: the Hague, 2– 4 October 1989. Marine Policy, 14(4):361–362, July 1990.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090057X.

Garrod:1991:OSA

[GW91] Brian Garrod and David Whitmarsh. Overcapitalization and structural adjustment: Recent developments in UK fisheries. Marine Policy, 15(4):289–296, July 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190005V.

Gwiazda:1993:CFP

[Gwi93] Adam Gwiazda. The common fisheries policy: Economic aspects. Marine Policy, 17(4):251–255, July 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390081D.

Gorina-Ysern:2021:MSR

[GY21] Montserrat Gorina-Ysern. Marine scientific research activities as the legal basis for intellectual property claims? Marine Policy, 22(4–5):337–357, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97000420.

Haakonsen:1992:SAF

[Haa92] Jan M. Haakonsen. Seminar on Artisanal Fishermen’s Migrations, Cotonou, Benin, 3–5 December 1991. Marine Policy, 16 (3):237–238, May 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900865.

Hall:1993:PPL

[Hal93] M. R. Hall. Problems of pollution and litter in the seas around the Scottish Isles: the Shetland perspective. Marine Policy, 17(5):418–421, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390058B.

Hannesson:1995:FHS

[Han95] Rögnvaldur Hannesson. Fishing on the high seas: cooperation or competition? Marine Policy, 19(5):371–377, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9500008T.

Hatcher:1997:POD

[Hat97] Aaron C. Hatcher. Producers’ organizations and devolved fisheries management in the United Kingdom: Collective and individual quota systems. Marine Policy, 21(6):519–533, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000195.

Hawkins:1991:FRH

[Haw91] J. E. Hawkins. Freight rate hedging in the Pacific Rim. Marine Policy, 15(6):459–461, November 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190054F.

Hayashi:1996:AHO

[Hay96] Moritaka Hayashi. Archaeological and historical objects under the United Nations Convention on the Law of the Sea. Marine Policy, 20(4):291–296, July 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000139.

Hinds:1998:CRO

[HB98] L. O. Hinds and G. B. Bacon. CIDA regional ocean initiative workshop: Lessons learned. Marine Policy, 22(6):539– 543, November 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000293.

Hinds:1998:WCR

[HBMV98] Lennox O. Hinds, G. B. Bacon, Carlyle Mitchell, and Gary Vernon. A workshop on CIDA’s regional oceans initiatives. Marine Policy, 22(6):437–440, November 1998. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000311.

Song:1990:SSP

[hBS90] Yann huei Billy Song. Shipping and shipbuilding policies in PR China. Marine Policy, 14(1):53–70, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090037R.

Hempel:1998:DKP

[Hem98] Gotthilf Hempel. Dr N. K. Panikkar: learning research by doing. Marine Policy, 22(3):247–254, May 1998. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000116.

Herbert:1995:FRG

[Her95]  Glen J. Herbert.  Fisheries relations in the Gulf of

Maine: Implications of an arbitrated maritime boundary. Marine Policy, 19(4):301–316, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500021W.

Hey:1991:PAI

[Hey91] Ellen Hey. The precautionary approach: Implications of the revision of the Oslo and Paris Conventions. Marine Policy, 15 (4):244–254, July 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190002S.

Hey:1992:BRF

[Hey92] Ellen Hey. Book review: The forum fisheries agency achievements, challenges and prospects: edited by Richard Herr, Institute of Pacific Studies, University of the South Pacific, 1990, xvi + 446 pp. US$24 (available from: the Forum Fisheries Agency, Information Officer, PO Box 629, Honiara, Solomon Islands). Marine Policy, 16(3):235–237, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X92900854.

Hilborn:1996:CPF

[HG96] Ray Hilborn and Don Gunderson. Chaos and paradigms for fisheries management. Marine Policy, 20(1):87–89, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9681921X.

Holland:1999:DFV

[HGG99] Dan Holland, Eyjolfur Gudmundsson, and John Gates. Do fishing vessel buyback programs work: a survey of the evidence. Marine Policy, 23(1):47–69, January 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000165.

Healey:1998:PFI

[HH98] M. C. Healey and T. Hennessey. The paradox of fairness: the impact of escalating complexity on fishery management. Marine Policy, 22(2):109–118, March 1998. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000377.

Hilden:1997:CBF

[Hil97] Mikael Hildén. Conflicts between fisheries and seabirds — Management options using decision analysis. Marine Policy, 21(2):143–153, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X96000528.

Hinds:1992:WMF

[Hin92] Lennox Hinds. World marine fisheries: Management and development problems. Marine Policy, 16(5):394–403, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290007C.

Hoagland:1991:EAM

[HKP91] Porter Hoagland and Hauke L. Kite-Powell. European advanced marine electronic instrumentation: a US perspective. Marine Policy, 15(6):431–454, November 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190052D.

Hakapaa:1999:IPP

[HM99] K. Hakapää and E. J. Molenaar. Innocent passage — past and present. Marine Policy, 23(2):131–145, March 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000323.

Hofman:1993:CCA

[Hof93] Robert J. Hofman. Convention for the conservation of Antarctic marine living resources. Marine Policy, 17(6):534–536, November 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390016V.

Holliday:1993:PSM

[Hol93] Steve Holliday. Port and shipping management practice: role of the oil industry and charterers. Marine Policy, 17(5):387– 390, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X93900525.

Hong:1995:MPR

[Hon95] Seoung-Yong Hong. Marine policy in the Republic of Korea. Marine Policy, 19(2):97–113, March 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400009H.

Honneland:1998:ARF

[Høn98] Geir B. Hønneland. Autonomy and regionalisation in the fisheries management of northwestern Russia. Marine Policy, 22(1):57–65, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X97000298.

Hinds:1998:OMA

[HP98] Lennox O. Hinds and Renald Piché. Oceans, Marine Affairs, and Fisheries (OMAF) database emphasis: South– South network module. Marine Policy, 22(6):537–538, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000281.

Halliday:1992:DMM

[HPB92] R. G. Halliday, F. G. Peacock, and D. L. Burke. Development of management measures for the groundfish fishery in Atlantic Canada: a case study of the Nova Scotia inshore fleet. Marine Policy, 16(6):411–426, November 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X92900706.

Herrick:1997:AFE

[HRS97] Samuel F. Herrick, Byron Rader, and Dale Squires. Access fees and economic benefits in the Western Pacific United States purse seine tuna fishery. Marine Policy, 21(1):83–96, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X96000486.

Hearns:1996:MPC

[HS96] Glen S. Hearns and William G. Stormont. Managing Potential Conflicts in the South China Sea. Marine Policy, 20(2):177– 181, March 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X00800020.

Song:1999:CMT

[hS99] Yann huei Song. China’s missile tests in the Taiwan Strait: Relevant international law questions. Marine Policy, 23(1): 81–102, January 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000190.

Huebert:1995:PVT

[Hue95] Rob Huebert. Polar vision or tunnel vision the making of Canadian Arctic waters policy: the making of Canadian Arctic waters policy. Marine Policy, 19(4):343–363, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9500011T.

Hughes:1992:EMP

[Hug92] Adrian Hughes. Economic measures to protect the environment. Marine Policy, 16(1):36–42, January 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290053R.

Hughes:1994:OOG

[Hug94] Harold Hughes. Offshore oil and gas: the UK’s second most valuable resource. Marine Policy, 18(6):464–471, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900671.

Hunt:1997:MSP

[Hun97] Colin Hunt. Management of the South Pacific tuna fishery. Marine Policy, 21(2):155–171, March 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X9600053X.

Hunt:1999:FFT

[Hun99] Colin Hunt. Fiji’s fisheries: their contribution to development and their future. Marine Policy, 23(6):571–585, November 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X9800044X.

Hutchinson:1996:TUC

[Hut96] Gillian Hutchinson. Threats to underwater cultural heritage. The problems of unprotected archaeological and historic sites, wrecks and objects found at sea. Marine Policy, 20(4):287– 290, July 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000176.

Ijlstra:1990:BRA

[Ijl90] Ton Ijlstra. Book review: Annual review of ocean affairs: Law and policy: Main Documents 1985–1987, Vols 1 and 2, UNIFO Publishers, Sarasota, FL, 1989, 926 pp + index. Marine Policy, 14(1):97–98, January 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090046T.

Istra:1993:NSP

[Ist93] T. I. J. Istra. North Sea pollution: Vessel-source pollution, environmental management and the establishment of EEZs. Marine Policy, 17(2):130–135, March 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390027Z.

Jakobsen:1999:DLC

[Jak99] Stig-Erik Jakobsen. Development of local capitalism in Norwegian fish farming. Marine Policy, 23(2):117–130, March 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X9800030X.

Janis:1992:LST

[Jan92] Mark W. Janis. The law of the sea tribunal and the ICJ: Some notions about utility. Marine Policy, 16(2):102–106, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290032K.

Juda:1990:PCO

[JB90] Lawrence Juda and R. H. Burroughs. The prospects for comprehensive ocean management. Marine Policy, 14(1):23– 35, January 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090035P.

Jorgensen-Dahl:1992:BRC

[JD92] Arnfinn Jorgensen-Dahl. Book review: The challenge of Arctic shipping: Science, environmental assessment, and human values, edited by David L. VanderZwaag and Cynthia Lamson, McGill-Queen’s university press, Montreal and Kingston, Canada, 1990, 282 pp. Marine Policy, 16(2):142, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290040V.

Jistra:1995:FES

[Jis95] Ton I. Jistra. The First European Seas at Risk Conference, Copenhagen, 26–28 October 1994. Marine Policy, 19(5):447– 448, September 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9590001B.

Jentoft:1995:UPF

[JM95] Svein Jentoft and Bonnie McCay. User participation in fisheries management: lessons drawn from international experiences. Marine Policy, 19(3):227–246, May 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400010P.

Jentoft:2021:STF

[JMW21] Svein Jentoft, Bonnie J. McCay, and Douglas C. Wilson. Social theory and fisheries co-management. Marine Policy, 22(4–5): 423–436, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000407.

Johnson:1992:FED

[Joh92] Ian Johnson. (1) Financing environmental development through the Global environmental facility. Marine Policy, 16(1):69–71, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290068Z.

Jones:1997:GGR

[Jon97] Roy Jones. Geostrategy and geotactics: Recent naval port developments at Garden Island, Western Australia. Marine Policy, 21(4):363–376, July 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X97000201.

Jones:1999:MNR

[Jon99] Peter J. S. Jones. Marine nature reserves in Britain: past lessons, current status and future issues. Marine Policy, 23(4– 5):375–396, July 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000566.

Joyner:1996:MEP

[Joy96] Christopher C. Joyner. The 1991 Madrid environmental protection protocol. Contributions to marine pollution law. Marine Policy, 20(3):183–197, May 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000127.

Jusoh:1998:ACC

[Jus98] M. Mazlan Jusoh. ASEAN–Canada Cooperative Programme on Marine Science — phase II (CPMS-II): All phases of the project cycle. Marine Policy, 22(6):493–503, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000244.

Kusuma-Atmadja:1991:SRI

[KA91] Mochtar Kusuma-Atmadja. Sovereign rights over Indonesian natural resources: an archipelagic concept of rational and sustainable resource management. Marine Policy, 15(6):383–412, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190050L.

Kuperan:1994:SSC

[KA94] K. Kuperan and Nik Mustapha Raja Abdullah. Small-scale coastal fisheries and co-management. Marine Policy, 18(4): 306–313, July 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900450.

Kusuma-Atmadja:1996:LIA

[KAP96] Mochtar Kusuma-Atmadja and Tommy H. Purwaka. Legal and institutional aspects of coastal zone management in Indonesia. Marine Policy, 20(1):63–86, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X95000344.

Kaplan:1999:SGC

[Kap99] Ilene M. Kaplan. Suspicion, growth and co-management in the commercial fishing industry: the financial settlers of New Bedford. Marine Policy, 23(3):227–241, May 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000414.

Kaplan:2021:RCN

[Kap21] Ilene M. Kaplan. Regulation and compliance in the New England Conch Fishery: a case for co-management. Marine Policy, 22(4–5):327–335, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000487.

Kariotis:1990:CGE

[Kar90] Theodore C. Kariotis. The case for a Greek exclusive economic zone in the Aegean Sea. Marine Policy, 14(1):3–14, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090033N.

Karagiannakos:1996:TAC

[Kar96] A. Karagiannakos. Total Allowable Catch (TAC) and quota management system in the European Union. Marine Policy, 20 (3):235–248, May 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96869867.

Kasoulides:1990:U

[Kas90] George C. Kasoulides. Update. Marine Policy, 14(1):84– 85, January 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090040X.

Kbaier:1994:ILA

[Kba94] Rouchdy Kbaier. Involvement of local authorities in the protection of coastal areas: the city of Dakar. Marine Policy, 18(2): 199–208, March 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900302.

Kebe:1990:WDA

[Keb90] Moctar Kebe. Waste disposal in Africa. Marine Policy, 14(3): 251–253, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900834.

Kent:1992:APS

[Ken92] Paul E. Kent. Asia’s port sector: the need to institutionalize training. Marine Policy, 16(5):371–378, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290005A.

Kendra:1997:SDV

[Ken97] James M. Kendra. Seaport development versus environmental preservation: the case of Sears Island, Maine, USA. Marine Policy, 21(5):409–424, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97000043.

Kesteven:1995:CCC

[Kes95] G. L. Kesteven. Chaos, complexity and community management. Marine Policy, 19(3):247–248, May 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9590824E.

Kesteven:1997:FCM

[Kes97a] G. L. Kesteven. Fisheries co-management: a comparative analysis. Marine Policy, 21(2):195–196, March 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97832253.

Kesteven:1997:MRR

[Kes97b] G. L. Kesteven. MSY revisited. A realistic approach to fisheries management and administration. Marine Policy, 21(1): 73–82, January 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X96000425.

Ketkar:1995:PMR

[Ket95] Kusum W. Ketkar. Protection of marine resources: the US Oil Pollution Act of 1990 and the future of the maritime industry. Marine Policy, 19(5):391–400, September 1995.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X95000104.

Keyuan:1999:IME

[Key99] Zou Keyuan. Implementing marine environmental protection law in China: progress, problems and prospects. Marine Policy, 23(3):207–225, May 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X98000438.

Kirkegaard:2021:PIM

[KG21] Ian R. Kirkegaard and Donald F. Gartside. Performance indicators for management of marine recreational fisheries. Marine Policy, 22(4–5):413–422, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000037.

Kimball:1990:SSP

[Kim90] Lee Kimball. The Seventh Session of the Preparatory Commission for the 1982 UN Convention on the Law of the Sea: Kingston, Jamaica, 27 February–23 March 1989, and UN Headquarters, New York, USA, 14 August–1 September 1989. Marine Policy, 14(1):88–93, January 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090042P.

Kim:1993:BRC

[Kim93a] Hyun-Soo Kim. Book review: China’s practice in the law of the sea: by Jeanette Greenfield, Clarendon Press, Oxford, 1992, xv + 321pp. Marine Policy, 17(2):146–147, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900352.

Kimbal:1993:UOA

[Kim93b] Lee A. Kimbal. UNCED and the oceans agenda: the process forward. Marine Policy, 17(6):491–500, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390012R.

Kimball:1994:LE

[Kim94] Lee A. Kimball. Letter to the editor. Marine Policy, 18(5):436– 437, September 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900396.

King:1994:BSM

[Kin94] John King. Beyond the shoreline: the marine estate. Marine Policy, 18(6):457–463, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900663.

King:1995:ICS

[Kin95] John King. An inquiry into the causes of shipwrecks: Its implications for the prevention of pollution. Marine Policy, 19(6):469–475, November 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X95000296.

King:1999:NDS

[Kin99] John King. New directions in shipbuilding policy. Marine Policy, 23(3):191–205, May 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X98000426.

Kittichaisaree:1990:DOL

[Kit90] Kriangsak Kittichaisaree. Development of ocean law, policy and management in Thailand. Marine Policy, 14(4):315– 323, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090053T.

Koroleva:1991:SMP

[KK91] N. D. Koroleva and V. A. Kiselev. Soviet marine pollution legislation: Prevention of pollution from ships and the LOS Convention. Marine Policy, 15(1):49–54, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190043B.

Kolossovskiy:1993:PUC

[Kol93] Igor K. Kolossovskiy. Prospects for universality of the UN convention on the law of the sea. Marine Policy, 17(1):4– 10, January 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390003L.

Koroleva:1990:RPE

[Kor90] N. D. Koroleva. The right of pursuit from the exclusive economic zone. Marine Policy, 14(2):137–141, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90901006.

Koroleva:1992:EDR

[Kor92] N. D. Koroleva. Ecological damage: Responsibility for pollution of the marine environment. Marine Policy, 16(2):86– 89, March 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290029O.

Kite-Powell:1995:ETS

[KPG95] Hauke L. Kite-Powell and Arthur G. Gaines. Evaluation of a technological standard for electronic chart systems. Marine Policy, 19(3):185–197, May 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400004C.

Kriwoken:1991:GBR

[Kri91] Lorne K. Kriwoken. Great Barrier Reef Marine Park: Intergovernmental relations. Marine Policy, 15(5):349–362, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9190090X.

Kuronuma:1993:IMI

[KT93] Yoshihiro Kuronuma and Clement A. Tisdell. Institutional management of an international mixed good: the IWC and socially optimal whale harvests. Marine Policy, 17(4):235– 250, July 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390080M.

Kullenberg:1994:TEU

[Kul94] Gunnar Kullenberg. Tropical environments: Understanding and predicting change. Marine Policy, 18(2):138–147, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900191.

Kullenberg:1998:CBM

[Kul98] Gunnar Kullenberg. Capacity building in marine research and ocean observations: a perspective on why and how. Marine Policy, 22(3):185–195, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000050.

Kutting:1994:MPI

[Küt94] Gabriela Kütting. Mediterranean pollution: International cooperation and the control of pollution from land-based sources. Marine Policy, 18(3):233–247, May 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900027.

Kolodkin:1990:LRS

[KV90] A. L. Kolodkin and M. E. Volosov. The legal regime of the Soviet Arctic: Major issues. Marine Policy, 14(2):158– 168, March 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090103X.

Kimball:1991:ESP

[KW91] Lee Kimball and Dolores M. Wesson. The eighth session of the Preparatory Commission for the 1982 UN Convention on the Law of the Sea: Kingston, Jamaica, 5–30 March 1990 and New York, USA, 13–31 August 1990. Marine Policy, 15(3): 220–223, May 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190065J.

Kwiatkowska:1990:PLS

[Kwi90a] Barbara Kwiatkowska. 2nd Penataran on the law of the sea: Bandung, Indonesia, 8–19 January 1990. Marine Policy, 14(4):

368–370, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900605.

Kwiatkowska:1990:IMAa

[Kwi90b] Barbara Kwiatkowska. Institutional marine affairs cooperation in developing State regions: Part 1: General problems and prospects. Marine Policy, 14(5):385–398, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X90900029.

Kwiatkowska:1990:IMAb

[Kwi90c] Barbara Kwiatkowska. Institutional marine affairs cooperation in developing State regions: Part 2: the Indian Ocean and IOMAC. Marine Policy, 14(5):399–420, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090003A.

Kwiatkowska:1993:BRS

[Kwi93a] Barbara Kwiatkowska. Book review: Sea-bed energy and minerals: the international legal regime: Vol 1. The Continental Shelf, by E. D. Brown, Martinus Nijhoff, Dordrecht, Boston, London, 1992, liv + 484pp, US$186, £105. Marine Policy, 17(2):141–145, March 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390032X.

Kwiatkowska:1993:OAL

[Kwi93b] Barbara Kwiatkowska. Ocean affairs and the law of the sea in Africa: Towards the 21st century. Marine Policy, 17(1):11– 43, January 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390004M.

Lalonde:1990:RWS

[Lal90] Brice Lalonde. Reduction of waste at source. Marine Policy, 14 (3):224–227, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900786.

Langdon:1990:BRL

[Lan90a] J. B. R. L. Langdon. Book review: The law of naval warfare: a collection of agreements and documents with commentaries: edited by Professor Natalino Ronzitti, Martinus Nijhoff, Dordrecht, 1988, 888 pp. £117.00. Marine Policy, 14(1):95– 96, January 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090044R.

Langdon:1990:ETP

[Lan90b] J. B. R. L. Langdon. The extent of transit passage: Some practical anomalies. Marine Policy, 14(2):130–136, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090099D.

Larminie:1993:BIW

[Lar93] F. G. Larminie. The Braer incident: workshop presentation. Marine Policy, 17(5):483–485, September 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390070J.

Laughlin:1993:CAI

[Lau93] Thomas L. Laughlin. Chapter 17 of Agenda 21: Implementing data and information aspects. Marine Policy, 17(6):557– 560, November 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390018X.

Laughton:1994:UOW

[Lau94] Anthony S. Laughton. Using the ocean wisely: the waste problem. Marine Policy, 18(6):453–456, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900655.

Lawson:1990:BRD

[Law90a] Rowena Lawson. Book review: Development and social changes in the Pacific Islands: edited by A. D. Couper, Routledge, London, 1989, 203 pp, £30.00. Marine Policy, 14(4): 370–371, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090061U.

Lawson:1990:CMD

[Law90b] Rowena Lawson. Conference on Management and Development Strategies in South Pacific fisheries, to celebrate the Tenth Anniversary of the South Pacific Forum Fisheries Agency: Honiara, Solomon Isles, September 1989. Marine Policy, 14(5):456, September 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090010O.

Lawn:1993:AEE

[Law93a] Dan Lawn. The Alaskan experience of environmental management: Need history be a path to the future? Marine Policy, 17(5):412–417, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390057A.

Lawson:1993:BRS

[Law93b] Rowena Lawson. Book review: A study of international fisheries research: World Bank, UNDP, CEC and FAO, World Bank Policy and Research Series 19, World Bank, Washington, 1992. Marine Policy, 17(4):322–324, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390088K.

Lawson:1993:BRR

[Law93c] Rowena Lawson. Book review: Review of fisheries in OECD member countries: OECD, Paris, 1993, 347pp + 170 tables. Marine Policy, 17(4):322, July 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390087J.

Lax:1992:LDR

[Lax92] Michael D. Lax. Laytime and demurrage: Recent developments. Marine Policy, 16(2):118–126, March 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290036O.

LePensec:1992:MDA

[Le 92] Louis Le Pensec. (5) Multi-disciplinary approach. Marine Policy, 16(1):62, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9290064V.

Lee:1990:KSP

[Lee90] Tae-Woo Lee. Korean shipping policy: the role of government. Marine Policy, 14(5):421–437, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090004B.

Leinster:1993:AHE

[Lei93] Paul Leinster. Assessing human exposures to environmental pollution arising from the Braer incident. Marine Policy, 17(5):457–467, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390066C.

Levy:1993:NOP

[Lév93] Jean-Pierre Lévy. A national ocean policy: an elusive quest. Marine Policy, 17(2):75–80, March 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390023V.

Lodge:1992:MTC

[Lod92] Michael Lodge. Minimum terms and conditions of access: Responsible fisheries management measures in the South Pacific region. Marine Policy, 16(4):277–305, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290045Q.

Lodge:1998:DPA

[Lod98] Michael W. Lodge. The development of the Palau Arrangement for the management of the Western Pacific Purse Seine Fishery. Marine Policy, 22(1):1–28, January 1998.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000249.

Long:1990:EAT

[Lon90] William L. Long. Economic aspects of transport and disposal of hazardous wastes. Marine Policy, 14(3):198–204, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090073Z.

Longhurst:1991:BRL

[Lon91] Alan R. Longhurst. Book review: Large marine ecosystems: Patterns, processes and yields, edited by K. Sherman, L. M. Alexander and B. D. Gold, American Association for the Advancement of Science, Washington, DC, 242 pp. Marine Policy, 15(5):377–378, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190097U.

Lowe:1990:BRU

[Low90] A. V. Lowe. Book review: United Nations Convention on the law of the sea 1982: a commentary. Volume 2 Editor-in-Chief Myron H. Nordquist, volume editors Shabtai Rosenne and Louis B. Sohn Martinus Nijhoff, Dordrecht/Boston/ London, 1989, 497pp. Marine Policy, 14(5):459–460, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090013H.

Lowe:1994:BRP

[Low94] Vaughan Lowe. Book review: Port state jurisdiction and control: Evolution of the port state regime: George C. Kasoulides, Nijhoff 1993 298 pp £78.00 ISBN 0-7923-2281-9. Marine Policy, 18(5):438–439, September 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900418.

Labajos:1996:CIF

[LRA96] Carlos Angel Pérez Labajos, Beatriz Blanco Rojo, and Rafael Mazas Arranz. The Canary Islands fishing policy. Marine Policy, 20(6):463–474, November 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000371.

Lalwani:1999:DMI

[LS99] C. S. Lalwani and T. Stojanovic. The development of marine information systems in the UK. Marine Policy, 23(4–5): 427–438, July 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9900007X.

Ludwig:1995:BNE

[LV95] Noel A. Ludwig and Mark J. Valencia. Building North-east Asian maritime regimes: Will Japan take the lead? Marine Policy, 19(2):83–96, March 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X94000019.

loannidis:1990:IPU

[lW90] Chris loannidis and David Whitmarsh. Import penetration and UK fisheries policy. Marine Policy, 14(5):451–453, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090007E.

Miller:1991:CZT

[MA91] Marc L. Miller and Jan Auyong. Coastal zone tourism: a potent force affecting environment and society. Marine Policy, 15 (2):75–99, March 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190008Y.

Mackay:1993:WPC

[Mac93a] David W. Mackay. Water pollution control: the case for local control and accountability. Marine Policy, 17(5):422–425, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390059C.

Mackenzie:1993:SFM

[Mac93b] L. D. M. Mackenzie. Sea fisheries management: the EC position. Marine Policy, 17(5):343–346, September 1993.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900433.

Mahmoudi:1991:PWT

[Mah91] Said Mahmoudi. Passage of warships through the Strait of Hormuz. Marine Policy, 15(5):338–348, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190089T.

Mahmoudi:1997:LPP

[Mah97] Said Mahmoudi. Legal protection of the Persian Gulf’s marine environment. Marine Policy, 21(1):53–62, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X96000474.

Maiava:1998:PFC

[Mai98] Iosefa Maiava. Programme formulation case study Canada– South Pacific Ocean Development Programme (C-SPODP). Marine Policy, 22(6):455–467, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000219.

Makovskaia:1992:DCR

[Mak92]  A. A. Makovskaia.  Development of contractual rela-

tions in maritime transport: 1991 Soviet civil legislation. Marine Policy, 16(2):115–117, March 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9290035N.

Marriott:1990:TWF

[Mar90] Alan Marriott. Third World fisheries departments. Marine Policy, 14(5):453–455, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090008F.

Marti:1991:UPB

[Mar91] Bruce E. Marti. US/Pacific basin containerized foreign trade: Flag of convenience participation. Marine Policy, 15(3):193– 198, May 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190062G.

Marti:1992:PPC

[Mar92] Bruce E. Marti. Passenger perceptions of cruise itineraries: a Royal Viking Line case study. Marine Policy, 16(5):360– 370, September 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900049.

Marston:1993:SPM

[Mar93] Geoffrey Marston. St Pierre-Miquelon arbitration: Canada-France maritime delimitation award. Marine Policy, 17(3): 155–170, May 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390074D.

Marriott:1997:FIR

[Mar97] Seán P. Marriott. Fisheries institutional reform in developing countries. Marine Policy, 21(5):435–444, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000158.

Matthiasson:1992:PDR

[Mat92] Thorolfur Matthíasson. Principles for distribution of rent from a ‘commons’. Marine Policy, 16(3):210–231, May 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290082Z.

McCarter:1999:SII

[McC99] Capt Philip McCarter. STCW ’95: implementation issues: What is the pass mark? Marine Policy, 23(1):11–24, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000468.

McElroy:1991:JSP

[McE91a] J. K. McElroy. The Java Sea purse seine fishery: a modernday ‘tragedy of the commons’? Marine Policy, 15(4):255– 271, July 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190003T.

McElroy:1991:SJS

[McE91b] J. K. McElroy. Seminar on the Java Sea pelagic fishery assessment project: Central Research Institute for Fisheries, Slipi, Jakarta, Indonesia, 21 March 1991. Marine Policy, 15(6):462– 464, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190055G.

McElroy:1991:BRM

[McE91c] Seamus McElroy. Book review: The marketing of marine products from the South Pacific: edited by P. W. Philipson, Institute of Pacific Studies of the University of the South Pacific, 1989, 307 pp. Marine Policy, 15(2):136–137, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X91900132.

McElroy:1992:FRM

[McE92] Seamus McElroy. Failure to resolve marine boundary disputes raises tensions in SE Asia. Marine Policy, 16(6):488– 490, November 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290075Z.

McElroy:1993:TAF

[McE93a] Seamus McElroy. Third Asian fisheries forum: Singapore 26–30 October 1992. Marine Policy, 17(2):140–141, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390031W.

McElroy:1993:WBF

[McE93b] Seamus McElroy. World Bank fisheries report. Marine Policy, 17(2):136–137, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X93900282.

McIntyre:1993:CPS

[McI93] A. D. McIntyre. Control of pollution of the sea. Marine Policy, 17(5):394–398, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X93900547.

McIntyre:1995:EMO

[McI95] A. D. McIntyre. Environmental monitoring of the oceans. Marine Policy, 19(6):497–502, November 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X95000322.

McConville:1997:EIU

[MG97] James McConville and David Glen. The employment implications of the United Kingdom’s merchant fleet’s decline. Marine Policy, 21(3):267–276, May 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000092.

Mirovitskaya:1992:FET

[MH92] Natalia S. Mirovitskaya and J. Christopher Haney. Fisheries exploitation as a threat to environmental security: the North Pacific Ocean. Marine Policy, 16(4):243–258, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290043O.

Mitchell:1997:FMG

[Mit97] Carlyle L. Mitchell. Fisheries management in the Grand Banks, 1980–1992 and the straddling stock issue. Marine Policy, 21(1):97–109, January 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X96000462.

Mitchell:2021:SOD

[Mit21] C. L. Mitchell. Sustainable oceans development: the Canadian approach. Marine Policy, 22(4–5):393–412, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000025.

Mizukami:1991:FPS

[Miz91] Chiyuki Mizukami. Fisheries problems in the South Pacific region. Marine Policy, 15(2):111–121, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X91900109.

Mohamed:1991:NMM

[Moh91] Mohd Ibrahim Hj Mohamed. National management of Malaysian fisheries. Marine Policy, 15(1):2–14, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190038D.

Morgan:1995:OFQ

[Mor95] G. R. Morgan. Optimal fisheries quota allocation under a transferable quota (TQ) management system. Marine Policy, 19(5):379–390, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500005Q.

Morris:1999:DRU

[Mor99] Anthony J. Morris. Discharge regulation of the UK nuclear industry. Marine Policy, 23(4–5):359–373, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000499.

Marcadon:1997:GNP

[MP97] Jacques Marcadon and David Pinder. Geostrategy, naval ports and local economic development: Brest in an era of strategic reappraisal. Marine Policy, 21(4):391–408, July 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000213.

Moncrieff:1993:ESI

[MS93] James Moncrieff and Charles H. Simpson. Economic and social impacts. Marine Policy, 17(5):469–472, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390067D.

McConville:1996:RAD

[MT96] J. McConville and K. W. Timmermann. Regulation and age discrimination in the dry-bulk cargo trades. Marine Policy, 20(5):389–395, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X96000310.

McElroy:1992:SPL

[MU92] James Kevin McElroy and Jacobus C. B. Uktolseja. Skipjack pole-and-line operations in east Indonesia: a comparative analysis of catch performance. Marine Policy, 16(6):451– 462, November 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290073X.

Murray:1994:MEC

[Mur94] L. E. Murray. Maritime enforcement: the Canadian federal government’s marine fleets and the Navy’s mission. Marine Policy, 18(6):521–529, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900752.

Muttagi:1994:CPB

[Mut94] P. K. Muttagi. Controlling pollution in Bombay and Calcutta. Marine Policy, 18(2):209–216, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900310.

Nadelson:1992:EEZ

[Nad92] Robert Nadelson. The exclusive economic zone: State claims and the LOS convention. Marine Policy, 16(6):463–487, November 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290074Y.

Nauen:1994:WAR

[Nau94] Cornelia E. Nauen. Why is aquatic resource systems research needed?: a response to Rowena Lawson’s review of World Bank, UNDP, CEC and FAO, a Study of International Fisheries Research, Marine Policy, Vol 17, No 4, pp 322–324. Marine Policy, 18(3):284–286, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9490006X.

Nauen:1995:RPP

[Nau95] Comelia E. Nauen. Regional Postharvest Programme gets into gear in West Africa. Marine Policy, 19(5):445–446, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9598049L.

Neiland:1991:SFF

[Nei91] Arthur E. Neiland. Scottish fish farming conference and exhibition: Glasgow, 29–30 January 1991. Marine Policy, 15(3): 226–228, May 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190067L.

Neiland:1992:ACE

[Nei92] Arthur E. Neiland. 4th Annual Conference of the European Association of Fisheries Economists (EAFE): Salerno, Italy, 22–24 April 1992. Marine Policy, 16(5):404–406, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290008D.

Nicholson:1990:EN

[Nic90] Jennifer Nicholson. Editor’s note. Marine Policy, 14(1): 2, January 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090032M.

Nielsen:1992:SPD

[Nie92] Jesper Raakjær Nielsen. Structural problems in the Danish fishing industry: Institutional and socio-economic factors as barriers to adjustment. Marine Policy, 16(5):349–359, September 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X92900038.

Nielsen:1994:PFM

[Nie94] Jesper Raakjær Nielsen. Participation in fishery management policy making: National and EC regulation of Danish fishermen. Marine Policy, 18(1):29–40, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900876.

Nijpels:1990:SDR

[Nij90] Ed H. T. M. Nijpels. Sustainable development and risk management. Marine Policy, 14(3):219–223, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900775.

Noelke:1993:HAA

[Noe93] Eberhard Noelke. The harbour authority approach to the environment: role of the European Harbour Masters Association. Marine Policy, 17(5):380–386, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900514.

Nollkaemper:1991:DSB

[Nol91] André Nollkaemper. Deep sea-bed mining and the protection of the marine environment. Marine Policy, 15(1):55– 66, January 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190044C.

Nollkaemper:1993:APS

[Nol93] André Nollkaemper. Agenda 21 and prevention of seabased marine pollution: a spurious relationship? Marine Policy, 17(6):537–556, November 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9390017W.

Norrena:1994:SCW

[Nor94] E. J. Norrena. Stewardship of coastal waters and protected spaces: Canada’s approach. Marine Policy, 18(2):153–160, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900213.

Nowak:1992:SFH

[Now92] Robert Nowak. Scotia–Fundy herring purse seine fishery: Some economic considerations. Marine Policy, 16(3):183– 196, May 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900809.

Nowell:1993:HEC

[Now93] David Nowell. Humber estuary conference: Hull, UK, 5 November 1992. Marine Policy, 17(2):137–138, March 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900293.

Nielsen:1999:FAW

[NR99] D. Nielsen and S. Roberts. Fatalities among the world’s merchant seafarers (1990–1994). Marine Policy, 23(1):71–80, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000177.

Nakken:1996:RPO

[NSS96] Odd Nakken, Per Sandberg, and Stein Ivar Steinshamn. Reference points for optimal fish stock management: a lesson to be learned from the Northeast Arctic cod stock. Marine Policy, 20(6):447–462, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X96000334.

Nielsen:1997:FOF

[NV97] Jesper Raakjær Nielsen and Tomas Vedsmand. fishermen’s organisations in fisheries management. Perspectives for fisheries co-management based on Danish fisheries. Marine Policy, 21 (3):277–288, May 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000109.

ofArdbrecknish:1993:SFM

[oA93] Lord Mackay of Ardbrecknish. Sea fisheries management: workshop presentation. Marine Policy, 17(5):352–353, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900466.

G:1990:IPA

[oCKG90] Lord Callaghan of Cardiff K. G. Introductions: President’s address. Marine Policy, 14(3):184–187, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900694.

OKeefe:1996:PUC

[O’K96] Patrick J. O’Keefe. Protecting the underwater cultural heritage. The International Law Association Draft Convention. Marine Policy, 20(4):297–307, July 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000188.

Okemwa:1998:IOC

[Oke98] Ezekiel Okemwa. The Intergovernmental Oceanographic Commission of UNESCO and regional capacity building. Marine Policy, 22(3):197–207, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000062.

Okill:1993:EIW

[Oki93] David Okill. Environmental impacts on wildlife, land, food and soil. Marine Policy, 17(5):449–452, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390064A.

Omar:1992:MFP

[ONAK92] Ishak Hj Omar, Kusairi Mohd Noh, Nik Mustapha Raja Abdullah, and K. Kuperan. Malaysian fisheries policy: Search for new grounds. Marine Policy, 16(6):438–450, November 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290072W.

OReilly:1995:CCA

[O’R95] Lennox O’Reilly. Crisis in Canada’s Atlantic sea fisheries. Marine Policy, 19(4):271–283, July 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400013I.

Programme:1992:LBP

[oUotCEP92] U. N. E. P. Regional Co ordinating Unit of the Caribbean Environment Programme. Land-based pollution protocol to Cartagena convention. Marine Policy, 16(1):43–49, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9290054S.

Owen:1999:EMO

[Owe99] Jeanette Owen. The environmental management of oil tanker routes in UK waters. Marine Policy, 23(4–5):289–306, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000463.

Otto-Zimmermann:1994:LIA

[OZ94] Konrad Otto-Zimmermann. Local implementation of Agenda 21. Marine Policy, 18(2):112–115, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900159.

Parker:1993:PSM

[Par93] C. J. Parker. Port and shipping management: workshop presentation. Marine Policy, 17(5):391–393, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900536.

Payne:1994:GFW

[Pay94a] Robert L. Payne. A glance into the future of the world canned tuna trade. Marine Policy, 18(5):407–421, September 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900361.

Payoyo:1994:IIC

[Pay94b] Peter Bautista Payoyo. Implementation of international conventions through port state control: an assessment. Marine Policy, 18(5):379–392, September 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900345.

Pomeroy:1997:TTR

[PB97] Robert S. Pomeroy and Fikret Berkes. Two to tango: the role of government in fisheries co-management. Marine Policy, 21(5):465–480, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97000171.

Pomeroy:1997:CBC

[PC97] Robert S. Pomeroy and Melvin B. Carlos. Communitybased coastal resource management in the Philippines: a review and evaluation of programs and projects, 1984– 1994. Marine Policy, 21(5):445–464, September 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X9700016X.

Pernetta:1998:ORG

[Per98] John C. Pernetta. An overview of the role of the Global Environment Facility in international waters with reference to marine capacity building. Marine Policy, 22(3):235–246, May 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000104.

Pettit:1999:SAC

[Pet99] S. J. Pettit. The statutory approach to coastal defence in England and Wales. Marine Policy, 23(4–5):465–477, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000529.

Phillips:1990:IWC

[Phi90] Cassandra Phillips. International Whaling Commission 41st annual meeting: San Diego, CA, 12–16 June 1989. Marine Policy, 14(1):93–95, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9090043Q.

Pickering:1994:LCA

[Pic94a] Helen Pickering. Legislative changes after Piper Alpha: Royal society, London, UK, 1 March 1994. Marine Policy, 18(5): 435, September 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900388.

Pickering:1994:PCZ

[Pic94b] Helen Pickering. Practical coastal zone management: Alternatives and strategies. Marine Policy, 18(5):393–406, September 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900353.

Pickering:1996:ARB

[Pic96] Helen Pickering. Artificial reefs of bulk waste materials: a scientific and legal review of the suitability of using the cement stabilised by-products of coal-fired power stations. Marine Policy, 20(6):483–497, November 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X9600036X.

Pickering:1999:NEO

[Pic99] H. Pickering. A new era for the offshore oil and gas industry on the UKCS. Marine Policy, 23(4–5):329–346, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98900471.

Plant:1990:ILA

[Pla90] G. Plant. International legal aspects of vessel traffic services. Marine Policy, 14(1):71–81, January 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090038S.

Plant:1995:ELV

[Pla95] Glen Plant. A European lawyer’s view of the Government response to the Donaldson Report. Marine Policy, 19(6):453– 467, November 1995. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X95000355.

Plant:1996:NRT

[Pla96] G. Plant. Navigation regime in the Turkish straits for merchant ships in peacetime: Safety, environmental protection and high politics. Marine Policy, 20(1):15–27, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9500009U.

Parsons:1996:CCC

[PM96] L. Scott Parsons and Jean-Jacques Maguire. Comments on chaos, complexity and community management of fisheries. Marine Policy, 20(2):175–176, March 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X95000436.

Polacheck:1990:APN

[Pol90] Tom Polacheck. Another perspective on the need for international skipjack and yellowfin tuna management. Marine Policy, 14(6):526–529, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090021I.

Pomeroy:1991:SSF

[Pom91] Robert S. Pomeroy. Small-scale fisheries management and development: Towards a community-based approach. Marine Policy, 15(1):39–48, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9190042A.

Pontecorvo:1991:WFS

[Pon91] G. Pontecorvo. Workshop on fisheries statistics: Woods Hole Oceanographic Institution, 21–23 October 1990. Marine Policy, 15(5):372–373, September 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9190093Q.

Potier:1994:CEC

[Pot94] Michel Potier. Cost effectiveness in coastal zone management: OECD experience. Marine Policy, 18(2):121–126, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900175.

Potts:1999:NSA

[Pot99] J. S. Potts. The non-statutory approach to coastal defence in England and Wales: Coastal Defence Groups and Shoreline Management Plans. Marine Policy, 23(4–5):479–500, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X98000530.

Pomeroy:1995:ITF

[PP95] Robert S. Pomeroy and Michael D. Pido. Initiatives towards fisheries co-management in the Philippines: the case of San Miguel Bay. Marine Policy, 19(3):213–226, May 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400008G.

Poggie:1995:PVS

[PPJ95] J. Poggie, R. Pollnac, and S. Jones. Perceptions of vessel safety regulations: a southern New England fishery. Marine Policy, 19(5):411–418, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500015X.

Prescott:1990:BRR

[Pre90] Victor Prescott. Book review: Relevant circumstances and maritime delimitation: by Malcolm D. Evans, Oxford Monographs in International Law, Clarendon Press, Oxford, UK, 1989, 257pp. £35.00. Marine Policy, 14(6):538, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090026N.

Prescot:1991:BRO

[Pre91] Victor Prescot. Book review: Ocean changes in global change: Introductory Geographical Analysis, by Adalberto Vallega, Pubblicazione 44 Dell’Instituto di Scienze Geografiche, universita Degli Studi di Genova, 1990, 136pp. Marine Policy, 15(6): 465, November 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190057I.

Pearse:1992:HRU

[PW92] Peter H. Pearse and Carl J. Walters. Harvesting regulation under quota management systems for ocean fisheries: Decision making in the face of natural variability, weak information, risks and conflicting incentives. Marine Policy, 16(3): 167–182, May 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900795.

Queirolo:1997:NEC

[QJZ97] Lewis E. Queirolo, Richard S. Johnston, and Zhengkun Zhang. The nature and evolution of cooperative fishing arrangements in extended jurisdiction zones. Marine Policy, 21(3):255– 266, May 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000079.

Ramsden:1990:LAE

[Ram90] Edward Ramsden. Local authority experience in the UK. Marine Policy, 14(3):286–289, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X90900915.

Rodgers:1999:DRM

[RB99] P. Rodgers and P. Bertram. Dispersion of revenue in mixed fisheries. Marine Policy, 23(1):37–46, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000153.

Redgwell:1990:CDEa

[Red90a] Catherine Redgwell. Current developments in the EC. Marine Policy, 14(1):82–83, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9090039T.

Redgwell:1990:CDEb

[Red90b] Catherine Redgwell. Current developments in the EC. Marine Policy, 14(5):449–450, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090006D.

Redgwell:1990:CDEc

[Red90c] Catherine Redgwell. Current developments in the EC. Marine Policy, 14(6):530–532, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090022J.

Redgwell:1990:GSL

[Red90d] Catherine Redgwell. The greening of salvage law. Marine Policy, 14(2):142–150, March 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9090101V.

Redgwell:1990:OPC

[Red90e] Catherine Redgwell. ‘Oil pollution — claims & liability’, organized by IBC legal studies & services: London, 22 November 1989. Marine Policy, 14(4):366–368, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090059Z.

Redgwell:1991:CDE

[Red91a] Catherine Redgwell. Current developments in the EC. Marine Policy, 15(1):68–69, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9190046E.

Redgwell:1991:CDU

[Red91b] Catherine Redgwell. Current developments: UK practice. Marine Policy, 15(5):370–371, September 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190092P.

Redgwell:1991:OPE

[Red91c] Catherine Redgwell. ‘Oil pollution: Environmental and commercial concerns including a USA perspective’: organized by IBC Legal Studies and Services Limited London, 8 November 1990. Marine Policy, 15(3):223–226, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190066K.

Redgwell:1992:COP

[Red92] Catherine Redgwell. Compensation for oil pollution damage: Quantifying environmental harm. Marine Policy, 16(2):90– 98, March 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290030S.

Reid:2021:MIB

[Rei21] Chris Reid. Managing innovation in the British Herring Fishery: the role of the Herring Industry Board 1945– 1977. Marine Policy, 22(4–5):281–295, ???? 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000475.

Remond-Gouilloud:1990:ILC

[RG90] Martine Remond-Gouilloud. Insurance, liability and compensation. Marine Policy, 14(3):236–242, May 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090080B.

Rothwell:1996:FIP

[RH96] Donald R. Rothwell and Marcus Haward. Federal and international perspectives on Australia’s maritime claims. Marine Policy, 20(1):29–46, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X95000274.

Ridolfi:1997:GNP

[Rid97] Giovanni Ridolfi. Geostrategy and naval power in the Mediterranean: Changing scenarios, continuing conflict. Marine Policy, 21(4):377–390, July 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X97000225.

Rimmer:1997:UWP

[Rim97] Peter J. Rimmer. US Western Pacific geostrategy: Subic Bay before and after withdrawal. Marine Policy, 21(4):325– 344, July 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000122.

Ritchie:1993:EIB

[Rit93] William Ritchie. Environmental impacts of the Braer oil spill and development of a strategy for the monitoring of change and recovery. Marine Policy, 17(5):434–440, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900628.

Rothwell:1994:LSP

[RK94] Donald R. Rothwell and Stuart Kaye. Law of the sea and the polar regions: Reconsidering the traditional norms. Marine Policy, 18(1):41–58, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X94900884.

Roginko:1992:EME

[RL92] Alexei Yu. Roginko and Matthew J. LaMourie. Emerging marine environmental protection strategies for the Arctic. Marine Policy, 16(4):259–276, July 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9290044P.

Roach:1996:SWM

[Roa96] J. Ashley Roach. Sunken warships and military aircraft. Marine Policy, 20(4):351–354, July 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X96000164.

Robson:1993:SFM

[Rob93] Godfrey Robson. Sea fisheries management: the UK position. Marine Policy, 17(5):347–349, September 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X93900444.

Rossi:1990:SFA

[Ros90] Sir Hugh Rossi. Summary and future action. Marine Policy, 14 (3):290–294, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900926.

Rothwell:1990:ISI

[Rot90] Donald R. Rothwell. The Indonesian straits incident: Transit or archipelagic sea lanes passage? Marine Policy, 14(6):491– 506, November 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090019N.

Rothwell:1992:CSS

[Rot92] Donald R. Rothwell. Coastal State sovereignty and innocent passage: the voyage of the Lusitania Expresso. Marine Policy, 16(6):427–437, November 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9290071V.

Snoussi:1998:MCB

[SA98] Maria Snoussi and Larry Awosika. Marine capacity building in North and West Africa. Marine Policy, 22(3):209– 215, May 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000074.

Saharuddin:1995:DMM

[Sah95] Abdul Hamid Saharuddin. Development and management of Malaysian marine fisheries: Technical conservation measures. Marine Policy, 19(2):115–126, March 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400011G.

Saul:1998:CFR

[Sau98] Hugh Saul. CARICOM Fisheries Resource Assessment and Management Programme (CFRAMP): Project monitoring and evaluation. Marine Policy, 22(6):477–491, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000232.

Side:1993:CCA

[SBH93] J. Side, M. Baine, and K. Hayes. Current controls for abandonment and disposal of offshore installations at sea. Marine Policy, 17(5):354–362, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X93900477.

Squires:1998:ITQ

[SCC^{+}98] Dale Squires, Harry Campbell, Stephen Cunningham, Christopher Dewees, R. Quentin Grafton, Samuel F. Herrick, James Kirkley, Sean Pascoe, Kjell Salvanes, Bruce Shallard, Bruce Turris, and Niels Vestergaard. Individual transferable quotas in multispecies fisheries. Marine Policy, 22(2):135–159, March 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000390.

Schrank:1995:EFJ

[Sch95] William E. Schrank. Extended fisheries jurisdiction: origins of the current crisis in Atlantic Canada’s fisheries. Marine Policy, 19(4):285–299, July 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9400003B.

Schug:1996:IMB

[Sch96] Donald M. Schug. International maritime boundaries and indigenous people. The case of the Torres Strait. Marine Policy, 20(3):209–222, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X95000469.

Schalk:1998:MMN

[Sch98a] Peter H. Schalk. Management of marine natural resources through by biodiversity informatics. Marine Policy, 22(3): 269–280, May 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9800013X.

Schrank:1998:FCS

[Sch98b] William E. Schrank. The failure of Canadian seasonal fishermen’s unemployment insurance reform during the 1960s and 1970s. Marine Policy, 22(1):67–81, January 1998. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000304.

Scovazzi:1990:BRM

[Sco90] Tullio Scovazzi. Book review: The 200-mile exclusive economic zone in the new law of the sea: by B. Kwiatkowska, Martinus Nijhoff, Dordrecht, 1989, 397 pp + xviii. Marine Policy, 14(4): 373–374, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090064X.

Scoullos:1992:VEE

[Sco92] M. J. Scoullos. (7) Views of the European Environmental Bureau. Marine Policy, 16(1):65–66, January 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290066X.

Scobbie:1994:IGF

[Sco94] Iain Scobbie. The ICJ and the Gulf of Fonseca: When two implies three but entails one. Marine Policy, 18(3):249–262, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900035.

Scovazzi:1995:MRR

[Sco95] Tullio Scovazzi. Management regimes and responsibility for international straits: With special reference to the Mediterranean Straits. Marine Policy, 19(2):137–152, March 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400007F.

Seeberg-Elverfeldt:1997:RPI

[SE97a] Niels-J. Seeberg-Elverfeldt. The role of private industry in implementing the Baltic Sea Joint Comprehensive Environmental Action Programme. Marine Policy, 21(5):481–491, September 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000316.

Steins:1997:IAU

[SE97b] Nathalie A. Steins and Victoria M. Edwards. Institutional analysis of UK coastal fisheries: Implications of overlapping regulations for fisheries management. Marine Policy, 21(6):535–544, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97000237.

Sebek:1990:GEF

[Seb90] Viktor Sebek. Guest Editor’s foreword. Marine Policy, 14 (3):183, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900683.

Sebek:1992:GEF

[Seb92] Viktor Sebek. Guest Editor’s foreword. Marine Policy, 16(1): 3–4, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290050Y.

Sebek:1994:I

[Seb94] Viktor Sebek. Introduction. Marine Policy, 18(2):92–93, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900108.

Senior:1996:FFS

[Sen96] Gordon Senior. Fisheries in the future: Sustainability or extinction? Marine Policy, 20(1):91–95, January 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96819221.

Shaw:1992:SCL

[Sha92] Richard Shaw. The 1989 Salvage Convention and Lloyd’s Open Form 1990: Are they working? Marine Policy, 16(2):127– 129, March 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290037P.

Sharp:1997:RAT

[Sha97] Basil M. H. Sharp. From regulated access to transferable harvesting rights: Policy insights from New Zealand. Marine Policy, 21(6):501–517, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97800334.

Schumacher:1996:LBM

[SHG96] Mary Schumacher, Porter Hoagland, and Arthur Gaines. Land-based marine pollution in the Caribbean Incentives and prospects for an effective regional protocol. Marine Policy, 20(2):99–121, March 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X95000447.

Side:1997:FNS

[Sid97] Jonathan Side. The future of North Sea oil industry abandonment in the light of the Brent Spar decision. Marine Policy, 21(1):45–52, January 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X96000450.

Skjaerseth:1992:TED

[Skj92] Jon Birger Skjaerseth. Towards the end of dumping in the North Sea: an example of effective international problem solving? Marine Policy, 16(2):130–140, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290038Q.

Skogsfors:1994:RCF

[Sko94] Sven-Erik Skogsfors. Role of citizens in formulating water quality policies. Marine Policy, 18(2):189–191, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900272.

Smith:1999:CSM

[SL99] H. D. Smith and C. S. Lalwani. The call of the sea: the marine knowledge industry in the UK. Marine Policy, 23(4–5): 397–412, July 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X99000068.

Stel:1996:BCA

[SM96] Jan H. Stel and Brian F. Mannix. A benefit-cost analysis of a regional global ocean observing system: Seawatch Europe. Marine Policy, 20(5):357–376, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000292.

Smet:1990:MCT

[Sme90] Miet Smet. Management and control of transfrontier waste movements in Belgium. Marine Policy, 14(3):228–235, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900797.

Smet:1992:NS

[Sme92] Miet Smet. (1) The North Sea. Marine Policy, 16(1):50– 51, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290055T.

Smith:1993:MME

[Smi93] Hance Smith. Managing the marine environment: the Shetland conference. Marine Policy, 17(5):332–333, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390040A.

Smith:1994:BRA

[Smi94] H. D. Smith. Book review: Atlas for marine policy in East Asian seas: edited by Joseph Morgan and Mark J. Valencia, University of California Press, Berkeley/Los Angeles/ Oxford, 1992, vi + 152 pp. Marine Policy, 18(1):80, January 1994. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X94900922.

Smith:1995:EMS

[Smi95a] Hance D. Smith. The environmental management of shipping. Marine Policy, 19(6):503–508, November 1995. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X95000333.

Smith:1995:SSE

[Smi95b] Hance D. Smith. Shipping, safety and the environment: a post-Donaldson analysis. Marine Policy, 19(6):451–452, November 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X95000285.

Smith:1999:RMS

[Smi99a] Hance D. Smith. The regional management of the seas around the United Kingdom. Marine Policy, 23(4–5):525– 535, July 1999. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X99000123.

Smith:1999:HGS

[Smi99b] Hence D. Smith. The historical geography of the seas around the British Isles. Marine Policy, 23(4–5):275–287, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000111.

Sen:1996:FCM

[SN96] Sevaly Sen and Jesper Raakjaer Nielsen. Fisheries comanagement: a comparative analysis. Marine Policy, 20 (5):405–418, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X96000280.

Sen:1997:CFC

[SN97] Sevaly Sen and Jesper Raakjoer Nielsen. Commentary on “Fisheries co-management: a comparative analysis”. Marine Policy, 21(6):545–546, November 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97800346.

Sorokin:1992:ICA

[Sor92] A. V. Sorokin. International cooperation in aquaculture: Legal aspects. Marine Policy, 16(2):99–101, March 1992. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9290031J.

Sosmena:1994:MHH

[Sos94] Gaudioso C. Sosmena. Marine health hazards in Southeast Asia. Marine Policy, 18(2):175–182, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900248.

Smith:1996:BAI

[SP96] A. K. Smith and D. A. Pollard. The best available information — some case studies from NSW, Australia, of conservation-related management responses which impact on recreational fishers. Marine Policy, 20(3):261–267, May 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000115.

Smith:1997:P

[SP97a] Hance Smith and David Pinder. Preface. Marine Policy, 21 (4):289, July 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000274.

Smith:1997:GNP

[SP97b] Hance D. Smith and David Pinder. Geostrategy and naval port systems: Frameworks for analysis. Marine Policy, 21(4): 291–308, July 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000262.

Symes:1997:IFM

[SP97c] David Symes and Jeremy Phillipson. Inshore fisheries management in the UK: Sea fisheries committees and the challenge of marine environmental management. Marine Policy, 21(3): 207–224, May 1997. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X9700002X.

Squires:1994:SGM

[Squ94] Dale Squires. Sources of growth in marine fishing industries. Marine Policy, 18(1):5–18, January 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9490085X.

Srivastava:1990:RIM

[Sri90] Chandrika Prasad Srivastava. The role of the International Maritime Organisation. Marine Policy, 14(3):243–246, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900812.

Slack:1997:GRG

[SS97] Brian Slack and John Starr. Geostrategy, resources and geopolitics: Implications for the US Navy and naval port system. Marine Policy, 21(4):309–323, July 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000110.

Soegiarto:1998:IEM

[SS98] Aprilani Soegiarto and Jan H. Stel. The Indonesian experience in marine capacity building. Marine Policy, 22(3):255– 267, May 1998. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X98000128.

Stamp:1993:ECC

[Sta93] Preben S. Stamp. European cooperation to combat marine oil pollution. Marine Policy, 17(5):430–433, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X93900617.

Stacey:1994:LRG

[Sta94] Michael L. Stacey. Legislation, regulation and government administration. Marine Policy, 18(6):501–505, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900736.

Steiner:1993:GIP

[Ste93a] Rick Steiner. Government, industry and public management of the seas in the 21st century. Marine Policy, 17(5):399– 403, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X93900558.

Stel:1993:TEO

[Ste93b] Jan H. Stel. Towards a European office for marine science and technology. Marine Policy, 17(4):309–321, July 1993. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390086I.

Stokke:1991:ATS

[Sto91a] Olav Schram Stokke. The Antarctic Treaty System in world politics: the Fridtjof Nansen Institute, Oslo, 21–23 May 1990. Marine Policy, 15(2):137–141, March 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X91900143.

Stokke:1991:TFJ

[Sto91b] Olav Schram Stokke. Transnational fishing: Japan’s changing strategy. Marine Policy, 15(4):231–243, July 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190001R.

Stormont:1994:MPC

[Sto94] William G. Stormont. Managing potential conflicts in the South China Sea. Marine Policy, 18(4):353–356, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900485.

Strathclyde:1994:CWQ

[Str94] Lord Strathclyde. Coastal water quality in the North Sea. Marine Policy, 18(2):161–164, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900221.

Sturmey:1990:PCL

[Stu90] S. G. Sturmey. Pragmatic competition and liner shipping. Marine Policy, 14(1):36–52, January 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090036Q.

Stubbs:1994:UCG

[Stu94] Captain B. B. Stubbs. The US coast guard: a unique instrument of US national security. Marine Policy, 18(6):506– 520, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900744.

Sueyoshi:1994:SDI

[Sue94] Koichi Sueyoshi. From sea of death to international environmental leadership: the case of Kitakyushu City. Marine Policy, 18(2):195–198, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900299.

Sumaila:1999:RGT

[Sum99] Ussif Rashid Sumaila. A review of game-theoretic models of fishing. Marine Policy, 23(1):1–10, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000456.

Sunderland:1994:MAN

[Sun94] Russell Sunderland. Marine administration: a new structure at the UK Department of Transport. Marine Policy, 18(6):498– 500, November 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900728.

Sun:1995:PRC

[Sun95] Kuan-Ming Sun. Policy of the Republic of China towards the South China Sea: recent developments. Marine Policy, 19(5):401–409, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500016Y.

Sun:1996:FTS

[Sun96] Kuan-Ming Sun. Freeze the tropical seas. An ice-cool prescription for the burning Spratly issues! Marine Policy, 20(3): 199–208, May 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000103.

Suter:1990:BRM

[Sut90] Keith D. Suter. Book review: Managing the frozen south: the creation and evolution of the Antarctic treaty system: by M. J. Peterson, University of California Press, Berkeley, CA, 1988, 283 pp. Marine Policy, 14(1):100, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090049W.

Suter:1993:TGT

[Sut93a] Keith Suter. Timor Gap treaty: the continuing controversy. Marine Policy, 17(4):294–302, July 1993. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9390084G.

Sutherland:1993:MCS

[Sut93b] George Sutherland. Management and control of shipping around Shetland. Marine Policy, 17(5):371–379, September 1993. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9390050D.

Tabai:1994:GCC

[Tab94] Ieremia Tabai. Global climatic change: Impact on coastal environments of the South Pacific. Marine Policy, 18(2):183– 185, March 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900256.

Tango:1990:BRM

[Tan90] Gerardo G. Tango. Book review: Marine policy for America, second edition: by Gerard Mangone, Taylor and Francis, New York and London, 1988, 349 pp. $40. Marine Policy, 14(6):534– 536, November 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090024L.

Tan:1998:ACF

[Tan98] Sen Min Tan. ASEAN–Canada Fisheries Post-Harvest Technology Project — Phase II Project Implementation. Marine Policy, 22(6):469–476, November 1998. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000220.

Cheng:1991:ILS

[tCCT91] Hsiang tai Cheng, George K. Criner, and Ralph E. Townsend. International lobster supply: Benefits of Canadian seasonality. Marine Policy, 15(5):363–369, September 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190091O.

Thanopoulou:2021:WPF

[Tha21] Helen A. Thanopoulou. What price the flag? The terms of competitiveness in shipping. Marine Policy, 22(4–5):359– 374, ???? 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X97000432.

Thebaud:1997:TMF

[Thé97] Olivier Thébaud. Transboundary marine fisheries management. Recent developments and elements of analysis. Marine Policy, 21(3):237–253, May 1997. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X97000055.

Thomas:1994:NOC

[Tho94] B. J. Thomas. The need for organizational change in seaports. Marine Policy, 18(1):69–78, January 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900906.

Till:1991:BRM

[Til91] Geoffrey Till. Book review: Maritime strategy and European security: Eric Grove, Common security studies no 2, Brassey’s, London, 1990, pp 168, £19.95. Marine Policy, 15(5):376, September 1991. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9190096T.

Tolba:1990:GAH

[Tol90] Mostafa K. Tolba. The global agenda and the hazardous wastes challenge. Marine Policy, 14(3):205–209, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900742.

Tolba:1992:PU

[Tol92] Mostafa Tolba. (4) Policy of UNEP. Marine Policy, 16(1): 61, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290063U.

Topfer:1990:ADC

[Töp90] Klaus Töpfer. Assistance to developing countries in the implementation of international rules. Marine Policy, 14(3):259– 263, May 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X90900856.

Topfer:1992:B

[Töp92] Klaus Töpfer. (3) The Baltic. Marine Policy, 16(1):53– 54, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290057V.

Townsend:1992:BIT

[Tow92] Ralph E. Townsend. Bankable individual transferable quotas. Marine Policy, 16(5):345–348, September 1992. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X92900027.

Towfighi:1994:IPM

[Tow94] Parviz S. Towfighi. Integrated planning and management of coastal areas. Marine Policy, 18(2):107–111, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900140.

Townsend:1995:FSG

[Tow95a] Ralph E. Townsend. Fisheries self-governance: corporate or cooperative structures? Marine Policy, 19(1):39–45, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9592571N.

Townsend:1995:TDS

[Tow95b] Ralph E. Townsend. Transferable dynamic stock rights. Marine Policy, 19(2):153–158, March 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9400006E.

Treves:1990:BRL

[Tre90] Tullio Treves. Book review: The law of deep sea-bed mining: by Said Mahmoudi Almquist & Wicksell International, Stockholm, 1987, 362pp. Marine Policy, 14(5):457–459, September 1990. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9090012G.

Trippier:1990:WMD

[Tri90] David A. Trippier. Waste management and the development of standards. Marine Policy, 14(3):214–218, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X90900764.

Trippier:1992:SP

[Tri92] David Trippier. (1) Science and policy. Marine Policy, 16(1): 58, January 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X92900603.

Valencia:1990:SJT

[Val90] Mark J. Valencia. Sea of Japan: Transnational marine resource issues and possible cooperative responses. Marine Policy, 14(6):507–525, November 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9090020R.

Valero:1994:SAD

[Val94] Gerardo M. C. Valero. Spratly archipelago dispute: Is the question of sovereignty still relevant? Marine Policy, 18(4): 314–344, July 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900469.

Vallega:1995:TSM

[Val95] Adalberto Vallega. Towards the sustainable management of the Mediterranean Sea. Marine Policy, 19(1):47–64, January 1995. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9592572O.

VanErmen:1990:PEH

[Van90] Raymond Van Ermen. Protecting the environment from hazardous substances. Marine Policy, 14(3):278–282, May 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090089A.

VanDyke:1991:MEW

[Van91] Jon M. Van Dyke. Military exclusion and warning zones on the high seas. Marine Policy, 15(3):147–169, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190059K.

VanDyke:1992:BRR

[Van92] Jon M. Van Dyke. Book review: The regime of islands in international law: Hiran J. Jayewardene, Martinus Nijhoff, Dordrecht, 1990, 572pp, US$159. Marine Policy, 16(5):406– 407, September 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290009E.

VanDyke:1996:ASD

[Van96] Jon M. Van Dyke. The Aegean Sea dispute: options and avenues. Marine Policy, 20(5):397–404, September 1996. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96000255.

vandenMuijsenberg:1994:CDM

[vdM94] J. H. A. van den Muijsenberg. Contaminated dredged material: Cross-border initiatives by the port of Rotterdam. Marine Policy, 18(2):192–194, March 1994. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900280.

vanEekelen:1994:NCW

[vE94] Willem van Eekelen. Naval cooperation in WEU. Marine Policy, 18(6):534–537, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900779.

Verlaan:1997:NSM

[Ver97] Philomène A. Verlaan. New seafloor mapping technology and article 76 of the 1982 United Nations Convention on the law of the sea. Marine Policy, 21(5):425–434, September 1997. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000080.

Vartanov:1997:NLC

[VH97] Raphael Vartanov and Charles D. Hollister. Nuclear legacy of the cold war. Russian policy and ocean disposal. Marine Policy, 21(1):1–15, January 1997. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X96000449.

Wagner:1995:SCP

[Wag95] Eric Wagner. Submarine cables and protections provided by the law of the sea. Marine Policy, 19(2):127–136, March 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9400002A.

Wilson:1996:CPM

[WAK96] James A. Wilson, James Acheson, and Peter Kleban. Chaos and parametric management. Marine Policy, 20(5):429–438, September 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000267.

Wall:1993:IOP

[Wal93] James F. Wall. Intergovernmental oil pollution liability and compensation: theory and practice. Marine Policy, 17(5):473– 478, September 1993. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9390068E.

Wilson:1994:CCC

[WAMK94] James A. Wilson, James M. Acheson, Mark Metcalfe, and Peter Kleban. Chaos, complexity and community management of fisheries. Marine Policy, 18(4):291–305, July 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900442.

Watt:1990:BRE

[Wat90a] D. C. Watt. Book review: Expansion of third world navies: by Michael Morris, Macmillan, London, 1987, xiii + 294 pp, £35.00. Marine Policy, 14(1):96–97, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090045S.

Watt:1990:IMP

[Wat90b] D. Cameron Watt. An integrated marine policy: a meaningful concept? Marine Policy, 14(4):299–304, July 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090051R.

Watt:1990:BROa

[Wat90c] Donald Cameron Watt. Book review: Ocean yearbook 6: edited by Elisabeth Mann Borgese and Norton Ginsburg, University of Chicago Press, Chicago, IL, 1987. Marine Policy, 14 (4):372, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090062V.

Watt:1990:BROb

[Wat90d] Donald Cameron Watt. Book review: Ocean yearbook 7: edited by Elisabeth Mann Borgese, Norton Ginsburg and Joseph R. Morgan, University of Chicago Press, Chicago, IL, 1987. Marine Policy, 14(4):372–373, July 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090063W.

Williams:1991:PEC

[WBN91] D. C. Williams, Stephen C. Bushardt, and Edward Nissan. Packaging for export: a case study of underutilized species of finfish. Marine Policy, 15(1):15–22, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190039E.

Woodhatch:1999:GAF

[WC99] Libby Woodhatch and Kevin Crean. The gentleman’s agreements: a fisheries management case study from the Southwest of England. Marine Policy, 23(1):25–35, January 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X98000013.

Wright:1991:DNF

[WD91] Andrew Wright and David J. Doulman. Drift-net fishing in the South Pacific: From controversy to management. Marine Policy, 15(5):303–329, September 1991. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190081L.

Wecker:1990:LSI

[Wec90] Miranda Wecker. Law of the Sea Institute annual meeting: Noordwijk aan Zee, The Netherlands, 12–15 June 1989. Marine Policy, 14(4):362–365, July 1990. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090058Y.

Wettestad:1994:SPI

[Wet94] Jørgen Wettestad. Science, politics and institutional design: the case of the North-East Atlantic land-based pollution regime. Marine Policy, 18(3):219–232, May 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X94900019.

Whitmarsh:1990:TCM

[Whi90] David Whitmarsh. Technological change and marine fisheries development. Marine Policy, 14(1):15–22, January 1990. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9090034O.

Williams:1996:PME

[Wil96] Caroline Williams. Protecting the marine environment from land-based activities: a global programme of action. Marine Policy, 20(1):95–97, January 1996. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X96819233.

Wallace:1998:ESC

[WLD98] I. F. Wallace, R. K. Lindner, and D. D. Dole. Evaluating stock and catchability trends: Annual average catch per unit effort is an inadequate indicator of stock and catchability trends in fisheries. Marine Policy, 22(1):45–55, January 1998. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X97000286.

Wooldridge:1999:EMP

[WMH99] Christopher F. Wooldridge, Christopher McMullen, and Vicki Howe. Environmental management of ports and harbours — implementation of policy through scientific monitoring. Marine Policy, 23(4–5):413–425, July 1999. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000554.

Whitmarsh:1999:RBC

[WNJ99] David Whitmarsh, James Northen, and Shabbar Jaffry. Recreational benefits of coastal protection: a case study. Marine Policy, 23(4–5):453–463, July 1999. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0308597X98000517.

Wonham:1996:SRR

[Won96] J. Wonham. Some recent regulatory developments in IMO for which there are corresponding requirements in the United Nations convention on the law of the sea. A challenge to be met by the states parties? Marine Policy, 20(5):377–388, September 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X96000309.

Wonham:2021:ASB

[Won21] J. Wonham. Agenda 21 and sea-based pollution: Apportunity or apathy? Marine Policy, 22(4–5):375–391, ???? 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X97000493.

Woods:1994:GOO

[Woo94] John Woods. The global ocean observing system. Marine Policy, 18(6):445–452, November 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X94900647.

Wood-Thomas:1994:LBM

[WT94] B. C. Wood-Thomas. Land-based marine pollution and coastal zone management: Role of state and local government in the USA. Marine Policy, 18(2):165–174, March 1994. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9490023X.

Wang:1992:MFR

[WyZ92] Stanley D. H. Wang and Bing yi Zhan. Marine fishery resource management in PR China. Marine Policy, 16(3):197– 209, May 1992. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9290081Y.

Yates:1995:PHT

[YCS95] John Yates, Paul Cunningham, and Denis Smith. Polar hydrocarbon technologies: future developments. Marine Policy, 19(5):419–436, September 1995. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0308597X9500006R.

Yates:1994:MST

[YR94] John Yates and Ganpat S. Roonwal. Marine science and technology in India: Current status. Marine Policy, 18(1):59– 68, January 1994. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X94900892.

Young:1996:RIF

[YSM96] James A. Young, Andrew P. Smith, and James F. Muir. Representing the individual fishermen: an attitudinal perspective on one PO’s membership. Marine Policy, 20(2):157– 169, March 1996. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X96000061.

Yu:1991:MFM

[Yu91] Huming Yu. Marine fishery management in PR China. Marine Policy, 15(1):23–32, January 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X9190040I.

Zhijie:1990:CZP

[ZC90] Fan Zhijie and P. P. Côté. Coastal zone of PR China: Management approaches and institutions. Marine Policy, 14(4): 305–314, July 1990. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0308597X9090052S.

Zhijie:1991:PDM

[ZC91] Fan Zhijie and R. P. Côté. Population, development and marine pollution in China. Marine Policy, 15(3):210–219, May 1991. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X9190064I.

Zhixiong:1991:CED

[Zhi91] Wang Zhixiong. China and the exploitation of deep seabed polymetallic nodules. Marine Policy, 15(2):132–135, March 1991. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X9190012Z.